### v2

# InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection

Abstract
The prevalence of sarcasm in social media, conveyed through text-image combinations, presents significant challenges for sentiment analysis and intention mining. Current multi-modal sarcasm detection methods have been proven to struggle with biases from spurious cues, leading to a superficial understanding of the complex interactions between text and image. To address these issues, we propose InterCLIP-MEP, a robust framework for multi-modal sarcasm detection. InterCLIP-MEP introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone, enhancing sample representations by embedding cross-modality information in each encoder. Furthermore, a novel training strategy is designed to adapt InterCLIP for a Memory-Enhanced Predictor (MEP). MEP uses dynamic dual-channel memory to store valuable historical knowledge of test samples and then leverages this memory as a non-parametric classifier to derive the final prediction. By using InterCLIP to encode text-image interactions more effectively and incorporating MEP, InterCLIP-MEP offers a more robust recognition of multi-modal sarcasm. Experiments demonstrate that InterCLIP-MEP achieves state-of-the-art performance on the MMSD2.0 benchmark. Code and data are available at https://github.com/CoderChen01/InterCLIP-MEP.
Introduction
Sarcasm, with its inherent subtlety and complexity, plays a significant role in human communication by often conveying irony, mockery, or hidden intentions (Muecke 1982; Gibbs and O’Brien 1991; Gibbs and Colston 2007). The automatic detection of sarcasm from text has been a significant research focus, aiding tasks such as sentiment analysis and intent mining (Pang, Lee et al. 2008; Tsur, Davidov, and Rappoport 2010; Bouazizi and Ohtsuki 2015; Poria et al. 2016; Amir et al. 2016; Wu et al. 2018; Tay et al. 2018; Baziotis et al. 2018). With the rapid development of social media platforms like Twitter and Reddit, the need for multi-modal sarcasm detection has become increasingly apparent. Users frequently use text-image combinations to convey their messages, making it essential to develop methods to interpret these multi-modal communications. The task of multi-modal sarcasm detection presents significant challenges, requiring the system to accurately capture and comprehend the intricate interplay between textual and visual information to identify sarcasm cues.
A pivotal development in this field was the introduction of the Multi-Modal Sarcasm Detection (MMSD) benchmark by Cai, Cai, and Wan (2019). This benchmark has become a cornerstone for subsequent research, inspiring a range of advanced techniques aimed at improving sarcasm detection performance (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Liu, Wang, and Li 2022; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024). The above methods have explored various approaches, including the use of Transformer-based models for high-quality representation extraction (Pan et al. 2020), OCR and object detection technologies to enhance image understanding (Qiao et al. 2023), and the incorporation of external emotional knowledge (Liu, Wang, and Li 2022). Graph neural networks have also been utilized to construct dependency and cross-modal graphs for improved sarcasm identification (Liu, Wang, and Li 2022; Liang et al. 2021, 2022).
Figure 1: Comparison of InterCLIP-MEP and other multi-modal sarcasm detection pipelines.
Despite these advancements, the existing methods have shown limitations, often learning biases from spurious cues present in the MMSD benchmark (Qin et al. 2023). To address this, Qin et al. (2023) introduced MMSD2.0, a refined benchmark that eliminates these spurious cues and corrects mislabeled samples. Their work demonstrated the need for more reliable benchmarks and robust detection methods.
Building on these insights, we propose InterCLIP-MEP, a novel framework designed to enhance the reliability of multi-modal sarcasm detection. We present a comparison of existing methods with InterCLIP-MEP in Figure 1. As shown in Figure 1(a) and Figure 1(b), some existing methods use separate uni-modal encoders, such as BERT (Devlin et al. 2019) for text and ViT (Dosovitskiy et al. 2020) for visual modalities, as backbones (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024). These encoded samples are fused using various techniques to obtain the final sample representations. We argue that this approach makes it difficult for the model to capture the subtle interactive information between text and images. Although Qin et al. (2023) utilized a multi-modal pre-trained backbone, CLIP (Radford et al. 2021), to encode samples and demonstrated notable results on the MMSD2.0 benchmark, the original CLIP only aligns text and images directly, lacking the nuanced semantic space required for sarcasm detection. Therefore, InterCLIP-MEP introduces a refined variant, Interactive CLIP (InterCLIP), as the backbone to enhance sample encoding for this specific task. Inspired by the foundational work of Ganz et al. (2024), InterCLIP enhances the capture of interactive information between text and image by embedding representations from one modality into the encoder of the other, leading to a more nuanced understanding of multi-modal sarcasm cues.
After encoding the samples, as depicted in Figure 1(a), many approaches directly identify sarcastic samples through a classification module. This often leads to high prediction entropy and incorrect results, especially for hard cases with greater uncertainty. To the best of our knowledge, few methods consider leveraging historical knowledge of samples to aid in detection. As shown in Figure 1(b), only Wei et al. (2024) constructed a static semantic space that stores representations of all training samples. When identifying sarcastic samples, they query $k$-nearest neighbors in this semantic space and use a voting mechanism to derive the final predictions. However, their static semantic space struggles to handle out-of-distribution issues.
As shown in Figure 1(c), InterCLIP-MEP uses a novel training strategy to adapt InterCLIP for a Memory-Enhanced Predictor (MEP). Based on the sample representations encoded by InterCLIP, InterCLIP-MEP trains a classification module to assign pseudo-labels to samples. Simultaneously, it trains a projection module to map samples into a latent space. In this latent space, same-class samples are pulled closer together, while different-class samples are pushed further apart. Additionally, InterCLIP-MEP fine-tunes the weights of the self-attention modules in InterCLIP’s text and vision encoders using LoRA (Hu et al. 2022), building upon the original CLIP weights. Building on InterCLIP and its training strategy, MEP leverages historical knowledge of test samples to identify sarcasm. It maintains a dynamic, fixed-length dual-channel memory, with one channel storing historical information of non-sarcastic samples and the other storing information of sarcastic samples. MEP uses the learned InterCLIP to extract sample representations, calculate pseudo-labels with the classification module, and compute prediction entropy, while also obtaining projection features through the projection module. Based on the pseudo-labels and prediction entropy, MEP dynamically updates the projected features of samples into the corresponding memory channels, treating these features as the historical knowledge of the test samples. Specifically, MEP consistently updates the memory with projected features of samples that have lower prediction entropy. MEP utilizes this memory as a non-parametric classifier to compute the final prediction for each sample.
Our experiments demonstrate the robustness and innovation of InterCLIP-MEP in accurately identifying sarcasm through sophisticated interaction and memory mechanisms. Our contributions can be summarized as follows:
  • We propose InterCLIP-MEP, a novel framework for reliable multi-modal sarcasm detection, which introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone, improving the encoding of text-image interactions by embedding cross-modality information within each encoder.
  • We design a novel training strategy to adapt InterCLIP for a Memory-Enhanced Predictor (MEP). This strategy enables the integration of dynamic dual-channel memory, which stores valuable historical knowledge of test samples and leverages this memory for more precise and informed predictions.
  • Our extensive experiments on the more reliable benchmark, MMSD2.0, demonstrate that InterCLIP-MEP outperforms existing state-of-the-art methods.
Figure 2: Overview of our framework. The left part of the figure shows the overall structure and training process of Interactive-CLIP. The right part illustrates the process of MEP identifying sarcastic samples.
Related Work
Early work in sarcasm detection primarily focused on textual data (Bouazizi and Ohtsuki 2015; Poria et al. 2016; Amir et al. 2016; Wu et al. 2018; Tay et al. 2018; Baziotis et al. 2018). With the rise of social media, detecting sarcasm from text-image data has become increasingly challenging, leading to the development of several multi-modal approaches (Schifanella et al. 2016; Cai, Cai, and Wan 2019; Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Liu, Wang, and Li 2022; Qin et al. 2023; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024). Schifanella et al. (2016) were among the first to use multi-modal posts from social platforms to identify sarcasm clues from both text and images. Cai, Cai, and Wan (2019) introduced the MMSD benchmark for multi-modal sarcasm detection, demonstrating its effectiveness with a hierarchical fusion model incorporating image attributes. This benchmark has been widely used in subsequent studies (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Liu, Wang, and Li 2022; Qin et al. 2023; Wen, Jia, and Yang 2023; Tian et al. 2023; Wei et al. 2024). Further research expanded on MMSD, exploring dependencies and external knowledge, using advanced models, and introducing graph-based methods to enhance performance (Xu, Zeng, and Mao 2020; Pan et al. 2020; Liang et al. 2021, 2022; Liu, Wang, and Li 2022). However, MMSD has been proven to contain spurious cues that could bias models (Qin et al. 2023). To address this, Qin et al. (2023) introduced MMSD2.0 benchmark, which removed these cues and corrected mislabeled samples. They observed a significant performance drop when re-evaluating state-of-the-art methods on MMSD2.0 and proposed the need for more stable multi-modal sarcasm detection methods. Motivated by the work of Qin et al. (2023), we introduce the InterCLIP-MEP framework for more reliable multi-modal sarcasm detection.
Methodology
In this paper, we introduce InterCLIP-MEP, a novel framework for multi-modal sarcasm detection that combines interactive cross-modal encoding with memory-enhanced prediction. An overview of InterCLIP-MEP is illustrated in Figure 2. Our approach leverages a variant of CLIP, called Interactive CLIP (InterCLIP), which enhances the original CLIP architecture by embedding representations from one modality into the encoder of the other modality, resulting in more interactive representations. InterCLIP enables a nuanced understanding of the interaction between text and image. To facilitate memory-enhanced prediction, we design a novel training strategy, integrating a classification module and a projection module, with InterCLIP serving as the backbone. These modules are trained to identify sarcastic samples while constructing a latent space where features of the same class are clustered together and those of different classes are distinctly separated. For the prediction phase, InterCLIP-MEP employs a Memory-Enhanced Predictor (MEP) that leverages historical knowledge for final predictions. Instead of relying solely on the classification module to identify sarcasm, MEP compares the features of the current sample with those stored in memory, thereby enhancing the accuracy of sarcasm detection.
Interactive CLIP
We introduce Interactive CLIP (InterCLIP), which can condition text or image encoder to interact with representations from the other modality, enhancing the capture of interactive information crucial for identifying sarcasm.
The input to InterCLIP is a text-image pair $\mathcal{P}=(T,I)$, where $T$ represents a piece of text and $I$ represents an image. Here, for simplicity, we do not consider the case of batch inputs. The text encoder $\mathcal{T}$ extracts the vanilla text representation $\mathbf{F}_{t}$:
$\mathbf{F}_{t}=\mathcal{T}(T)=\{h^{t}_{\textrm{bos}}(t_{\textrm{bos}}),\\ h^{t}_{1}(t_{1}),\dots,h^{t}_{n}(t_{n}),h^{t}_{\textrm{eos}}(t_{\textrm{eos}})\},$ (1)
where $t_{i}$ denotes a text token, $n$ is the length of $T$ after tokenization, $t_{\textrm{bos}}$ and $t_{\textrm{eos}}$ are special tokens required by the text encoder. Here, $h^{t}_{i}(\cdot)\in\mathbb{R}^{d_{t}}$ represents the $d_{t}$-dimensional encoded representation of the corresponding token $t_{i}$, with $i$ ranging from 1 to $n$, including the beginning-of-sequence (bos) and end-of-sequence (eos) tokens. The vision encoder $\mathcal{V}$ extracts the vanilla image representation $\mathbf{F}_{v}$:
$\mathbf{F}_{v}=\mathcal{V}(I)=\{h^{v}_{\textrm{cls}}(p_{\textrm{cls}}),h^{v}_{1}(p_{1}),\dots,h^{v}_{m}(p_{m})\},$ (2)
where $I$ is processed into multiple patches $p_{i}$, $m$ is the number of patches, and $p_{\textrm{cls}}$ is a special token required by the visual encoder. Here, $h^{v}_{i}(\cdot)\in\mathbb{R}^{d_{v}}$ represents the $d_{v}$-dimensional encoded representation of the corresponding $p_{i}$, with $i$ ranging from 1 to $m$, including the classification (cls) token. Specifically, both $\mathbf{F}_{t}$ and $\mathbf{F}_{v}$ are representations from the final layer outputs of their respective encoders. Conditioning on $\mathbf{F}_{t}$ or $\mathbf{F}_{v}$, we can obtain the interactive text representation $\tilde{\mathbf{F}}_{t}$ or the interactive image representation $\tilde{\mathbf{F}}_{v}$:
$\tilde{\mathbf{F}}_{t}=\mathcal{T}(T|\mathbf{F}_{v}),\tilde{\mathbf{F}}_{v}=\mathcal{V}(V|\mathbf{F}_{t}).$ (3)
We use $\tilde{h}^{t}_{i}(\cdot)\in\mathbb{R}^{d_{t}}$ and $\tilde{h}^{v}_{i}(\cdot)\in\mathbb{R}^{d_{v}}$ to denote the re-encoded interactive representations of each text token and image patch, respectively.
Specifically, we condition only the top-$n$ self-attention layers of the text or vision encoder, where $n$ is a hyperparameter that will be analyzed in subsequent sections. The structure of the conditioned self-attention layers is illustrated in Figure 3. We denote the input representations to the self-attention layers as $\mathbf{H}_{t/v}$, which are the output representations from the previous layer. The previous layer can either be conditioned or non-conditioned. Due to the dimensional mismatch between the embedded representations $\mathbf{F}_{v/t}$ and the corresponding encoder representation space, we introduce an adapting projection layer $\mathcal{F}_{t/v}$ to project $\mathbf{F}_{v/t}$ into the appropriate representation space.
To fuse the input representations $\mathbf{H}_{t/v}$ with the projected embedded representations $\mathbf{F}_{v/t}^{{}^{\prime}}=\mathcal{F}_{t/v}(\mathbf{F}_{v/t})$, we concatenate them and feed them into the attention layer to obtain the transformed representations. We then extract the transformed input representations $\mathbf{H}^{{}^{\prime}}_{t/v}$ from the output. Following Ganz et al. (2024), we apply a gated projection layer $\mathcal{G}_{t/v}$ along with the self-attention’s projection head $\mathcal{H}_{t/v}$ using a learnable gating mechanism to compute the self-attention output representation $\mathbf{H}^{{}^{\prime\prime}}_{t/v}$.
Given the similarity between the self-attention layers of the vision encoder and the text encoder, we use the text encoder $\mathcal{T}$ to illustrate the process as follows:
$\mathbf{F}_{v}^{{}^{\prime}}=\mathcal{F}_{t}(\mathbf{F}_{v}),\mathbf{F}_{v}\in\mathbb{R}^{{m}\times{d_{v}}},\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{{m}\times{d_{t}}},$ (4)
$\mathbf{H}^{{}^{\prime}}_{t}=\textrm{Attn}_{t}(\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}})_{\textrm{[:n]}},\\ \mathbf{H}_{t},\mathbf{H}^{{}^{\prime}}_{t}\in\mathbb{R}^{{n}\times{d_{t}}},\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{{(n+m)}\times{d_{t}}},$ (5)
$\mathbf{H}^{{}^{\prime\prime}}_{t}=\mathcal{H}_{t}(\mathbf{H}^{{}^{\prime}}_{t})+\mathcal{G}_{t}(\mathbf{H}^{{}^{\prime}}_{t})\cdot\textrm{tanh}(\beta_{t}),\mathbf{H}^{{}^{\prime\prime}}_{t}\in\mathbb{R}^{{n}\times{d_{t}}}.$ (6)
Here, $\oplus$ denotes the concatenation operation, and $\beta_{t}$ is a learnable parameter initialized to 0 to ensure training stability. The subsequent computation follows the original CLIP (Radford et al. 2021), ultimately yielding the interactive representations $\tilde{\mathbf{F}}_{t}$.
Figure 3: Structure of the conditional self-attention layer.
InterCLIP has three distinct modes to achieve the final fused feature $\tilde{h}^{f}\in\mathbb{R}^{d_{t}+d_{v}}$:
  • T2V: The text encoder extracts the vanilla text representation $\mathbf{F}_{t}$. This text representation is then embedded into the vision encoder to obtain the conditioned vision representation $\tilde{\mathbf{F}}_{v}$. The final fused representation $\tilde{h}^{f}$ is formed by concatenating $h^{t}_{\textrm{eos}}(t_{\textrm{eos}})$ and $\tilde{h}^{v}_{\textrm{cls}}(p_{\textrm{cls}})$.
  • V2T: The vision encoder extracts the vanilla image representation $\mathbf{F}_{v}$. This image representation is then embedded into the text encoder to obtain the conditioned text representation $\tilde{\mathbf{F}}_{t}$. The final fused representation $\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}(t_{\textrm{eos}})$ and $h^{v}_{\textrm{cls}}(p_{\textrm{cls}})$.
  • Two-way (TW): The text and vision encoders extract the vanilla text and image representations $\mathbf{F}_{t}$ and $\mathbf{F}_{v}$, respectively. These representations are then embedded into the vision and text encoders to obtain the conditioned image representation $\tilde{\mathbf{F}}_{v}$ and text representation $\tilde{\mathbf{F}}_{t}$. The final fused representation $\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}(t_{\textrm{eos}})$ and $\tilde{h}^{v}_{\textrm{cls}}(p_{\textrm{cls}})$.
We will analyze the effectiveness of these three modes of interaction in the experimental analysis.
Training Strategy
As shown in Figure 2 (left), to adapt InterCLIP for memory-enhanced prediction, we introduce a novel training strategy. Using InterCLIP as the backbone to obtain fused features of the samples, we introduce a classification module and a projection module.
The classification module $\mathcal{F}_{c}$ calculates the probability of a sample being sarcastic or non-sarcastic:
$\hat{y}=\textrm{softmax}(\mathcal{F}_{c}(\tilde{H}^{f})),$ (7)
where $\tilde{H}^{f}\in\mathbb{R}^{{N}\times(d_{t}+d_{v})}$ represents the fused representations of a batch of samples, and $N$ denotes the batch size. We optimize $\mathcal{F}_{c}$ using binary cross-entropy loss:
$\mathcal{L}^{c}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i,1})+(1-y_{i})\log(1-\hat{y}_{i,0})\right],$ (8)
where $y_{i}$ denotes the label of the $i$-th sample, with sarcastic labeled as 1 and non-sarcastic as 0, and $\hat{y}_{i}$ denotes the prediction for the $i$-th sample. The projection module $\mathcal{F}_{p}$ maps $\tilde{H}^{f}$ into a latent feature space:
$\hat{H}^{f}=\textrm{norm}(\mathcal{F}_{p}(\tilde{H}^{f})),\quad\hat{H}^{f}\in\mathbb{R}^{{N}\times d_{f}},$ (9)
where $\textrm{norm}(\cdot)$ denotes L2 normalization, and $d_{f}$ represents the dimension of the projected features. In this space, cosine distance between features of the same class is minimized, while the distance between features of different classes is maximized. We use a label-aware cosine similarity loss to optimize $\mathcal{F}_{p}$:
$\mathcal{L}^{p}=\textrm{mean}(\hat{H}^{f}_{P}\cdot\hat{H}^{f^{T}}_{N})+\textrm{mean}(1-\hat{H}^{f}_{P}\cdot\hat{H}^{f^{T}}_{P})\\ +\textrm{mean}(1-\hat{H}^{f}_{N}\cdot\hat{H}^{f^{T}}_{N}),$ (10)
where $\hat{H}^{f}_{P}$ and $\hat{H}^{f}_{N}$ represent the projected fused features of positive and negative samples, respectively.
We fully train the modules $\mathcal{F}_{c}$, $\mathcal{F}_{p}$, the adapting projection layer $\mathcal{F}_{t/v}$, the gated projection layer $\mathcal{G}_{t/v}$, and the parameters $\beta_{t/v}$. We use LoRA (Hu et al. 2022) to fine-tune parts of the weight matrices $\mathbf{W}$ in the self-attention modules of all encoders, specifically various combinations of $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$. We consider $\mathbf{W}$ and the rank $r$ of LoRA as hyperparameters for our study. All modules are optimized by minimizing the joint loss:
$\mathcal{L}=\mathcal{L}^{c}+\mathcal{L}^{p}.$ (11)
Memory-Enhanced Predictor
Algorithm 1 Memory-Enhanced Predictor
Input: Memory size $L$, Learned InterCLIP model, classification module $\mathcal{F}_{c}$ and projection module $\mathcal{F}_{p}$
Output: Final prediction $\hat{y}^{p}$
1:  Initialize memory $\mathcal{M}\in\mathbf{0}^{2\times L\times d_{f}}$
2:  Initialize index $\mathcal{I}\in\mathbf{0}^{2}$
3:  Initialize entropy records $\mathcal{C}\in\mathbf{0}^{2\times L}$
4:  for $i\leftarrow 1$ to $N_{\textrm{test}}$ do
5:     $\tilde{h}^{f}_{i}\leftarrow\textrm{InterCLIP}(\mathcal{P}_{i})$
6:     $\hat{y}_{i}\leftarrow\textrm{softmax}(\mathcal{F}_{c}(\tilde{h}^{f}_{i}))$
7:     $\ell_{\textrm{pse}_{i}}\leftarrow\arg\max_{j}(\hat{y}_{i,j}),j\in\{0,1\}$
8:     $c_{i}\leftarrow-\hat{y}_{i,0}\log{\hat{y}_{i,0}}-\hat{y}_{i,1}\log{\hat{y}_{i,1}}$
9:     $\hat{h}^{f}_{i}\leftarrow\textrm{norm}(\mathcal{F}_{p}(\tilde{h}^{f}_{i}))$
10:     if $\mathcal{I}[\ell_{\textrm{pse}_{i}}]<L$ then
11:        $\mathcal{M}[\ell_{\textrm{pse}_{i}}][\mathcal{I}[\ell_{\textrm{pse}_{i}}]]\leftarrow\hat{h}^{f}_{i}$
12:        $\mathcal{C}[\ell_{\textrm{pse}_{i}}][\mathcal{I}[\ell_{\textrm{pse}_{i}}]]\leftarrow c_{i}$
13:        $\mathcal{I}[\ell_{\textrm{pse}_{i}}]\leftarrow\mathcal{I}[\ell_{\textrm{pse}_{i}}]+1$
14:     else
15:        $j\leftarrow\textrm{GetMaxIdx}(\mathcal{C}[\ell_{\textrm{pse}_{i}}])$
16:        if $c_{i}<\mathcal{C}[\ell_{\textrm{pse}_{i}}][j]$ then
17:           $\mathcal{M}[\ell_{\textrm{pse}_{i}}][j]\leftarrow\hat{h}^{f}_{i}$
18:           $\mathcal{C}[\ell_{\textrm{pse}_{i}}][j]\leftarrow c_{i}$
19:        end if
20:     end if
21:     $\textrm{logits}\leftarrow[\sum_{k=0}^{\mathcal{I}[0]}(\hat{h}^{f}_{i}\mathcal{M}[0]^{T})_{k},\sum_{k=0}^{\mathcal{I}[1]}(\hat{h}^{f}_{i}\mathcal{M}[1]^{T})_{k}]$
22:     $\hat{y}^{p}_{i}\leftarrow\textrm{softmax}(\textrm{logits})$
23:     yield $\hat{y}^{p}_{i}$
24:  end for
As depicted in Figure 2 (right), we present a Memory-Enhanced Predictor (MEP) that builds upon the learned InterCLIP, along with the classification module and the projection module, leveraging the valuable historical knowledge of test samples to enhance sarcasm detection.
The detailed computational process of MEP is provided in Algorithm 1, where $N_{\textrm{test}}$ denotes the number of test samples. MEP uses the trained InterCLIP to extract fused features of the samples. It utilizes the classification module $\mathcal{F}_{c}$ to assign a pseudo-label $\ell_{\textrm{pse}_{i}}$ to each sample $\mathcal{P}_{i}$ and the projection module $\mathcal{F}_{p}$ to obtain the sample’s projected feature $\hat{h}^{f}_{i}$. To store valuable projected features of test samples as historical knowledge, MEP maintains a dynamic fixed-length dual-channel memory $\mathcal{M}\in\mathcal{R}^{2\times L\times d_{f}}$, where $L$ is the memory length per channel. The first channel stores projected features of non-sarcastic samples, while the second channel stores those of sarcastic samples. Based on the pseudo-label $\ell_{\textrm{pse}_{i}}$, the appropriate memory channel $\mathcal{M}[\ell_{\textrm{pse}_{i}}]$ is selected for updating. If the selected channel has available space, the sample’s projected features are added directly, and the prediction entropy is recorded. If the memory is full, the prediction entropy of all samples in the memory is compared with that of the current sample, replacing the samples with the highest entropy as necessary. Finally, the current sample’s projected features are combined with the historical features stored in both memory channels $\mathcal{M}$ using cosine similarity to yield the final prediction.
All Sarcastic Non-sarcastic
Train 19812 9572 10240
Validation 2410 1042 1368
Test 2409 1037 1372
Table 1: Statistics of the dataset from the MMSD2.0 benchmark (Qin et al. 2023).
Modality Method Acc. (%) F1 (%) P (%) R (%)
Text TextCNN* (Kim 2014) 71.61 69.52 64.62 75.22
Bi-LSTM* (Graves and Schmidhuber 2005) 72.48 68.05 68.02 68.08
SMSD* (Xiong et al. 2019) 73.56 69.97 68.45 71.55
RoBERTa* (Liu et al. 2019) 79.66 76.21 76.74 75.70
Image ResNet* (He et al. 2015) 65.50 57.58 61.17 54.39
ViT* (Dosovitskiy et al. 2020) 72.02 69.72 65.26 74.83
Text-Image HFM* (Cai, Cai, and Wan 2019) 70.57 66.88 64.84 69.05
Att-BERT* (Pan et al. 2020) 80.03 77.04 76.28 77.82
CMGCN* (Liang et al. 2022) 79.83 76.90 75.82 78.01
HKE* (Liu, Wang, and Li 2022) 76.50 72.25 73.48 71.07
DIP (Wen, Jia, and Yang 2023) 80.59 78.23 75.52 81.14
DynRT (Tian et al. 2023) 70.37 68.55 63.02 75.15
Multi-view CLIP* (Qin et al. 2023) 85.64 84.10 80.33 88.24
G2SAM (Wei et al. 2024) 79.43 78.07 72.04 85.20
InterCLIP-MEP w/o Inter ($L=1024$) 86.05 84.81 79.83 90.45
InterCLIP-MEP w/ TW ($L=128$) 85.51 84.26 79.15 90.07
InterCLIP-MEP w/ V2T ($L=640$) 86.26 85.00 80.17 90.45
InterCLIP-MEP w/ T2V ($L=1024$) 86.72 85.61 80.20 91.80
Table 2: Main results. We use * to indicate that the results are taken from Qin et al. (2023). $\dagger$ indicates the results obtained by running this method on the MMSD2.0 benchmark (Qin et al. 2023).
Experiment
Experimental Settings
Datasets.
We use the dataset from the MMSD2.0 benchmark (Qin et al. 2023), which is an improved version of the dataset originally introduced by Cai, Cai, and Wan (2019). We present the statistics of the dataset in Table 1.
Metrics.
Following Liu, Wang, and Li (2022), we use accuracy (Acc.), precision (P), recall (R), and F1-score (F1) as metrics to evaluate the performance of our model.
Baselines.
We compare the effectiveness of the InterCLIP-MEP framework against several uni-modal and multi-modal methods. For text modality methods, we compare with TextCNN (Kim 2014), Bi-LSTM (Graves and Schmidhuber 2005), SMSD (Xiong et al. 2019), and RoBERTa (Liu et al. 2019). For image modality methods, we compare with ResNet (He et al. 2015) and ViT (Dosovitskiy et al. 2020). For state-of-the-art multi-modal methods, we compare with the following:
  • HFM (Cai, Cai, and Wan 2019): A hierarchical fusion model integrating text and image features.
  • Att-BERT (Pan et al. 2020): A model using an attention mechanism to capture sarcasm incongruity.
  • CMGCN (Liang et al. 2022): A method that constructs a cross-modal graph and employs a graph convolutional network.
  • HKE (Liu, Wang, and Li 2022): A hierarchical framework utilizing multi-head cross-attention and graph neural networks.
  • DIP (Wen, Jia, and Yang 2023): A dual incongruity perceiving network that extracts sarcastic information from factual and affective levels.
  • DynRT (Tian et al. 2023): A dynamic routing transformer network for multi-modal sarcasm detection.
  • Multi-view CLIP (Qin et al. 2023): A method adapted to CLIP for capturing multi-view sarcasm cues.
  • G2SAM (Wei et al. 2024): A method using global graph-based semantic awareness and leveraging historical knowledge from training samples for multi-modal sarcasm detection.
Param. Value
$r$ 8
top-$n$ 4
$d_{f}$ 1024
$\mathbf{W}$ $W_{k},W_{v},W_{o}$
Epoch 3
Batch size 64
Learning rate 5e-4
LoRA learning rate 1e-4
Table 3: Hyperparameter settings.
Implementation Details
We implement model training and testing using PyTorch Lightning [^1] https://lightning.ai/. We construct InterCLIP leveraging the Transformers library (Wolf et al. 2020), specifically utilizing the clip-vit-base-patch32222https://huggingface.co/openai/clip-vit-base-patch32 model from the HuggingFace Hub [^3] https://huggingface.co/models. We optimize model parameters using AdamW (Loshchilov and Hutter 2017), setting the learning rate to 1e-4 for the LoRA fine-tuning modules and 5e-4 for other trainable modules. We employ a cosine annealing scheduler with warmup to dynamically adjust the learning rate, with warmup steps constituting the first 20% of the total optimization steps and the minimum learning rate set to 1% of the initial rate. For the modules $\mathcal{G}_{t/v}$, $\mathcal{F}_{t/v}$, $\mathcal{F}_{c}$, and $\mathcal{F}_{p}$, we use simple multi-layer perceptrons (MLPs). We perform training with a batch size of 64 for 3 epochs. We run all experiments on a machine with an NVIDIA RTX 4090 GPU.
Main Results
The main results are shown in Table 2. To validate the effectiveness of our InterCLIP-MEP framework, we conduct experiments using the original CLIP as the backbone instead of InterCLIP, referred to as InterCLIP-MEP w/o Inter. We compare this configuration with three interaction modes of InterCLIP: InterCLIP-MEP w/ V2T, InterCLIP-MEP w/ T2V, and InterCLIP-MEP w/ TW.
For each experiment, we condition only the top four layers of the self-attention modules, with the projection dimension $d_{f}$ set to 1024. We set the rank $r$ of LoRA to 8, fine-tuning the self-attention module weight matrices $\mathbf{W}$, specifically $W_{k}$, $W_{v}$, and $W_{o}$. For the memory size $L$ of the MEP, we select the optimal size from $\mathbf{L}=\{128,256,384,512,640,768,896,1024,1152,1280\}$. Relevant hyperparameters are summarized in Table 3.
Our framework consistently outperforms or matches the performance of state-of-the-art methods, whether using InterCLIP or the original CLIP as the backbone, as shown in Table 2. This demonstrates the effectiveness of our training strategy and MEP. Methods InterCLIP-MEP w/ V2T and InterCLIP-MEP w/ T2V outperform InterCLIP-MEP w/o Inter, indicating that InterCLIP more effectively captures text-image interactions. Additionally, InterCLIP-MEP w/ T2V outperforms InterCLIP-MEP w/ V2T, likely due to the complexity of the visual space, which poses challenges for the projection layer in mapping visual representations into the text encoder space effectively. We observe that InterCLIP-MEP w/ TW performs worse than InterCLIP-MEP w/o Inter, possibly because embedding representations within both encoders increases learning difficulty compared to unidirectional methods. In summary, InterCLIP with T2V interaction, combined with our training strategy and MEP yields the most promising results.
Method Acc. (%) F1 (%)
InterCLIP-MEP w/o Inter
Baseline 86.05 84.81
w/o Proj 85.76 84.43
w/o MEP 85.39 83.99
w/o LoRA 82.44 77.73
InterCLIP-MEP w/ TW
Baseline 85.51 84.26
w/o Proj 85.43 84.05
w/o MEP 85.22 83.79
w/o LoRA 76.42 74.37
InterCLIP-MEP w/ V2T
Baseline 86.26 85.00
w/o Proj 85.68 84.22
w/o MEP 86.26 84.78
w/o LoRA 73.31 72.22
InterCLIP-MEP w/ T2V
Baseline 86.72 85.61
w/o Proj 86.22 84.51
w/o MEP 86.26 84.82
w/o LoRA 75.13 71.79
Table 4: Ablation study results of InterCLIP-MEP. Baseline represents the results of the corresponding method without any ablation.
Ablation Study
To further validate the effectiveness of InterCLIP-MEP, we remove the projection module $\mathcal{F}_{p}$ and train only the classification module $\mathcal{F}_{c}$ for prediction, denoted as w/o Proj. To test the necessity of using LoRA (Hu et al. 2022) for fine-tuning, we keep the rest of InterCLIP-MEP unchanged and freeze all self-attention weight matrices of InterCLIP, denoted as w/o LoRA, always selecting the optimal memory size $L$ for the MEP during inference. To evaluate the effectiveness of the MEP, we train both $\mathcal{F}_{p}$ and $\mathcal{F}_{c}$ but use only $\mathcal{F}_{c}$ during inference, denoted as w/o MEP.
Table 4 reports all results. All variants show performance declines compared to the baseline, demonstrating the importance of each module in the InterCLIP-MEP framework. For methods InterCLIP-MEP w/ TW and InterCLIP-MEP w/o Inter, the w/o MEP variant performed worse than the w/o Proj variant. However, for methods InterCLIP-MEP w/ T2V and InterCLIP-MEP w/ V2T, the w/o MEP variant performs better than the w/o Proj variant. This suggests that backbones with strong image-text interaction capabilities benefit from training the classification module $\mathcal{F}_{c}$ along with the projection module $\mathcal{F}_{p}$, even without using MEP during inference. We also find that not using LoRA to fine-tune the self-attention modules results in significant performance loss, indicating that the original CLIP’s vanilla space is not directly suitable for the sarcasm detection task.
Hyperparameter Study
In this section, we investigate the impact of different hyperparameter settings on model performance. Keeping the other hyperparameters constant as shown in Table 3, we fine-tune all possible weight matrices $\mathbf{W}$ for methods using original CLIP and different interaction modes of Interactive-CLIP as the backbone. We consistently select the optimal memory size from $\mathbf{L}$ for MEP. We calculate the average metrics for different methods and different weight matrices.
$\mathbf{W}$ Mean Acc. (%) Mean F1 (%)
$W_{q}$ 85.14 83.99
$W_{k}$ 85.09 84.00
$W_{v}$ 85.34 84.24
$W_{o}$ 85.39 84.23
$W_{q},W_{k}$ 85.36 84.16
$W_{q},W_{v}$ 85.63 84.40
$W_{q},W_{o}$ 85.73 84.49
$W_{k},W_{o}$ 85.67 84.43
$W_{v},W_{o}$ 85.73 84.54
$W_{k},W_{v}$ 85.70 84.51
$W_{q},W_{k},W_{o}$ 85.92 84.63
$W_{q},W_{v},W_{o}$ 85.87 84.60
$W_{q},W_{k},W_{v}$ 86.05 84.75
$W_{k},W_{v},W_{o}$ 86.14 84.92
$W_{q},W_{k},W_{v},W_{o}$ 85.82 84.55
Table 5: Average results of fine-tuning different weight matrices $\mathbf{W}$ across four baseline methods.
Method Mean Acc. (%) Mean F1 (%)
w/o Inter 85.49 84.32
w/ TW 85.66 84.44
w/ V2T 85.62 84.41
w/ T2V 85.78 84.55
Table 6: Average results of four baseline methods for fine-tuning different weight matrices $\mathbf{W}$.
The results are presented in Table 5 and Table 6. We observe that fine-tuning the weight matrices $W_{k},W_{v},W_{o}$ and using the T2V interaction mode of Interactive-CLIP are the best choices for our InterCLIP-MEP framework.
We further investigate the method using Interactive-CLIP with T2V interaction as the backbone. Keeping the other hyperparameters constant, we condition different top-n layers of the self-attention modules. We also study the impact of different projection dimensions $d_{f}$, different LoRA ranks $r$, and different memory sizes $L$ on the InterCLIP-MEP w/ T2V method. We present all results in Figure 4.
Figure 4: Hyperparameter study curves for InterCLIP-MEP w/ T2V. Figure 4(d) compares results with those from using only the classification module $\mathcal{F}_{c}$ for prediction.
Figure 4(a) shows that conditioning the top four self-attention layers yields the best results. From Figure 4(b), a rank of 8 is optimal. Figure 4(c) indicates the projection dimension is best at 256 or 1024. Figure 4(d) reveals that a memory size of 640 in MEP outperforms others, confirming the value of historical sample knowledge.
Figure 5: Case study of InterCLIP-MEP.
Figure 6: Visual comparison between Interactive-CLIP and original CLIP.
Case Study
As shown in Figure 5, we select three examples to further demonstrate the robustness of InterCLIP-MEP. We observe that direct predictions through the classification module $\mathcal{F}_{c}$ result in high entropy and incorrect outcomes. However, because the InterCLIP-MEP framework leverages the Memory-Enhanced Predictor (MEP) to utilize historical knowledge of test samples during inference, This approach effectively mitigates the issue for cases where the classification module fails to correctly identify the results by using the historical knowledge of the test samples.
Visualization
To further validate that Interactive-CLIP can more effectively capture the interactive information between text and images compared to original CLIP, thereby aiding in the detection of sarcasm cues, we visualize the areas of focus during the inference process of the visual model in Figure 6. We observe that in the first example, which complains about the unpleasant odor caused by the sea lions, Interactive-CLIP focuses more accurately on the location of the sea lions compared to the original CLIP. In the second example, which complains about traffic congestion, Interactive-CLIP correctly focuses on the distribution of cars on the road, whereas the original CLIP’s focus is scattered.
Conclusion and Future Work
In this paper, we propose a novel framework for robust multi-modal sarcasm detection, termed InterCLIP-MEP. We introduce Interactive CLIP (InterCLIP) as the backbone for extracting sample representations. Unlike the original CLIP, InterCLIP embeds representations from different modalities into the encoder, effectively capturing the intricate interactions between text and image. For inference, we introduce a Memory-Enhanced Predictor (MEP) that leverages historical knowledge of test samples to enhance inference, providing a more robust alternative to directly using the classification module for sarcasm recognition. This approach significantly improves the detection of sarcastic samples. Our experimental results demonstrate that InterCLIP-MEP achieves state-of-the-art performance on the MMSD2.0 benchmark. In future work, we plan to extend our approach to sarcasm recognition tasks across additional modalities, such as speech and video.
Acknowledgments
The research described in the paper received funding from the Anhui Provincial Natural Science Foundation through Grant No. 2308085MF220. Additionally, it was co-supported by the Anhui University Natural Science Foundation, which provided funding through Grant Nos. 2022AH050972 and KJ2021A0516.

### v3

@@
 The prevalence of sarcasm in social media, conveyed through text-image combinations, presents significant challenges for sentiment analysis and intention mining.
-Current multi-modal sarcasm detection methods have been proven to struggle with biases from spurious cues, leading to a superficial understanding of the complex interactions between text and image.
-To address these issues, we propose InterCLIP-MEP, a robust framework for multi-modal sarcasm detection.
-InterCLIP-MEP introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone, enhancing sample representations by embedding cross-modality information in each encoder.
-Furthermore, a novel training strategy is designed to adapt InterCLIP for a Memory-Enhanced Predictor (MEP).
-MEP uses dynamic dual-channel memory to store valuable historical knowledge of test samples and then leverages this memory as a non-parametric classifier to derive the final prediction.
-By using InterCLIP to encode text-image interactions more effectively and incorporating MEP, InterCLIP-MEP offers a more robust recognition of multi-modal sarcasm.
-Experiments demonstrate that InterCLIP-MEP achieves state-of-the-art performance on the MMSD2.0 benchmark.
-Code and data are available at https://github.com/CoderChen01/InterCLIP-MEP.
+Existing multi-modal sarcasm detection methods have been proven to overestimate performance, as they struggle to effectively capture the intricate sarcastic cues that arise from the interaction between an image and text.
+To address these issues, we propose InterCLIP-MEP, a novel framework for multi-modal sarcasm detection.
+Specifically, we introduce an Interactive CLIP (InterCLIP) as the backbone to extract text-image representations, enhancing them by embedding cross-modality information directly within each encoder, thereby improving the representations to capture text-image interactions better.
+Furthermore, an efficient training strategy is designed to adapt InterCLIP for our proposed Memory-Enhanced Predictor (MEP).
+MEP uses a dynamic, fixed-length dual-channel memory to store historical knowledge of valuable test samples during inference.
+It then leverages this memory as a non-parametric classifier to derive the final prediction, offering a more robust recognition of multi-modal sarcasm.
+Experiments demonstrate that InterCLIP-MEP achieves state-of-the-art performance on the MMSD2.0 benchmark, with an accuracy improvement of 1.08% and an F1 score improvement of 1.51% over the previous best method.
 Introduction
 Sarcasm, with its inherent subtlety and complexity, plays a significant role in human communication by often conveying irony, mockery, or hidden intentions (Muecke 1982; Gibbs and O’Brien 1991; Gibbs and Colston 2007).
+Figure 1: Comparative analysis of multi-modal sarcasm detection pipelines.
+Panels (a), (b), and (c) depict the pipeline comparisons between our InterCLIP-MEP method and other existing methods.
+Panel (d) utilizes GradCAM to visually demonstrate the superior ability of the InterCLIP backbone over the traditional CLIP backbone in identifying multi-modal sarcasm cues, which are marked with red boxes.
 The automatic detection of sarcasm from text has been a significant research focus, aiding tasks such as sentiment analysis and intent mining (Pang, Lee et al.
-2008; Tsur, Davidov, and Rappoport 2010; Bouazizi and Ohtsuki 2015; Poria et al.
-2016; Amir et al.
-2016; Wu et al.
-2018; Tay et al.
-2018; Baziotis et al.
-2018).
-With the rapid development of social media platforms like Twitter and Reddit, the need for multi-modal sarcasm detection has become increasingly apparent.
-Users frequently use text-image combinations to convey their messages, making it essential to develop methods to interpret these multi-modal communications.
-The task of multi-modal sarcasm detection presents significant challenges, requiring the system to accurately capture and comprehend the intricate interplay between textual and visual information to identify sarcasm cues.
-A pivotal development in this field was the introduction of the Multi-Modal Sarcasm Detection (MMSD) benchmark by Cai, Cai, and Wan (2019).
-This benchmark has become a cornerstone for subsequent research, inspiring a range of advanced techniques aimed at improving sarcasm detection performance (Xu, Zeng, and Mao 2020; Pan et al.
+2008; Tsur, Davidov, and Rappoport 2010; Bouazizi and Ohtsuki 2015).
+With the rapid development of social media platforms like Twitter and Reddit, users frequently use text-image combinations to convey their messages.
+Therefore, the need for multi-modal sarcasm detection has become increasingly apparent, which presents significant challenges, requiring the system to capture and comprehend the intricate interplay between textual and visual information to identify sarcasm cues.
+Despite the notable advancements made by existing studies (Xu, Zeng, and Mao 2020; Pan et al.
 2020; Liang et al.
-2021, 2022; Liu, Wang, and Li 2022; Wen, Jia, and Yang 2023; Tian et al.
-2023; Wei et al.
-2024).
-The above methods have explored various approaches, including the use of Transformer-based models for high-quality representation extraction (Pan et al.
-2020), OCR and object detection technologies to enhance image understanding (Qiao et al.
-2023), and the incorporation of external emotional knowledge (Liu, Wang, and Li 2022).
-Graph neural networks have also been utilized to construct dependency and cross-modal graphs for improved sarcasm identification (Liu, Wang, and Li 2022; Liang et al.
-2021, 2022).
-Figure 1: Comparison of InterCLIP-MEP and other multi-modal sarcasm detection pipelines.
-Despite these advancements, the existing methods have shown limitations, often learning biases from spurious cues present in the MMSD benchmark (Qin et al.
-2023).
-To address this, Qin et al.
-(2023) introduced MMSD2.0, a refined benchmark that eliminates these spurious cues and corrects mislabeled samples.
-Their work demonstrated the need for more reliable benchmarks and robust detection methods.
-Building on these insights, we propose InterCLIP-MEP, a novel framework designed to enhance the reliability of multi-modal sarcasm detection.
-We present a comparison of existing methods with InterCLIP-MEP in Figure 1.
-As shown in Figure 1(a) and Figure 1(b), some existing methods use separate uni-modal encoders, such as BERT (Devlin et al.
-2019) for text and ViT (Dosovitskiy et al.
-2020) for visual modalities, as backbones (Xu, Zeng, and Mao 2020; Pan et al.
+2021, 2022) in this field, Qin et al.
+(2023) have highlighted that the MMSD benchmark (Cai, Cai, and Wan 2019) employed in these works contains spurious cues.
+These cues lead to the development of biased models, thereby inflating the perceived effectiveness of these models in detecting sarcasm cues from multi-modal data.
+By introducing MMSD2.0, a refined benchmark that eliminates spurious cues and corrects mislabeled samples, Qin et al.
+(2023) demonstrates a substantial performance decline of existing methods when evaluated on MMSD2.0.
+Although they proposed a promising method that employs a Multi-view CLIP for reliable multi-modal sarcasm detection, it still leaves room for improvement, similar to their previous approaches.
+As shown in Figure 1(a) and Figure 1(b), some existing methods (Xu, Zeng, and Mao 2020; Pan et al.
 2020; Liang et al.
@@
 2023; Wei et al.
-2024).
-These encoded samples are fused using various techniques to obtain the final sample representations.
-We argue that this approach makes it difficult for the model to capture the subtle interactive information between text and images.
+2024) use separate uni-modal encoders, such as BERT (Devlin et al.
+2019) for text and ViT (Dosovitskiy et al.
+2020) for image, as the backbone to encode samples.
+This approach makes it difficult for the model to capture the subtle interactive information between text and image.
 Although Qin et al.
@@
 2021), to encode samples and demonstrated notable results on the MMSD2.0 benchmark, the original CLIP only aligns text and images directly, lacking the nuanced semantic space required for sarcasm detection.
-Therefore, InterCLIP-MEP introduces a refined variant, Interactive CLIP (InterCLIP), as the backbone to enhance sample encoding for this specific task.
-Inspired by the foundational work of Ganz et al.
-(2024), InterCLIP enhances the capture of interactive information between text and image by embedding representations from one modality into the encoder of the other, leading to a more nuanced understanding of multi-modal sarcasm cues.
-After encoding the samples, as depicted in Figure 1(a), many approaches directly identify sarcastic samples through a classification module.
-This often leads to high prediction entropy and incorrect results, especially for hard cases with greater uncertainty.
-To the best of our knowledge, few methods consider leveraging historical knowledge of samples to aid in detection.
-As shown in Figure 1(b), only Wei et al.
-(2024) constructed a static semantic space that stores representations of all training samples.
+Furthermore, as depicted in Figure 1(a), some approaches (Xu, Zeng, and Mao 2020; Pan et al.
+2020; Liang et al.
+2021, 2022; Wen, Jia, and Yang 2023; Tian et al.
+2023) directly classify sarcastic samples based on their representations.
+This often leads to high prediction entropy and erroneous outcomes, especially for hard cases with greater uncertainty.
+To the best of our knowledge, only Wei et al.
+(2024) considers leveraging historical knowledge of samples to aid in detection.
+As shown in Figure 1(b), they constructed a static semantic space that stores features of all training samples.
 When identifying sarcastic samples, they query $k$-nearest neighbors in this semantic space and use a voting mechanism to derive the final predictions.
-However, their static semantic space struggles to handle out-of-distribution issues.
-As shown in Figure 1(c), InterCLIP-MEP uses a novel training strategy to adapt InterCLIP for a Memory-Enhanced Predictor (MEP).
-Based on the sample representations encoded by InterCLIP, InterCLIP-MEP trains a classification module to assign pseudo-labels to samples.
-Simultaneously, it trains a projection module to map samples into a latent space.
+However, relying heavily on a static semantic space may limit its ability to adapt to various test samples.
+In this paper, building on these insights, we propose InterCLIP-MEP, a novel framework designed to enhance the reliability of multi-modal sarcasm detection.
+As shown in Figure 1(c), our InterCLIP-MEP introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone to enhance sample encoding for detecting multi-modal sarcasm.
+InterCLIP enhances the capture of interactive information between text and image by embedding representations from one modality into the encoder of the other, leading to a more nuanced understanding of multi-modal sarcasm cues.
+In Figure 1(d), we utilized GradCAM (Selvaraju et al.
+2016) visualization to compare the effectiveness of our InterCLIP backbone against the CLIP backbone in detecting multi-modal sarcasm cues.
+During inference, InterCLIP-MEP introduces a Memory-Enhanced Predictor (MEP) to build dynamically a non-parametric classifier from valuable historical features to identify multi-modal sarcasm.
+Specifically, we design an efficient training strategy to adapt InterCLIP for MEP.
+Based on the sample representations encoded by InterCLIP, we train a classification module to assign pseudo-labels to samples.
+Simultaneously, we train a projection module to map samples into a latent space.
 In this latent space, same-class samples are pulled closer together, while different-class samples are pushed further apart.
-Additionally, InterCLIP-MEP fine-tunes the weights of the self-attention modules in InterCLIP’s text and vision encoders using LoRA (Hu et al.
+Additionally, we fine-tune the weights of the self-attention modules in InterCLIP’s text and vision encoders using the low-rank adaptation (LoRA) technique (Hu et al.
 2022), building upon the original CLIP weights.
-Building on InterCLIP and its training strategy, MEP leverages historical knowledge of test samples to identify sarcasm.
+This approach allows for more efficient adaptation without significantly increasing the computational burden.
+Building on InterCLIP and the training strategy, MEP leverages historical knowledge of valuable test samples to identify multi-modal sarcasm.
 It maintains a dynamic, fixed-length dual-channel memory, with one channel storing historical information of non-sarcastic samples and the other storing information of sarcastic samples.
-MEP uses the learned InterCLIP to extract sample representations, calculate pseudo-labels with the classification module, and compute prediction entropy, while also obtaining projection features through the projection module.
-Based on the pseudo-labels and prediction entropy, MEP dynamically updates the projected features of samples into the corresponding memory channels, treating these features as the historical knowledge of the test samples.
-Specifically, MEP consistently updates the memory with projected features of samples that have lower prediction entropy.
-MEP utilizes this memory as a non-parametric classifier to compute the final prediction for each sample.
-Our experiments demonstrate the robustness and innovation of InterCLIP-MEP in accurately identifying sarcasm through sophisticated interaction and memory mechanisms.
-Our contributions can be summarized as follows:
-  • We propose InterCLIP-MEP, a novel framework for reliable multi-modal sarcasm detection, which introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone, improving the encoding of text-image interactions by embedding cross-modality information within each encoder.
-  • We design a novel training strategy to adapt InterCLIP for a Memory-Enhanced Predictor (MEP).
-This strategy enables the integration of dynamic dual-channel memory, which stores valuable historical knowledge of test samples and leverages this memory for more precise and informed predictions.
-  • Our extensive experiments on the more reliable benchmark, MMSD2.0, demonstrate that InterCLIP-MEP outperforms existing state-of-the-art methods.
+MEP then utilizes this memory as a non-parametric classifier to compute the final prediction for each sample, offering more robust and reliable multi-modal sarcasm detection.
+Overall, our contributions can be summarized as follows:
+  • We propose InterCLIP-MEP, a novel framework for reliable multi-modal sarcasm detection, which introduces Interactive CLIP (InterCLIP) improving the encoding of text-image interactions by embedding cross-modality information within each encoder and Memory-Enhanced Predictor (MEP) offering more robust and reliable multi-modal sarcasm detection.
+  • We develop an efficient training strategy that ensures computational efficiency and superior performance, bridging InterCLIP and MEP.
+  • Our extensive experiments on the more reliable benchmark, MMSD2.0, demonstrate that InterCLIP-MEP [^1] https://github.com/CoderChen01/InterCLIP-MEP outperforms existing state-of-the-art methods.
 Figure 2: Overview of our framework.
@@
 Related Work
-Early work in sarcasm detection primarily focused on textual data (Bouazizi and Ohtsuki 2015; Poria et al.
-2016; Amir et al.
-2016; Wu et al.
-2018; Tay et al.
-2018; Baziotis et al.
+Multi-modal sarcasm detection.
+Early work in sarcasm detection primarily focused on textual data (Bouazizi and Ohtsuki 2015; Amir et al.
+2016; Baziotis et al.
 2018).
-With the rise of social media, detecting sarcasm from text-image data has become increasingly challenging, leading to the development of several multi-modal approaches (Schifanella et al.
-2016; Cai, Cai, and Wan 2019; Xu, Zeng, and Mao 2020; Pan et al.
-2020; Liang et al.
-2021, 2022; Liu, Wang, and Li 2022; Qin et al.
-2023; Wen, Jia, and Yang 2023; Tian et al.
-2023; Wei et al.
-2024).
+With the rise of social media, detecting sarcasm from text-image data has become increasingly challenging, leading to the development of several multi-modal approaches.
 Schifanella et al.
@@
 2024).
-Further research expanded on MMSD, exploring dependencies and external knowledge, using advanced models, and introducing graph-based methods to enhance performance (Xu, Zeng, and Mao 2020; Pan et al.
-2020; Liang et al.
-2021, 2022; Liu, Wang, and Li 2022).
 However, MMSD has been proven to contain spurious cues that could bias models (Qin et al.
@@
 To address this, Qin et al.
-(2023) introduced MMSD2.0 benchmark, which removed these cues and corrected mislabeled samples.
+(2023) introduced the MMSD2.0 benchmark, which removed these cues and corrected mislabeled samples.
 They observed a significant performance drop when re-evaluating state-of-the-art methods on MMSD2.0 and proposed the need for more stable multi-modal sarcasm detection methods.
-Motivated by the work of Qin et al.
-(2023), we introduce the InterCLIP-MEP framework for more reliable multi-modal sarcasm detection.
+In this work, we introduce the InterCLIP-MEP framework towards more reliable multi-modal sarcasm detection.
+CLIP adaptation.
+The Contrastive Language-Image Pretraining (CLIP) model (Radford et al.
+2021) excels in vision-language tasks.
+Adapting CLIP for specific domains has shown substantial improvements, as demonstrated by Li, Shakhnarovich, and Yeh (2022) for phrase localization, Liang et al.
+(2023) for open-vocabulary semantic segmentation, and Wang et al.
+(2023) for action recognition.
+In this work, inspired by Ganz et al.
+(2024), we conditionally enhance both the text and visual encoders of CLIP, making it more effective in capturing the interplay between text and images to identify sarcasm.
+Unlike Ganz et al.
+(2024), who focused solely on embedding text into the visual encoder, we also explore embedding images into the text encoder.
+Furthermore, their approach is limited to general classification tasks and does not address the complexities of multi-modal sarcasm detection.
+Mermory-enhanced prediction.
+Inspired by cognitive science (Stokes 2015; Baddeley 2000), memory has been introduced to enhance neural networks (Weston, Chopra, and Bordes 2014; Sukhbaatar et al.
+2015).
+Several studies (Wu et al.
+2018; Wen, Jia, and Yang 2023) have used memory mechanisms to improve model training, and some (Zhang et al.
+2024; Wei et al.
+2024) leverage memory to store historical knowledge, enhancing prediction accuracy.
+In this work, we introduce a memory-enhanced predictor for multi-modal sarcasm detection.
+In contrast to other methods, our memory dynamically updates during testing, utilizing relevant historical information for improved accuracy and robustness.
 Methodology
-In this paper, we introduce InterCLIP-MEP, a novel framework for multi-modal sarcasm detection that combines interactive cross-modal encoding with memory-enhanced prediction.
 An overview of InterCLIP-MEP is illustrated in Figure 2.
-Our approach leverages a variant of CLIP, called Interactive CLIP (InterCLIP), which enhances the original CLIP architecture by embedding representations from one modality into the encoder of the other modality, resulting in more interactive representations.
-InterCLIP enables a nuanced understanding of the interaction between text and image.
-To facilitate memory-enhanced prediction, we design a novel training strategy, integrating a classification module and a projection module, with InterCLIP serving as the backbone.
-These modules are trained to identify sarcastic samples while constructing a latent space where features of the same class are clustered together and those of different classes are distinctly separated.
-For the prediction phase, InterCLIP-MEP employs a Memory-Enhanced Predictor (MEP) that leverages historical knowledge for final predictions.
-Instead of relying solely on the classification module to identify sarcasm, MEP compares the features of the current sample with those stored in memory, thereby enhancing the accuracy of sarcasm detection.
+Initially, we elaborate on the Interactive CLIP (InterCLIP) and its training strategy, followed by an in-depth explanation of the Memory-Enhanced Predictor (MEP).
 Interactive CLIP
-We introduce Interactive CLIP (InterCLIP), which can condition text or image encoder to interact with representations from the other modality, enhancing the capture of interactive information crucial for identifying sarcasm.
-The input to InterCLIP is a text-image pair $\mathcal{P}=(T,I)$, where $T$ represents a piece of text and $I$ represents an image.
+The input to Interactive CLIP (InterCLIP) is a text-image pair $\mathcal{P}=(T,I)$, where $T$ represents a piece of text and $I$ represents an image.
 Here, for simplicity, we do not consider the case of batch inputs.
@@
 We use $\tilde{h}^{t}_{i}(\cdot)\in\mathbb{R}^{d_{t}}$ and $\tilde{h}^{v}_{i}(\cdot)\in\mathbb{R}^{d_{v}}$ to denote the re-encoded interactive representations of each text token and image patch, respectively.
-Specifically, we condition only the top-$n$ self-attention layers of the text or vision encoder, where $n$ is a hyperparameter that will be analyzed in subsequent sections.
+In particular, we condition only the top-$n$ self-attention layers of the text or vision encoder, where $n$ is a hyperparameter that will be analyzed in the experiment section.
 The structure of the conditioned self-attention layers is illustrated in Figure 3.
@@
 Training Strategy
-As shown in Figure 2 (left), to adapt InterCLIP for memory-enhanced prediction, we introduce a novel training strategy.
+As shown in Figure 2 (left), to adapt InterCLIP for MEP, we introduce an efficient training strategy.
 Using InterCLIP as the backbone to obtain fused features of the samples, we introduce a classification module and a projection module.
@@
 where $\textrm{norm}(\cdot)$ denotes L2 normalization, and $d_{f}$ represents the dimension of the projected features.
-In this space, cosine distance between features of the same class is minimized, while the distance between features of different classes is maximized.
+In this space, the cosine distance between features of the same class is minimized, while the distance between features of different classes is maximized.
 We use a label-aware cosine similarity loss to optimize $\mathcal{F}_{p}$:
@@
 24:  end for
-As depicted in Figure 2 (right), we present a Memory-Enhanced Predictor (MEP) that builds upon the learned InterCLIP, along with the classification module and the projection module, leveraging the valuable historical knowledge of test samples to enhance sarcasm detection.
+As depicted in Figure 2 (right), we present the Memory-Enhanced Predictor (MEP) that builds upon the learned InterCLIP, along with the classification module and the projection module, leveraging the valuable historical knowledge of test samples to enhance sarcasm detection.
 The detailed computational process of MEP is provided in Algorithm 1, where $N_{\textrm{test}}$ denotes the number of test samples.
@@
 All Sarcastic Non-sarcastic
-Train 19812 9572 10240
-Validation 2410 1042 1368
-Test 2409 1037 1372
+Train 19,812 9,572 10,240
+Validation 2,410 1,042 1,368
+Test 2,409 1,037 1,372
 Table 1: Statistics of the dataset from the MMSD2.0 benchmark (Qin et al. 2023).
@@
 2023).
+Underlined values represent the best baseline for comparison.
+Bold values indicate those that surpass the baseline.
+w/o Inter w/ TW w/ V2T w/ T2V
+Variant Acc. (%) F1 (%) Acc. (%) F1 (%) Acc. (%) F1 (%) Acc. (%) F1 (%)
+Baseline 86.05 84.81 85.51 84.26 86.26 85.00 86.72 85.61
+w/o Proj 85.76 84.43 85.43 84.05 85.68 84.22 86.22 84.51
+w/o MEP 85.39 83.99 85.22 83.79 86.26 84.78 86.26 84.82
+w/o LoRA 82.44 77.73 76.42 74.37 73.31 72.22 75.13 71.79
+Table 3: Ablation study of InterCLIP-MEP. Baseline represents the results of the corresponding method without any ablation.
 Experiment
@@
   • G2SAM (Wei et al. 2024): A method using global graph-based semantic awareness and leveraging historical knowledge from training samples for multi-modal sarcasm detection.
-Param. Value
-$r$ 8
-top-$n$ 4
-$d_{f}$ 1024
-$\mathbf{W}$ $W_{k},W_{v},W_{o}$
-Epoch 3
-Batch size 64
-Learning rate 5e-4
-LoRA learning rate 1e-4
-Table 3: Hyperparameter settings.
-Implementation Details
-We implement model training and testing using PyTorch Lightning [^1] https://lightning.ai/.
-We construct InterCLIP leveraging the Transformers library (Wolf et al.
-2020), specifically utilizing the clip-vit-base-patch32222https://huggingface.co/openai/clip-vit-base-patch32 model from the HuggingFace Hub [^3] https://huggingface.co/models.
-We optimize model parameters using AdamW (Loshchilov and Hutter 2017), setting the learning rate to 1e-4 for the LoRA fine-tuning modules and 5e-4 for other trainable modules.
-We employ a cosine annealing scheduler with warmup to dynamically adjust the learning rate, with warmup steps constituting the first 20% of the total optimization steps and the minimum learning rate set to 1% of the initial rate.
-For the modules $\mathcal{G}_{t/v}$, $\mathcal{F}_{t/v}$, $\mathcal{F}_{c}$, and $\mathcal{F}_{p}$, we use simple multi-layer perceptrons (MLPs).
-We perform training with a batch size of 64 for 3 epochs.
-We run all experiments on a machine with an NVIDIA RTX 4090 GPU.
 Main Results
-The main results are shown in Table 2.
 To validate the effectiveness of our InterCLIP-MEP framework, we conduct experiments using the original CLIP as the backbone instead of InterCLIP, referred to as InterCLIP-MEP w/o Inter.
@@
 For the memory size $L$ of the MEP, we select the optimal size from $\mathbf{L}=\{128,256,384,512,640,768,896,1024,1152,1280\}$.
-Relevant hyperparameters are summarized in Table 3.
+The main results are shown in Table 2.
 Our framework consistently outperforms or matches the performance of state-of-the-art methods, whether using InterCLIP or the original CLIP as the backbone, as shown in Table 2.
@@
 In summary, InterCLIP with T2V interaction, combined with our training strategy and MEP yields the most promising results.
-Method Acc. (%) F1 (%)
-InterCLIP-MEP w/o Inter
-Baseline 86.05 84.81
-w/o Proj 85.76 84.43
-w/o MEP 85.39 83.99
-w/o LoRA 82.44 77.73
-InterCLIP-MEP w/ TW
-Baseline 85.51 84.26
-w/o Proj 85.43 84.05
-w/o MEP 85.22 83.79
-w/o LoRA 76.42 74.37
-InterCLIP-MEP w/ V2T
-Baseline 86.26 85.00
-w/o Proj 85.68 84.22
-w/o MEP 86.26 84.78
-w/o LoRA 73.31 72.22
-InterCLIP-MEP w/ T2V
-Baseline 86.72 85.61
-w/o Proj 86.22 84.51
-w/o MEP 86.26 84.82
-w/o LoRA 75.13 71.79
-Table 4: Ablation study results of InterCLIP-MEP. Baseline represents the results of the corresponding method without any ablation.
 Ablation Study
@@
 Hyperparameter Study
-In this section, we investigate the impact of different hyperparameter settings on model performance.
-Keeping the other hyperparameters constant as shown in Table 3, we fine-tune all possible weight matrices $\mathbf{W}$ for methods using original CLIP and different interaction modes of Interactive-CLIP as the backbone.
-We consistently select the optimal memory size from $\mathbf{L}$ for MEP.
-We calculate the average metrics for different methods and different weight matrices.
-$\mathbf{W}$ Mean Acc. (%) Mean F1 (%)
-$W_{q}$ 85.14 83.99
-$W_{k}$ 85.09 84.00
-$W_{v}$ 85.34 84.24
-$W_{o}$ 85.39 84.23
-$W_{q},W_{k}$ 85.36 84.16
-$W_{q},W_{v}$ 85.63 84.40
-$W_{q},W_{o}$ 85.73 84.49
-$W_{k},W_{o}$ 85.67 84.43
-$W_{v},W_{o}$ 85.73 84.54
-$W_{k},W_{v}$ 85.70 84.51
-$W_{q},W_{k},W_{o}$ 85.92 84.63
-$W_{q},W_{v},W_{o}$ 85.87 84.60
-$W_{q},W_{k},W_{v}$ 86.05 84.75
-$W_{k},W_{v},W_{o}$ 86.14 84.92
-$W_{q},W_{k},W_{v},W_{o}$ 85.82 84.55
-Table 5: Average results of fine-tuning different weight matrices $\mathbf{W}$ across four baseline methods.
-Method Mean Acc. (%) Mean F1 (%)
-w/o Inter 85.49 84.32
-w/ TW 85.66 84.44
-w/ V2T 85.62 84.41
-w/ T2V 85.78 84.55
-Table 6: Average results of four baseline methods for fine-tuning different weight matrices $\mathbf{W}$.
-The results are presented in Table 5 and Table 6.
-We observe that fine-tuning the weight matrices $W_{k},W_{v},W_{o}$ and using the T2V interaction mode of Interactive-CLIP are the best choices for our InterCLIP-MEP framework.
-We further investigate the method using Interactive-CLIP with T2V interaction as the backbone.
+In this section, we further investigate the method using Interactive-CLIP with T2V interaction as the backbone.
 Keeping the other hyperparameters constant, we condition different top-$n$ layers of the self-attention modules.
@@
 Figure 5: Case study of InterCLIP-MEP.
-Figure 6: Visual comparison between Interactive-CLIP and original CLIP.
-Case Study
+Figure 6: Visual comparison between InterCLIP and CLIP.
+Case Study and Visualization
 As shown in Figure 5, we select three examples to further demonstrate the robustness of InterCLIP-MEP.
-We observe that direct predictions through the classification module $\mathcal{F}_{c}$ result in high entropy and incorrect outcomes.
-However, because the InterCLIP-MEP framework leverages the Memory-Enhanced Predictor (MEP) to utilize historical knowledge of test samples during inference, This approach effectively mitigates the issue for cases where the classification module fails to correctly identify the results by using the historical knowledge of the test samples.
-Visualization
-To further validate that Interactive-CLIP can more effectively capture the interactive information between text and images compared to original CLIP, thereby aiding in the detection of sarcasm cues, we visualize the areas of focus during the inference process of the visual model in Figure 6.
-We observe that in the first example, which complains about the unpleasant odor caused by the sea lions, Interactive-CLIP focuses more accurately on the location of the sea lions compared to the original CLIP.
-In the second example, which complains about traffic congestion, Interactive-CLIP correctly focuses on the distribution of cars on the road, whereas the original CLIP’s focus is scattered.
-Conclusion and Future Work
+We observe that direct predictions through the classification module $\mathcal{F}_{c}$ result in high prediction entropy and incorrect outcomes.
+However, MEP effectively mitigates the issue for cases where $\mathcal{F}_{c}$ fails to correctly identify the results by using the historical knowledge of the test samples.
+To further validate that InterCLIP can more effectively capture the interactive information between text and images compared to original CLIP, thereby aiding in the detection of sarcasm cues, we visualize the areas of focus during the inference process of the visual model in Figure 6.
+We observe that in the first example, which complains about the unpleasant odor caused by the sea lions, InterCLIP focuses more accurately on the location of the sea lions compared to the original CLIP.
+In the second example, which complains about traffic congestion, InterCLIP correctly focuses on the distribution of cars on the road, whereas the original CLIP’s focus is scattered.
+Conclusion
 In this paper, we propose a novel framework for robust multi-modal sarcasm detection, termed InterCLIP-MEP.
-We introduce Interactive CLIP (InterCLIP) as the backbone for extracting sample representations.
+We introduce an Interactive CLIP (InterCLIP) as the backbone for extracting sample representations.
 Unlike the original CLIP, InterCLIP embeds representations from different modalities into the encoder, effectively capturing the intricate interactions between text and image.
-For inference, we introduce a Memory-Enhanced Predictor (MEP) that leverages historical knowledge of test samples to enhance inference, providing a more robust alternative to directly using the classification module for sarcasm recognition.
-This approach significantly improves the detection of sarcastic samples.
+For inference, we introduce a Memory-Enhanced Predictor (MEP) that leverages historical knowledge of valuable test samples to enhance inference, providing a more robust and reliable multi-modal sarcasm detection.
 Our experimental results demonstrate that InterCLIP-MEP achieves state-of-the-art performance on the MMSD2.0 benchmark.
-In future work, we plan to extend our approach to sarcasm recognition tasks across additional modalities, such as speech and video.
-Acknowledgments
-The research described in the paper received funding from the Anhui Provincial Natural Science Foundation through Grant No.
-2308085MF220.
-Additionally, it was co-supported by the Anhui University Natural Science Foundation, which provided funding through Grant Nos.
-2022AH050972 and KJ2021A0516.

### v4

@@
 Sarcasm, with its inherent subtlety and complexity, plays a significant role in human communication by often conveying irony, mockery, or hidden intentions (Muecke 1982; Gibbs and O’Brien 1991; Gibbs and Colston 2007).
-Figure 1: Comparative analysis of multi-modal sarcasm detection pipelines.
-Panels (a), (b), and (c) depict the pipeline comparisons between our InterCLIP-MEP method and other existing methods.
-Panel (d) utilizes GradCAM to visually demonstrate the superior ability of the InterCLIP backbone over the traditional CLIP backbone in identifying multi-modal sarcasm cues, which are marked with red boxes.
 The automatic detection of sarcasm from text has been a significant research focus, aiding tasks such as sentiment analysis and intent mining (Pang, Lee et al.
@@
 Therefore, the need for multi-modal sarcasm detection has become increasingly apparent, which presents significant challenges, requiring the system to capture and comprehend the intricate interplay between textual and visual information to identify sarcasm cues.
+Figure 1: An overview of the shortcomings of existing multi-modal sarcasm detection pipelines.
+In panels (a) and (b), we present two multi-modal sarcasm detection pipelines, with shortcomings indicated by a red question mark.
+In panel (c), we visually show an example of multi-modal sarcasm cues correctly or incorrectly recognized in a sarcasm sample.
 Despite the notable advancements made by existing studies (Xu, Zeng, and Mao 2020; Pan et al.
@@
 2023; Wei et al.
-2024) use separate uni-modal encoders, such as BERT (Devlin et al.
+2024) use dual unimodal encoders, such as BERT (Devlin et al.
 2019) for text and ViT (Dosovitskiy et al.
-2020) for image, as the backbone to encode samples.
+2021) for image, as the backbone to encode samples.
 This approach makes it difficult for the model to capture the subtle interactive information between text and image.
@@
 2021), to encode samples and demonstrated notable results on the MMSD2.0 benchmark, the original CLIP only aligns text and images directly, lacking the nuanced semantic space required for sarcasm detection.
+To illustrate the identification of multi-modal sarcasm cues, we present in Figure 1(c) an example of a sarcasm sample being correctly and incorrectly recognized.
 Furthermore, as depicted in Figure 1(a), some approaches (Xu, Zeng, and Mao 2020; Pan et al.
@@
 In this paper, building on these insights, we propose InterCLIP-MEP, a novel framework designed to enhance the reliability of multi-modal sarcasm detection.
-As shown in Figure 1(c), our InterCLIP-MEP introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone to enhance sample encoding for detecting multi-modal sarcasm.
+Our InterCLIP-MEP introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone to enhance sample encoding for detecting multi-modal sarcasm.
 InterCLIP enhances the capture of interactive information between text and image by embedding representations from one modality into the encoder of the other, leading to a more nuanced understanding of multi-modal sarcasm cues.
-In Figure 1(d), we utilized GradCAM (Selvaraju et al.
-2016) visualization to compare the effectiveness of our InterCLIP backbone against the CLIP backbone in detecting multi-modal sarcasm cues.
-During inference, InterCLIP-MEP introduces a Memory-Enhanced Predictor (MEP) to build dynamically a non-parametric classifier from valuable historical features to identify multi-modal sarcasm.
+During inference, Our InterCLIP-MEP introduces a Memory-Enhanced Predictor (MEP) to build dynamically a non-parametric classifier from valuable historical features to identify multi-modal sarcasm.
 Specifically, we design an efficient training strategy to adapt InterCLIP for MEP.
@@
 Figure 2: Overview of our framework.
-The left part of the figure shows the overall structure and training process of Interactive-CLIP.
-The right part illustrates the process of MEP identifying sarcastic samples.
+Training Interactive CLIP (InterCLIP).
+Vision and text representations are extracted using separate encoders and embedded into the top-n layers of the opposite modality’s encoder for interaction.
+The top-n layers are fine-tuned with LoRA, while the rest of the encoder remains frozen.
+Final vision and text representations are concatenated and used to train a classification module for sarcasm detection.
+A projection module is also trained to project representations into a latent space.
+Memory-Enhanced Predictor (MEP).
+During inference, InterCLIP generates interactive representations.
+The classification module assigns pseudo-labels, and the projection module provides projection features.
+MEP updates dynamic memory with these features and pseudo-labels.
+The final prediction is made by comparing the current sample’s projected feature with those in memory.
 Related Work
@@
 Schifanella et al.
-(2016) were among the first to use multi-modal posts from social platforms to identify sarcasm clues from both text and images.
+(2016) were among the first to use multi-modal posts from social platforms to identify sarcasm cues from both text and images.
 Cai, Cai, and Wan (2019) introduced the MMSD benchmark for multi-modal sarcasm detection, demonstrating its effectiveness with a hierarchical fusion model incorporating image attributes.
@@
 Furthermore, their approach is limited to general classification tasks and does not address the complexities of multi-modal sarcasm detection.
-Mermory-enhanced prediction.
+Memory-enhanced prediction.
 Inspired by cognitive science (Stokes 2015; Baddeley 2000), memory has been introduced to enhance neural networks (Weston, Chopra, and Bordes 2014; Sukhbaatar et al.
@@
 The text encoder $\mathcal{T}$ extracts the vanilla text representation $\mathbf{F}_{t}$:
-$\mathbf{F}_{t}=\mathcal{T}(T)=\{h^{t}_{\textrm{bos}}(t_{\textrm{bos}}),\\ h^{t}_{1}(t_{1}),\dots,h^{t}_{n}(t_{n}),h^{t}_{\textrm{eos}}(t_{\textrm{eos}})\},$ (1)
+$\mathbf{F}_{t}=\mathcal{T}(T)=\{h^{t}_{\textrm{bos}}(t_{\textrm{bos}}),h^{t}_{1}(t_{1}),\dots,h^{t}_{n}(t_{n}),h^{t}_{\textrm{eos}}(t_{\textrm{eos}})\},$ (1)
 where $t_{i}$ denotes a text token, $n$ is the length of $T$ after tokenization, $t_{\textrm{bos}}$ and $t_{\textrm{eos}}$ are special tokens required by the text encoder.
@@
 We use $\tilde{h}^{t}_{i}(\cdot)\in\mathbb{R}^{d_{t}}$ and $\tilde{h}^{v}_{i}(\cdot)\in\mathbb{R}^{d_{v}}$ to denote the re-encoded interactive representations of each text token and image patch, respectively.
-In particular, we condition only the top-$n$ self-attention layers of the text or vision encoder, where $n$ is a hyperparameter that will be analyzed in the experiment section.
-The structure of the conditioned self-attention layers is illustrated in Figure 3.
-We denote the input representations to the self-attention layers as $\mathbf{H}_{t/v}$, which are the output representations from the previous layer.
+To be specific, we condition only the top-$n$ self-attention layers of the text or vision encoder, where $n$ is a hyperparameter that will be analyzed in the experiment section.
+Figure 3 illustrates the structure of the conditioned self-attention layers.
+Given that the text and vision encoder in CLIP share a similar architecture, for simplicity, we denote the input representations to the self-attention layers of the text or vision encoder as $\mathbf{H}_{t/v}$, which are derived from the outputs of the previous layer.
 The previous layer can either be conditioned or non-conditioned.
@@
 Given the similarity between the self-attention layers of the vision encoder and the text encoder, we use the text encoder $\mathcal{T}$ to illustrate the process as follows:
-$\mathbf{F}_{v}^{{}^{\prime}}=\mathcal{F}_{t}(\mathbf{F}_{v}),\mathbf{F}_{v}\in\mathbb{R}^{{m}\times{d_{v}}},\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{{m}\times{d_{t}}},$ (4)
-$\mathbf{H}^{{}^{\prime}}_{t}=\textrm{Attn}_{t}(\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}})_{\textrm{[:n]}},\\ \mathbf{H}_{t},\mathbf{H}^{{}^{\prime}}_{t}\in\mathbb{R}^{{n}\times{d_{t}}},\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{{(n+m)}\times{d_{t}}},$ (5)
-$\mathbf{H}^{{}^{\prime\prime}}_{t}=\mathcal{H}_{t}(\mathbf{H}^{{}^{\prime}}_{t})+\mathcal{G}_{t}(\mathbf{H}^{{}^{\prime}}_{t})\cdot\textrm{tanh}(\beta_{t}),\mathbf{H}^{{}^{\prime\prime}}_{t}\in\mathbb{R}^{{n}\times{d_{t}}}.$ (6)
+$\begin{split}\mathbf{F}_{v}^{{}^{\prime}}&=\mathcal{F}_{t}(\mathbf{F}_{v}),\quad\mathbf{F}_{v}\in\mathbb{R}^{m\times d_{v}},\quad\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{m\times d_{t}},\\ \mathbf{H}^{{}^{\prime}}_{t}&=\textrm{Attn}_{t}(\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}})_{[:n]},\\ &\phantom{=}\mathbf{H}_{t},\mathbf{H}^{{}^{\prime}}_{t}\in\mathbb{R}^{n\times d_{t}},\quad\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{(n+m)\times d_{t}},\\ \mathbf{H}^{{}^{\prime\prime}}_{t}&=\mathcal{H}_{t}(\mathbf{H}^{{}^{\prime}}_{t})+\mathcal{G}_{t}(\mathbf{H}^{{}^{\prime}}_{t})\cdot\textrm{tanh}(\beta_{t}),\quad\mathbf{H}^{{}^{\prime\prime}}_{t}\in\mathbb{R}^{n\times d_{t}}.\end{split}$ (4)
 Here, $\oplus$ denotes the concatenation operation, and $\beta_{t}$ is a learnable parameter initialized to 0 to ensure training stability.
@@
 Memory-Enhanced Predictor
+As depicted in Figure 2 (right), we present the Memory-Enhanced Predictor (MEP) that builds upon the learned InterCLIP, along with the classification module and the projection module, leveraging the valuable historical knowledge of test samples to enhance sarcasm detection.
 Algorithm 1 Memory-Enhanced Predictor
@@
 24:  end for
-As depicted in Figure 2 (right), we present the Memory-Enhanced Predictor (MEP) that builds upon the learned InterCLIP, along with the classification module and the projection module, leveraging the valuable historical knowledge of test samples to enhance sarcasm detection.
 The detailed computational process of MEP is provided in Algorithm 1, where $N_{\textrm{test}}$ denotes the number of test samples.
@@
 RoBERTa* (Liu et al. 2019) 79.66 76.21 76.74 75.70
-Image ResNet* (He et al. 2015) 65.50 57.58 61.17 54.39
-ViT* (Dosovitskiy et al. 2020) 72.02 69.72 65.26 74.83
+Image ResNet* (He et al. 2016) 65.50 57.58 61.17 54.39
+ViT* (Dosovitskiy et al. 2021) 72.02 69.72 65.26 74.83
 Text-Image HFM* (Cai, Cai, and Wan 2019) 70.57 66.88 64.84 69.05
@@
 Baselines.
-We compare the effectiveness of the InterCLIP-MEP framework against several uni-modal and multi-modal methods.
+We compare the effectiveness of the InterCLIP-MEP framework against several unimodal and multi-modal methods.
 For text modality methods, we compare with TextCNN (Kim 2014), Bi-LSTM (Graves and Schmidhuber 2005), SMSD (Xiong et al.
@@
 For image modality methods, we compare with ResNet (He et al.
-2015) and ViT (Dosovitskiy et al.
-2020).
+2016) and ViT (Dosovitskiy et al.
+2021).
 For state-of-the-art multi-modal methods, we compare with the following:
@@
 We present all results in Figure 4.
-Figure 4: Hyperparameter study curves for InterCLIP-MEP w/ T2V. Figure 4(d) compares results with those from using only the classification module $\mathcal{F}_{c}$ for prediction.
+Figure 4: Hyperparameter study curves for InterCLIP-MEP w/ T2V. Panel (d) compares results with those from using only the classification module $\mathcal{F}_{c}$ for prediction.
 Figure 4(a) shows that conditioning the top four self-attention layers yields the best results.

### v5

@@
 Abstract
-The prevalence of sarcasm in social media, conveyed through text-image combinations, presents significant challenges for sentiment analysis and intention mining.
-Existing multi-modal sarcasm detection methods have been proven to overestimate performance, as they struggle to effectively capture the intricate sarcastic cues that arise from the interaction between an image and text.
-To address these issues, we propose InterCLIP-MEP, a novel framework for multi-modal sarcasm detection.
-Specifically, we introduce an Interactive CLIP (InterCLIP) as the backbone to extract text-image representations, enhancing them by embedding cross-modality information directly within each encoder, thereby improving the representations to capture text-image interactions better.
-Furthermore, an efficient training strategy is designed to adapt InterCLIP for our proposed Memory-Enhanced Predictor (MEP).
-MEP uses a dynamic, fixed-length dual-channel memory to store historical knowledge of valuable test samples during inference.
-It then leverages this memory as a non-parametric classifier to derive the final prediction, offering a more robust recognition of multi-modal sarcasm.
-Experiments demonstrate that InterCLIP-MEP achieves state-of-the-art performance on the MMSD2.0 benchmark, with an accuracy improvement of 1.08% and an F1 score improvement of 1.51% over the previous best method.
-Introduction
-Sarcasm, with its inherent subtlety and complexity, plays a significant role in human communication by often conveying irony, mockery, or hidden intentions (Muecke 1982; Gibbs and O’Brien 1991; Gibbs and Colston 2007).
-The automatic detection of sarcasm from text has been a significant research focus, aiding tasks such as sentiment analysis and intent mining (Pang, Lee et al.
-2008; Tsur, Davidov, and Rappoport 2010; Bouazizi and Ohtsuki 2015).
-With the rapid development of social media platforms like Twitter and Reddit, users frequently use text-image combinations to convey their messages.
-Therefore, the need for multi-modal sarcasm detection has become increasingly apparent, which presents significant challenges, requiring the system to capture and comprehend the intricate interplay between textual and visual information to identify sarcasm cues.
+Sarcasm in social media, often expressed through text-image combinations, poses challenges for sentiment analysis and intention mining.
+Current multi-modal sarcasm detection methods have been demonstrated to overly rely on spurious cues within the textual modality, revealing a limited ability to genuinely identify sarcasm through nuanced text-image interactions.
+To solve this problem, we propose InterCLIP-MEP, which introduces Interactive CLIP (InterCLIP) with an efficient training strategy to extract enriched text-image representations by embedding cross-modal information directly into each encoder.
+Additionally, we design a Memory-Enhanced Predictor (MEP) with a dynamic dual-channel memory that stores valuable test sample knowledge during inference, acting as a non-parametric classifier for robust sarcasm recognition.
+Experiments on two benchmarks demonstrate that InterCLIP-MEP achieves state-of-the-art performance, with significant accuracy and F1 score improvements on MMSD and MMSD2.0.
+Our code is available at https://github.com/CoderChen01/InterCLIP-MEP.
+InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection
+Junjie Chen1,3thanks: Junjie is a research assistant at the EPIC Lab of Shanghai Jiao Tong University, working remotely with Linfeng Zhang.
+Hang Yu2thanks: Co-corresponding author.
+Subin Huang3thanks: Corresponding author.
+Sanmin Liu3    Linfeng Zhang1 1Shanghai Jiao Tong University, China    2Shanghai University, China 3Anhui Polytechnic University, China jorji.chen@gmail.com,  yuhang@shu.edu.cn,  subinhuang@ahpu.edu.cn, sanmin.liu@ahpu.edu.cn,  zhanglinfeng@sjtu.edu.cn
+1 Introduction
+Sarcasm, with its subtlety and complexity, plays a key role in communication by conveying irony, mockery, or hidden meanings Muecke (1982); Gibbs and O’Brien (1991); Gibbs and Colston (2007).
+Automatically detecting sarcasm in text has become an important research area, supporting tasks like sentiment analysis and intent mining Pang et al.
+(2008); Tsur et al.
+(2010); Bouazizi and Ohtsuki (2015).
 Figure 1: An overview of the shortcomings of existing multi-modal sarcasm detection pipelines.
-In panels (a) and (b), we present two multi-modal sarcasm detection pipelines, with shortcomings indicated by a red question mark.
-In panel (c), we visually show an example of multi-modal sarcasm cues correctly or incorrectly recognized in a sarcasm sample.
-Despite the notable advancements made by existing studies (Xu, Zeng, and Mao 2020; Pan et al.
-2020; Liang et al.
-2021, 2022) in this field, Qin et al.
-(2023) have highlighted that the MMSD benchmark (Cai, Cai, and Wan 2019) employed in these works contains spurious cues.
-These cues lead to the development of biased models, thereby inflating the perceived effectiveness of these models in detecting sarcasm cues from multi-modal data.
-By introducing MMSD2.0, a refined benchmark that eliminates spurious cues and corrects mislabeled samples, Qin et al.
-(2023) demonstrates a substantial performance decline of existing methods when evaluated on MMSD2.0.
-Although they proposed a promising method that employs a Multi-view CLIP for reliable multi-modal sarcasm detection, it still leaves room for improvement, similar to their previous approaches.
-As shown in Figure 1(a) and Figure 1(b), some existing methods (Xu, Zeng, and Mao 2020; Pan et al.
-2020; Liang et al.
-2021, 2022; Wen, Jia, and Yang 2023; Tian et al.
-2023; Wei et al.
-2024) use dual unimodal encoders, such as BERT (Devlin et al.
-2019) for text and ViT (Dosovitskiy et al.
-2021) for image, as the backbone to encode samples.
-This approach makes it difficult for the model to capture the subtle interactive information between text and image.
-Although Qin et al.
-(2023) utilized a multi-modal pre-trained backbone, CLIP (Radford et al.
-2021), to encode samples and demonstrated notable results on the MMSD2.0 benchmark, the original CLIP only aligns text and images directly, lacking the nuanced semantic space required for sarcasm detection.
-To illustrate the identification of multi-modal sarcasm cues, we present in Figure 1(c) an example of a sarcasm sample being correctly and incorrectly recognized.
-Furthermore, as depicted in Figure 1(a), some approaches (Xu, Zeng, and Mao 2020; Pan et al.
-2020; Liang et al.
-2021, 2022; Wen, Jia, and Yang 2023; Tian et al.
-2023) directly classify sarcastic samples based on their representations.
-This often leads to high prediction entropy and erroneous outcomes, especially for hard cases with greater uncertainty.
-To the best of our knowledge, only Wei et al.
-(2024) considers leveraging historical knowledge of samples to aid in detection.
-As shown in Figure 1(b), they constructed a static semantic space that stores features of all training samples.
-When identifying sarcastic samples, they query $k$-nearest neighbors in this semantic space and use a voting mechanism to derive the final predictions.
-However, relying heavily on a static semantic space may limit its ability to adapt to various test samples.
-In this paper, building on these insights, we propose InterCLIP-MEP, a novel framework designed to enhance the reliability of multi-modal sarcasm detection.
-Our InterCLIP-MEP introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone to enhance sample encoding for detecting multi-modal sarcasm.
-InterCLIP enhances the capture of interactive information between text and image by embedding representations from one modality into the encoder of the other, leading to a more nuanced understanding of multi-modal sarcasm cues.
-During inference, Our InterCLIP-MEP introduces a Memory-Enhanced Predictor (MEP) to build dynamically a non-parametric classifier from valuable historical features to identify multi-modal sarcasm.
-Specifically, we design an efficient training strategy to adapt InterCLIP for MEP.
-Based on the sample representations encoded by InterCLIP, we train a classification module to assign pseudo-labels to samples.
-Simultaneously, we train a projection module to map samples into a latent space.
-In this latent space, same-class samples are pulled closer together, while different-class samples are pushed further apart.
-Additionally, we fine-tune the weights of the self-attention modules in InterCLIP’s text and vision encoders using the low-rank adaptation (LoRA) technique (Hu et al.
-2022), building upon the original CLIP weights.
-This approach allows for more efficient adaptation without significantly increasing the computational burden.
-Building on InterCLIP and the training strategy, MEP leverages historical knowledge of valuable test samples to identify multi-modal sarcasm.
-It maintains a dynamic, fixed-length dual-channel memory, with one channel storing historical information of non-sarcastic samples and the other storing information of sarcastic samples.
-MEP then utilizes this memory as a non-parametric classifier to compute the final prediction for each sample, offering more robust and reliable multi-modal sarcasm detection.
-Overall, our contributions can be summarized as follows:
-  • We propose InterCLIP-MEP, a novel framework for reliable multi-modal sarcasm detection, which introduces Interactive CLIP (InterCLIP) improving the encoding of text-image interactions by embedding cross-modality information within each encoder and Memory-Enhanced Predictor (MEP) offering more robust and reliable multi-modal sarcasm detection.
-  • We develop an efficient training strategy that ensures computational efficiency and superior performance, bridging InterCLIP and MEP.
-  • Our extensive experiments on the more reliable benchmark, MMSD2.0, demonstrate that InterCLIP-MEP [^1] https://github.com/CoderChen01/InterCLIP-MEP outperforms existing state-of-the-art methods.
+In panels (a) and (b), we present two main multi-modal sarcasm detection pipelines, with shortcomings indicated by a red question mark.
+In panel (c), we visually show an example of multi-modal sarcasm cues correctly or incorrectly recognized in a multi-modal sarcasm sample.
+With the rise of social media platforms like Twitter and Reddit, users often use text-image combinations to express their messages.
+As a result, multi-modal sarcasm detection is increasingly important, posing challenges in understanding the complex relationship between textual and visual cues to identify sarcasm.
+As shown in Figures 1(a) and 1(b), many methods rely on dual unimodal pre-trained encoders, such as ViTDosovitskiy et al.
+(2021) and BERT Devlin et al.
+(2019), as the backbone for encoding text-image pairs, followed by specific feature fusion Xu et al.
+(2020); Pan et al.
+(2020); Liang et al.
+(2021, 2022); Wen et al.
+(2023); Tian et al.
+(2023); Wei et al.
+(2024).
+However, this approach may not capture multi-modal sarcasm cues as effectively as multi-modal pre-trained models like CLIP Radford et al.
+(2021).
+In Figure 1(a), the use of a learnable classification head to predict labels from fused representations is common but often associated with high predictive entropy and significant uncertainty.
+Wei et al.
+(2024) pioneered the construction of a static semantic space using historical training samples, where more robust predictions are obtained during inference through KNN-based querying and voting, as illustrated in Figure 1(b).
+However, while CLIP has been proven to be effective in serving as a text-image encoder for multi-modal sarcasm detection Qin et al.
+(2023), it still struggles to capture multi-modal sarcasm cues due to the inherent inconsistency of sarcasm, which conflicts with CLIP’s direct alignment of text and image.
+Furthermore, relying on a static semantic space for inference is ill-suited to handle the dynamic nature of evolving sample distributions.
+In fact, Qin et al.
+(2023) have shown that many models rely on spurious cues in the MMSD benchmark Cai et al.
+(2019), leading to biased results.
+Building on the limitations of prior multi-modal sarcasm detection approaches, we propose Interactive CLIP (InterCLIP) as the backbone, embedding cross-modal representations directly into text and vision encoders to enhance the understanding of multi-modal sarcasm cues (Figure 2, left).
+To complement this, we design a Memory-Enhanced Predictor (MEP) that dynamically utilizes historical test sample features to create a more adaptive and reliable non-parametric classifier for final predictions (Figure 2, right).
+Together, these components form the proposed framework, InterCLIP-MEP.
+Furthermore, InterCLIP-MEP employs an efficient training strategy that fine-tunes cross-modal interactions through a lightweight adaptation mechanism, ensuring computational efficiency while delivering state-of-the-art performance (Figure 2, left).
+Overall, our contributions are as follows:
+  • We introduce InterCLIP-MEP [^1] Our code is available in the supplementary material., a novel framework for multi-modal sarcasm detection, which combines Interactive CLIP (InterCLIP) for enhanced text-image interaction encoding and Memory-Enhanced Predictor (MEP) for more robust and reliable sarcasm predictions.
+  • We propose an efficient training strategy that significantly reduces computational overhead compared to state-of-the-art methods.
+By introducing approximately 20.6x fewer trainable parameters, our approach reduces GPU memory usage by about 2.5x and accelerates computation by roughly 8.7x with a batch size of 128, all while maintaining superior performance on a single NVIDIA RTX 4090 GPU.
+  • Through extensive experiments on the MMSD and MMSD2.0 benchmarks, we show that InterCLIP-MEP improves accuracy by 1.08% and F1 score by 1.51% over state-of-the-art methods, especially on MMSD2.0.
+2 Related Work
+Early research in sarcasm detection focused primarily on text data Bouazizi and Ohtsuki (2015); Amir et al.
+(2016); Baziotis et al.
+(2018).
+With the rise of social media, detecting sarcasm in text-image pairs has become more challenging, driving the development of multi-modal approaches.
+Schifanella et al.
+(2016) were among the first to explore multi-modal social media posts for identifying sarcasm cues.
+Building on this, Cai et al.
+(2019) introduced the MMSD benchmark, demonstrating the effectiveness of a hierarchical fusion model that integrates image features.
+This benchmark has since become a foundation for multi-modal sarcasm detection, inspiring subsequent studies Xu et al.
+(2020); Pan et al.
+(2020); Liang et al.
+(2021, 2022); Liu et al.
+(2022); Qin et al.
+(2023); Wen et al.
+(2023); Tian et al.
+(2023); Wei et al.
+(2024).
+However, the MMSD benchmark was later found to contain spurious cues that could lead to model bias Qin et al.
+(2023).
+To mitigate this, Qin et al.
+(2023) introduced the MMSD2.0 benchmark, which removes these cues and corrects mislabeled samples.
+Re-evaluations on MMSD2.0 revealed significant performance drops in existing models, emphasizing the need for more robust approaches.
+In parallel, Tang et al.
+(2024) explored the use of large language models (LLMs) in multi-modal sarcasm detection, incorporating instruction templates and retrieval modules.
+While promising, the performance improvements were modest compared to the substantial computational cost.
+In this work, we present InterCLIP-MEP, a lightweight and efficient framework that achieves competitive performance without the high resource demands of LLM-based approaches.
+By overcoming the limitations of current methods, our approach offers a practical and scalable solution for multi-modal sarcasm detection.
 Figure 2: Overview of our framework.
-Training Interactive CLIP (InterCLIP).
-Vision and text representations are extracted using separate encoders and embedded into the top-n layers of the opposite modality’s encoder for interaction.
+(I) Training Interactive CLIP (InterCLIP): Vision and text representations are extracted using separate encoders and embedded into the top-$n$ layers of the opposite modality’s encoder for interaction.
 The top-$n$ layers are fine-tuned with LoRA, while the rest of the encoder remains frozen.
-Final vision and text representations are concatenated and used to train a classification module for sarcasm detection.
+Final vision and text representations are concatenated and used to train a classification module for identifying multi-modal sarcasm.
 A projection module is also trained to project representations into a latent space.
-Memory-Enhanced Predictor (MEP).
-During inference, InterCLIP generates interactive representations.
+(II) Memory-Enhanced Predictor (MEP): During inference, InterCLIP generates interactive representations.
 The classification module assigns pseudo-labels, and the projection module provides projection features.
 MEP updates dynamic memory with these features and pseudo-labels.
-The final prediction is made by comparing the current sample’s projected feature with those in memory.
-Related Work
-Multi-modal sarcasm detection.
-Early work in sarcasm detection primarily focused on textual data (Bouazizi and Ohtsuki 2015; Amir et al.
-2016; Baziotis et al.
-2018).
-With the rise of social media, detecting sarcasm from text-image data has become increasingly challenging, leading to the development of several multi-modal approaches.
-Schifanella et al.
-(2016) were among the first to use multi-modal posts from social platforms to identify sarcasm cues from both text and images.
-Cai, Cai, and Wan (2019) introduced the MMSD benchmark for multi-modal sarcasm detection, demonstrating its effectiveness with a hierarchical fusion model incorporating image attributes.
-This benchmark has been widely used in subsequent studies (Xu, Zeng, and Mao 2020; Pan et al.
-2020; Liang et al.
-2021, 2022; Liu, Wang, and Li 2022; Qin et al.
-2023; Wen, Jia, and Yang 2023; Tian et al.
-2023; Wei et al.
-2024).
-However, MMSD has been proven to contain spurious cues that could bias models (Qin et al.
-2023).
-To address this, Qin et al.
-(2023) introduced the MMSD2.0 benchmark, which removed these cues and corrected mislabeled samples.
-They observed a significant performance drop when re-evaluating state-of-the-art methods on MMSD2.0 and proposed the need for more stable multi-modal sarcasm detection methods.
-In this work, we introduce the InterCLIP-MEP framework towards more reliable multi-modal sarcasm detection.
-CLIP adaptation.
-The Contrastive Language-Image Pretraining (CLIP) model (Radford et al.
-2021) excels in vision-language tasks.
-Adapting CLIP for specific domains has shown substantial improvements, as demonstrated by Li, Shakhnarovich, and Yeh (2022) for phrase localization, Liang et al.
-(2023) for open-vocabulary semantic segmentation, and Wang et al.
-(2023) for action recognition.
-In this work, inspired by Ganz et al.
-(2024), we conditionally enhance both the text and visual encoders of CLIP, making it more effective in capturing the interplay between text and images to identify sarcasm.
-Unlike Ganz et al.
-(2024), who focused solely on embedding text into the visual encoder, we also explore embedding images into the text encoder.
-Furthermore, their approach is limited to general classification tasks and does not address the complexities of multi-modal sarcasm detection.
-Memory-enhanced prediction.
-Inspired by cognitive science (Stokes 2015; Baddeley 2000), memory has been introduced to enhance neural networks (Weston, Chopra, and Bordes 2014; Sukhbaatar et al.
-2015).
-Several studies (Wu et al.
-2018; Wen, Jia, and Yang 2023) have used memory mechanisms to improve model training, and some (Zhang et al.
-2024; Wei et al.
-2024) leverage memory to store historical knowledge, enhancing prediction accuracy.
-In this work, we introduce a memory-enhanced predictor for multi-modal sarcasm detection.
-In contrast to other methods, our memory dynamically updates during testing, utilizing relevant historical information for improved accuracy and robustness.
-Methodology
+The final prediction of the current sample is made by comparing its projected feature with those in memory.
+3 Methodology
 An overview of InterCLIP-MEP is illustrated in Figure 2.
 Initially, we elaborate on the Interactive CLIP (InterCLIP) and its training strategy, followed by an in-depth explanation of the Memory-Enhanced Predictor (MEP).
-Interactive CLIP
+3.1 Interactive CLIP
 The input to Interactive CLIP (InterCLIP) is a text-image pair $\mathcal{P}=(T,I)$, where $T$ represents a piece of text and $I$ represents an image.
 Here, for simplicity, we do not consider the case of batch inputs.
-The text encoder $\mathcal{T}$ extracts the vanilla text representation $\mathbf{F}_{t}$:
-$\mathbf{F}_{t}=\mathcal{T}(T)=\{h^{t}_{\textrm{bos}}(t_{\textrm{bos}}),h^{t}_{1}(t_{1}),\dots,h^{t}_{n}(t_{n}),h^{t}_{\textrm{eos}}(t_{\textrm{eos}})\},$ (1)
+The text encoder $\mathcal{T}$ extracts the vanilla text representations $\mathbf{F}_{t}$:
+$\mathbf{F}_{t}$ $=\mathcal{T}(T)$
+=$=\{h^{t}_{\textrm{bos}}(t_{\textrm{bos}}),h^{t}_{1}(t_{1}),\dots,h^{t}_{n}(t_{n}),h^{t}_{\textrm{eos}}(t_{\textrm{eos}})\},$ (1)
 where $t_{i}$ denotes a text token, $n$ is the length of $T$ after tokenization, $t_{\textrm{bos}}$ and $t_{\textrm{eos}}$ are special tokens required by the text encoder.
 Here, $h^{t}_{i}(\cdot)\in\mathbb{R}^{d_{t}}$ represents the $d_{t}$-dimensional encoded representation of the corresponding token $t_{i}$, with $i$ ranging from 1 to $n$, including the beginning-of-sequence (bos) and end-of-sequence (eos) tokens.
-The vision encoder $\mathcal{V}$ extracts the vanilla image representation $\mathbf{F}_{v}$:
+The vision encoder $\mathcal{V}$ extracts the vanilla image representations $\mathbf{F}_{v}$:
 $\mathbf{F}_{v}=\mathcal{V}(I)=\{h^{v}_{\textrm{cls}}(p_{\textrm{cls}}),h^{v}_{1}(p_{1}),\dots,h^{v}_{m}(p_{m})\},$ (2)
@@
 Specifically, both $\mathbf{F}_{t}$ and $\mathbf{F}_{v}$ are representations from the final layer outputs of their respective encoders.
-Conditioning on $\mathbf{F}_{t}$ or $\mathbf{F}_{v}$, we can obtain the interactive text representation $\tilde{\mathbf{F}}_{t}$ or the interactive image representation $\tilde{\mathbf{F}}_{v}$:
+Conditioning on $\mathbf{F}_{t}$ or $\mathbf{F}_{v}$, we can obtain the interactive text representations $\tilde{\mathbf{F}}_{t}$ or the interactive image representations $\tilde{\mathbf{F}}_{v}$:
 $\tilde{\mathbf{F}}_{t}=\mathcal{T}(T|\mathbf{F}_{v}),\tilde{\mathbf{F}}_{v}=\mathcal{V}(V|\mathbf{F}_{t}).$ (3)
@@
 Figure 3 illustrates the structure of the conditioned self-attention layers.
-Given that the text and vision encoder in CLIP share a similar architecture, for simplicity, we denote the input representations to the self-attention layers of the text or vision encoder as $\mathbf{H}_{t/v}$, which are derived from the outputs of the previous layer.
+Given that the text and vision encoder in CLIP share a similar architecture, for brevity, we denote the input representations to the self-attention layers of the text or vision encoder as $\mathbf{H}_{t/v}$, which are derived from the outputs of the previous layer.
 The previous layer can either be conditioned or non-conditioned.
@@
 Following Ganz et al.
-(2024), we apply a gated projection layer $\mathcal{G}_{t/v}$ along with the self-attention’s projection head $\mathcal{H}_{t/v}$ using a learnable gating mechanism to compute the self-attention output representation $\mathbf{H}^{{}^{\prime\prime}}_{t/v}$.
+(2024), we apply a gated projection layer $\mathcal{G}_{t/v}$ along with the self-attention’s projection head $\mathcal{H}_{t/v}$ using a learnable gating mechanism to compute the self-attention output representations $\mathbf{H}^{{}^{\prime\prime}}_{t/v}$.
 Given the similarity between the self-attention layers of the vision encoder and the text encoder, we use the text encoder $\mathcal{T}$ to illustrate the process as follows:
-$\begin{split}\mathbf{F}_{v}^{{}^{\prime}}&=\mathcal{F}_{t}(\mathbf{F}_{v}),\quad\mathbf{F}_{v}\in\mathbb{R}^{m\times d_{v}},\quad\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{m\times d_{t}},\\ \mathbf{H}^{{}^{\prime}}_{t}&=\textrm{Attn}_{t}(\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}})_{[:n]},\\ &\phantom{=}\mathbf{H}_{t},\mathbf{H}^{{}^{\prime}}_{t}\in\mathbb{R}^{n\times d_{t}},\quad\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{(n+m)\times d_{t}},\\ \mathbf{H}^{{}^{\prime\prime}}_{t}&=\mathcal{H}_{t}(\mathbf{H}^{{}^{\prime}}_{t})+\mathcal{G}_{t}(\mathbf{H}^{{}^{\prime}}_{t})\cdot\textrm{tanh}(\beta_{t}),\quad\mathbf{H}^{{}^{\prime\prime}}_{t}\in\mathbb{R}^{n\times d_{t}}.\end{split}$ (4)
-Here, $\oplus$ denotes the concatenation operation, and $\beta_{t}$ is a learnable parameter initialized to 0 to ensure training stability.
-The subsequent computation follows the original CLIP (Radford et al.
-2021), ultimately yielding the interactive representations $\tilde{\mathbf{F}}_{t}$.
-Figure 3: Structure of the conditional self-attention layer.
-InterCLIP has three distinct modes to achieve the final fused feature $\tilde{h}^{f}\in\mathbb{R}^{d_{t}+d_{v}}$:
-  • T2V: The text encoder extracts the vanilla text representation $\mathbf{F}_{t}$.
-This text representation is then embedded into the vision encoder to obtain the conditioned vision representation $\tilde{\mathbf{F}}_{v}$.
-The final fused representation $\tilde{h}^{f}$ is formed by concatenating $h^{t}_{\textrm{eos}}(t_{\textrm{eos}})$ and $\tilde{h}^{v}_{\textrm{cls}}(p_{\textrm{cls}})$.
-  • V2T: The vision encoder extracts the vanilla image representation $\mathbf{F}_{v}$.
-This image representation is then embedded into the text encoder to obtain the conditioned text representation $\tilde{\mathbf{F}}_{t}$.
-The final fused representation $\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}(t_{\textrm{eos}})$ and $h^{v}_{\textrm{cls}}(p_{\textrm{cls}})$.
-  • Two-way (TW): The text and vision encoders extract the vanilla text and image representations $\mathbf{F}_{t}$ and $\mathbf{F}_{v}$, respectively.
-These representations are then embedded into the vision and text encoders to obtain the conditioned image representation $\tilde{\mathbf{F}}_{v}$ and text representation $\tilde{\mathbf{F}}_{t}$.
-The final fused representation $\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}(t_{\textrm{eos}})$ and $\tilde{h}^{v}_{\textrm{cls}}(p_{\textrm{cls}})$.
-We will analyze the effectiveness of these three modes of interaction in the experimental analysis.
-Training Strategy
+$\begin{split}\mathbf{F}_{v}^{{}^{\prime}}&=\mathcal{F}_{t}(\mathbf{F}_{v}),\quad\mathbf{F}_{v}\in\mathbb{R}^{m\times d_{v}},\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{m\times d_{t}},\\ \mathbf{H}^{{}^{\prime}}_{t}&=\textrm{Attn}_{t}(\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}})_{[:n]},\\ &\phantom{=}\mathbf{H}_{t},\mathbf{H}^{{}^{\prime}}_{t}\in\mathbb{R}^{n\times d_{t}},\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{(n+m)\times d_{t}},\\ \mathbf{H}^{{}^{\prime\prime}}_{t}&=\mathcal{H}_{t}(\mathbf{H}^{{}^{\prime}}_{t})+\mathcal{G}_{t}(\mathbf{H}^{{}^{\prime}}_{t})\cdot\textrm{tanh}(\beta_{t}),\mathbf{H}^{{}^{\prime\prime}}_{t}\in\mathbb{R}^{n\times d_{t}}.\end{split}$ (4)
+Here, $\oplus$ denotes the concatenation operation, and $\beta_{t}$ is a learnable gating parameter initialized to 0 to ensure training stability.
+The subsequent computation follows the original CLIP Radford et al.
+(2021), ultimately yielding the interactive representations $\tilde{\mathbf{F}}_{t}$.
+Figure 3: Structure of the conditional self-attention.
+InterCLIP supports three interaction modes for fusing text and image features into the final representation $\tilde{h}^{f}\in\mathbb{R}^{d_{t}+d_{v}}$:
+  • T2V: Text representations $\mathbf{F}_{t}$ are embedded into the vision encoder to produce interactive image representations $\tilde{\mathbf{F}}_{v}$.
+$\tilde{h}^{f}$ is formed by concatenating $h^{t}_{\textrm{eos}}$ and $\tilde{h}^{v}_{\textrm{cls}}$.
+  • V2T: Image representations $\mathbf{F}_{v}$ are embedded into the text encoder to produce interactive text representations $\tilde{\mathbf{F}}_{t}$.
+$\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}$ and $h^{v}_{\textrm{cls}}$.
+  • Two-way (TW): Both text and image representations $\mathbf{F}_{t}$ and $\mathbf{F}_{v}$ are embedded into each other’s encoders, resulting in $\tilde{\mathbf{F}}_{t}$ and $\tilde{\mathbf{F}}_{v}$.
+$\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}$ and $\tilde{h}^{v}_{\textrm{cls}}$.
+We will analyze the effectiveness of these three interaction modes in the experimental analysis.
+Training Strategy.
 As shown in Figure 2 (left), to adapt InterCLIP for MEP, we introduce an efficient training strategy.
 Using InterCLIP as the backbone to obtain fused features of the samples, we introduce a classification module and a projection module.
-The classification module $\mathcal{F}_{c}$ calculates the probability of a sample being sarcastic or non-sarcastic:
-$\hat{y}=\textrm{softmax}(\mathcal{F}_{c}(\tilde{H}^{f})),$ (5)
-where $\tilde{H}^{f}\in\mathbb{R}^{{N}\times(d_{t}+d_{v})}$ represents the fused representations of a batch of samples, and $N$ denotes the batch size.
-We optimize $\mathcal{F}_{c}$ using binary cross-entropy loss:
-$\mathcal{L}^{c}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i,1})+(1-y_{i})\log(1-\hat{y}_{i,0})\right],$ (6)
+Given the fused features of a batch of samples $\tilde{H}^{f}\in\mathbb{R}^{{N}\times(d_{t}+d_{v})}$, the classification module $\mathcal{F}_{c}$ calculates the probabilities $\hat{y}$ of these samples being sarcastic or non-sarcastic:
+$\hat{y}=\textrm{softmax}(\mathcal{F}_{c}(\tilde{H}^{f})),\quad\hat{y}\in\mathbb{R}^{N\times 2},$ (5)
+where $N$ denotes the batch size. We optimize $\mathcal{F}_{c}$ using binary cross-entropy loss:
+$\mathcal{L}^{c}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i,1})+(1-y_{i})\log(1-\hat{y}_{i,0})\right],$
 where $y_{i}$ denotes the label of the $i$-th sample, with sarcastic labeled as 1 and non-sarcastic as 0, and $\hat{y}_{i}$ denotes the prediction for the $i$-th sample.
@@
 $\mathcal{L}^{p}=\textrm{mean}(\hat{H}^{f}_{P}\cdot\hat{H}^{f^{T}}_{N})+\textrm{mean}(1-\hat{H}^{f}_{P}\cdot\hat{H}^{f^{T}}_{P})\\ +\textrm{mean}(1-\hat{H}^{f}_{N}\cdot\hat{H}^{f^{T}}_{N}),$ (8)
-where $\hat{H}^{f}_{P}$ and $\hat{H}^{f}_{N}$ represent the projected fused features of positive and negative samples, respectively.
-We fully train the modules $\mathcal{F}_{c}$, $\mathcal{F}_{p}$, the adapting projection layer $\mathcal{F}_{t/v}$, the gated projection layer $\mathcal{G}_{t/v}$, and the parameters $\beta_{t/v}$.
-We use LoRA (Hu et al.
-2022) to fine-tune parts of the weight matrices $\mathbf{W}$ in the self-attention modules of all encoders, specifically various combinations of $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$.
+where $\hat{H}^{f}_{P}$ and $\hat{H}^{f}_{N}$ represent the projected features of positive and negative samples, respectively.
+We fully train the modules $\mathcal{F}_{c}$, $\mathcal{F}_{p}$, the adapting projection layers ($\mathcal{F}_{t}$ and $\mathcal{F}_{v}$), the gated projection layers ($\mathcal{G}_{t}$ and $\mathcal{G}_{v}$), and the learnable gating parameters ($\beta_{t}$ and $\beta_{v}$).
+We use LoRA Hu et al.
+(2022) to fine-tune parts of the weight matrices $\mathbf{W}$ in the self-attention modules of all encoders, specifically various combinations of $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$.
 We consider $\mathbf{W}$ and the rank $r$ of LoRA as hyperparameters for our study.
-All modules are optimized by minimizing the joint loss:
+All learnable parts are optimized by minimizing the joint loss:
 $\mathcal{L}=\mathcal{L}^{c}+\mathcal{L}^{p}.$ (9)
-Memory-Enhanced Predictor
-As depicted in Figure 2 (right), we present the Memory-Enhanced Predictor (MEP) that builds upon the learned InterCLIP, along with the classification module and the projection module, leveraging the valuable historical knowledge of test samples to enhance sarcasm detection.
 Algorithm 1 Memory-Enhanced Predictor
@@
 20:     end if
-21:     $\textrm{logits}\leftarrow[\sum_{k=0}^{\mathcal{I}[0]}(\hat{h}^{f}_{i}\mathcal{M}[0]^{T})_{k},\sum_{k=0}^{\mathcal{I}[1]}(\hat{h}^{f}_{i}\mathcal{M}[1]^{T})_{k}]$
+21:     $\textrm{logits}\leftarrow\left[\sum_{k=0}^{\mathcal{I}[0]}(\hat{h}^{f}_{i}\mathcal{M}[0]^{T})_{k},\sum_{k=0}^{\mathcal{I}[1]}(\hat{h}^{f}_{i}\mathcal{M}[1]^{T})_{k}\right]$
 22:     $\hat{y}^{p}_{i}\leftarrow\textrm{softmax}(\textrm{logits})$
@@
 24:  end for
+3.2 Memory-Enhanced Predictor
+As depicted in Figure 2 (right), we present the Memory-Enhanced Predictor (MEP) that builds upon the learned InterCLIP, along with the classification module and the projection module, leveraging the valuable historical knowledge of test samples to enhance multi-modal sarcasm detection.
 The detailed computational process of MEP is provided in Algorithm 1, where $N_{\textrm{test}}$ denotes the number of test samples.
@@
 If the selected channel has available space, the sample’s projected features are added directly, and the prediction entropy is recorded.
-If the memory is full, the prediction entropy of all samples in the memory is compared with that of the current sample, replacing the samples with the highest entropy as necessary.
-Finally, the current sample’s projected features are combined with the historical features stored in both memory channels $\mathcal{M}$ using cosine similarity to yield the final prediction.
-All Sarcastic Non-sarcastic
-Train 19,812 9,572 10,240
-Validation 2,410 1,042 1,368
-Test 2,409 1,037 1,372
-Table 1: Statistics of the dataset from the MMSD2.0 benchmark (Qin et al. 2023).
-Modality Method Acc. (%) F1 (%) P (%) R (%)
-Text TextCNN* (Kim 2014) 71.61 69.52 64.62 75.22
-Bi-LSTM* (Graves and Schmidhuber 2005) 72.48 68.05 68.02 68.08
-SMSD* (Xiong et al. 2019) 73.56 69.97 68.45 71.55
-RoBERTa* (Liu et al. 2019) 79.66 76.21 76.74 75.70
-Image ResNet* (He et al. 2016) 65.50 57.58 61.17 54.39
-ViT* (Dosovitskiy et al. 2021) 72.02 69.72 65.26 74.83
-Text-Image HFM* (Cai, Cai, and Wan 2019) 70.57 66.88 64.84 69.05
-Att-BERT* (Pan et al. 2020) 80.03 77.04 76.28 77.82
-CMGCN* (Liang et al. 2022) 79.83 76.90 75.82 78.01
-HKE* (Liu, Wang, and Li 2022) 76.50 72.25 73.48 71.07
-DIP (Wen, Jia, and Yang 2023) 80.59 78.23 75.52 81.14
-DynRT (Tian et al. 2023) 70.37 68.55 63.02 75.15
-Multi-view CLIP* (Qin et al. 2023) 85.64 84.10 80.33 88.24
-G2SAM (Wei et al. 2024) 79.43 78.07 72.04 85.20
-InterCLIP-MEP w/o Inter ($L=1024$) 86.05 84.81 79.83 90.45
-InterCLIP-MEP w/ TW ($L=128$) 85.51 84.26 79.15 90.07
-InterCLIP-MEP w/ V2T ($L=640$) 86.26 85.00 80.17 90.45
-InterCLIP-MEP w/ T2V ($L=1024$) 86.72 85.61 80.20 91.80
+If the memory is full, the prediction entropy of all samples in the memory is compared with that of the current sample.
+Samples with the highest entropy are replaced, ensuring the retained samples have lower entropy.
+Finally, the current sample’s projected feature is combined with the historical features stored in both memory channels $\mathcal{M}$ using cosine similarity to yield the final prediction.
+4 Experiment
+MMSD/MMSD2.0 Sarcastic Non-sarcastic All
+Train 8,642/9,576 11,174/10,240 19,816/19,816
+Validation 959/1,042 1,451/1,368 2,410/2,410
+Test 959/1,037 1,450/1,372 2,409/2,409
+Table 1: Statistics of MMSD and MMSD2.0.
+Method MMSD2.0 MMSD
+Acc. (%) F1 (%) P (%) R (%) Acc. (%) F1 (%) P (%) R (%)
+Text
+TextCNN Kim (2014) 71.61* 69.52* 64.62* 75.22* 80.03* 75.32* 74.29* 76.39*
+Bi-LSTM Zhou et al. (2016) 72.48* 68.05* 68.02* 68.08* 81.90* 77.53* 76.66* 78.42*
+SMSD Xiong et al. (2019) 73.56* 69.97* 68.45* 71.55* 80.90* 75.82* 76.46* 75.18*
+RoBERTa Liu et al. (2019) 79.66* 76.21* 76.74* 75.70* 93.97* 92.45* 90.39* 94.59*
+Image
+ResNet He et al. (2016) 65.50* 57.58* 61.17* 54.39* 64.76* 61.53* 54.41* 70.80*
+ViT Dosovitskiy et al. (2021) 72.02* 69.72* 65.26* 74.83* 67.83* 63.40* 57.93* 70.07*
+Text-Image
+HFM Cai et al. (2019) 70.57* 66.88* 64.84* 69.05* 83.44* 80.18* 76.57* 84.15*
+Att-BERT Pan et al. (2020) 80.03* 77.04* 76.28* 77.82* 86.05* 82.92* 80.87* 85.08*
+CMGCN Liang et al. (2022) 79.83* 76.90* 75.82* 78.01* 86.54* 84.09* -* -*
+HKE Liu et al. (2022) 76.50* 72.25* 73.48* 71.07* 87.36* 72.25* 81.84* 86.48*
+DIP Wen et al. (2023) 80.59 78.23 75.52 81.14 89.59 87.17 87.76 86.58
+DynRT Tian et al. (2023) 70.37 68.55 63.02 75.15 93.59 91.93 90.30 93.62
+Multi-view CLIP Qin et al. (2023) 85.64* 84.10* 80.33* 88.24* 88.33* 85.55* 82.66* 88.65*
+G2SAM Wei et al. (2024) 79.43 78.07 72.04 85.20 90.48 88.48 87.95 89.02
+InterCLIP-MEP (Ours)
+w/o Inter ($L_{2}=1024,L_{1}=1280$) 86.05 84.81 79.83 90.45 88.75 86.31 83.73 89.05
+w/ TW ($L_{2}=128,L_{1}=1152$) 85.51 84.26 79.15 90.07 88.54 86.32 82.25 90.82
+w/ V2T ($L_{2}=640,L_{1}=1024$) 86.26 85.00 80.17 90.45 88.92 86.66 83.21 90.41
+w/ T2V ($L_{2}=1024,L_{1}=1152$) 86.72 85.61 80.20 91.80 88.83 86.37 84.02 88.84
+InterCLIP-MEP w/ RoBERTa (Ours)
+w/o Inter ($L_{2}=640,L_{1}=128$) 77.21 75.55 70.20 81.77 93.94 92.54 90.77 94.37
+w/ TW ($L_{2}=896,L_{1}=256$) 81.98 80.78 74.69 87.95 93.73 92.28 90.48 94.16
+w/ V2T ($L_{2}=640,L_{1}=512$) 76.96 75.26 69.98 81.39 93.94 92.54 90.69 94.47
+w/ T2V ($L_{2}=1024,L_{1}=384$) 82.81 81.55 75.81 88.24 93.73 92.28 90.56 94.06
 Table 2: Main results.
@@
 (2023).
-$\dagger$ indicates the results obtained by running this method on the MMSD2.0 benchmark (Qin et al.
-2023).
-Underlined values represent the best baseline for comparison.
-Bold values indicate those that surpass the baseline.
+- indicates that results are not reported.
+$\dagger$ indicates our reproduced results.
+Underlined values represent the best multi-modal baseline for comparison.
+Bold values indicate those that surpass the underlined baseline.
+$L_{2}$ for MMSD2.0 and $L_{1}$ for MMSD denote the optimal MEP memory sizes.
+Method Accuracy (%) Trainable Parameters (M) Fitting Time / Epoch (s) Inference Time (s) GPU Memory Peak (GB)
+Multi-view CLIP  85.64  165  488  51  15.59
+DIP  80.59  196  OOM  OOM  OOM
+G2SAM  79.43  116  90  13  18.32
+DynRT  70.37  25  370  26  8.03
+InterCLIP-MEP  86.72  8  55  6  6.14
+Table 3: Efficiency comparison of different methods.
+To demonstrate the efficiency of InterCLIP-MEP, we selected several recent baselines for comparison.
+The analysis was conducted using the MMSD2.0 dataset on a single NVIDIA RTX 4090 GPU with a batch size of 128.
+In the table, Fitting Time / Epoch indicates the time required for each epoch during training and validation and OOM indicates Out of Memory, referring to GPU memory overflow.
 w/o Inter w/ TW w/ V2T w/ T2V
@@
 w/o LoRA 82.44 77.73 76.42 74.37 73.31 72.22 75.13 71.79
-Table 3: Ablation study of InterCLIP-MEP. Baseline represents the results of the corresponding method without any ablation.
-Experiment
-Experimental Settings
-Datasets.
-We use the dataset from the MMSD2.0 benchmark (Qin et al.
-2023), which is an improved version of the dataset originally introduced by Cai, Cai, and Wan (2019).
-We present the statistics of the dataset in Table 1.
-Metrics.
-Following Liu, Wang, and Li (2022), we use accuracy (Acc.), precision (P), recall (R), and F1-score (F1) as metrics to evaluate the performance of our model.
+Table 4: Ablation study of InterCLIP-MEP, with Baseline denoting results without ablation.
+4.1 Experimental Settings
+Datasets and metrics.
+Following Qin et al.
+(2023), we evaluate performance on MMSD Cai et al.
+(2019) and MMSD2.0 Qin et al.
+(2023) using accuracy (Acc.), precision (P), recall (R), and F1-score (F1) as metrics.
+We present the statistics of the two datasets in Table 1.
 Baselines.
 We compare the effectiveness of the InterCLIP-MEP framework against several unimodal and multi-modal methods.
-For text modality methods, we compare with TextCNN (Kim 2014), Bi-LSTM (Graves and Schmidhuber 2005), SMSD (Xiong et al.
-2019), and RoBERTa (Liu et al.
-2019).
-For image modality methods, we compare with ResNet (He et al.
-2016) and ViT (Dosovitskiy et al.
-2021).
-For state-of-the-art multi-modal methods, we compare with the following:
-  • HFM (Cai, Cai, and Wan 2019): A hierarchical fusion model integrating text and image features.
-  • Att-BERT (Pan et al. 2020): A model using an attention mechanism to capture sarcasm incongruity.
-  • CMGCN (Liang et al. 2022): A method that constructs a cross-modal graph and employs a graph convolutional network.
-  • HKE (Liu, Wang, and Li 2022): A hierarchical framework utilizing multi-head cross-attention and graph neural networks.
-  • DIP (Wen, Jia, and Yang 2023): A dual incongruity perceiving network that extracts sarcastic information from factual and affective levels.
-  • DynRT (Tian et al. 2023): A dynamic routing transformer network for multi-modal sarcasm detection.
-  • Multi-view CLIP (Qin et al. 2023): A method adapted to CLIP for capturing multi-view sarcasm cues.
-  • G2SAM (Wei et al. 2024): A method using global graph-based semantic awareness and leveraging historical knowledge from training samples for multi-modal sarcasm detection.
-Main Results
-To validate the effectiveness of our InterCLIP-MEP framework, we conduct experiments using the original CLIP as the backbone instead of InterCLIP, referred to as InterCLIP-MEP w/o Inter.
-We compare this configuration with three interaction modes of InterCLIP: InterCLIP-MEP w/ V2T, InterCLIP-MEP w/ T2V, and InterCLIP-MEP w/ TW.
+For text modality methods, we compare with TextCNN Kim (2014), Bi-LSTM Zhou et al.
+(2016), SMSD Xiong et al.
+(2019), and RoBERTa Liu et al.
+(2019).
+For image modality methods, we compare with ResNet He et al.
+(2016) and ViT Dosovitskiy et al.
+(2021).
+We compare with state-of-the-art multi-modal methods, including HFM Cai et al.
+(2019), Att-BERT Pan et al.
+(2020), CMGCN Liang et al.
+(2022), HKE Liu et al.
+(2022), DIP Wen et al.
+(2023), DynRT Tian et al.
+(2023), Multi-view CLIP Qin et al.
+(2023), and G2SAM Wei et al.
+(2024), which employ various techniques such as hierarchical fusion, graph neural networks, and dynamic routing for multi-modal sarcasm detection.
+4.2 Main Results
+To validate the effectiveness of our InterCLIP-MEP framework, we conduct experiments using the original CLIP as the backbone instead of InterCLIP, referred to as w/o Inter.
+We compare this configuration with three interaction modes of InterCLIP: w/ V2T, w/ T2V, and w/ TW.
 For each experiment, we condition only the top four layers of the self-attention modules, with the projection dimension $d_{f}$ set to 1024.
-We set the rank $r$ of LoRA to 8, fine-tuning the self-attention module weight matrices $\mathbf{W}$, specifically $W_{k}$, $W_{v}$, and $W_{o}$.
-For the memory size $L$ of the MEP, we select the optimal size from $\mathbf{L}=\{128,256,384,512,640,768,896,1024,1152,1280\}$.
+We set the LoRA rank $r$ to 8, fine-tuning the self-attention weight matrices $\mathbf{W}$, specifically $W_{k}$, $W_{v}$, and $W_{o}$.
+For the memory size $L$, we select the optimal size from a fixed set of candidate values $\mathbf{L}$  [^2] The fixed candidate values are {128, 256, 384, 512, 640, 768, 896, 1024, 1152, 1280}..
 The main results are shown in Table 2.
-Our framework consistently outperforms or matches the performance of state-of-the-art methods, whether using InterCLIP or the original CLIP as the backbone, as shown in Table 2.
+Performance on MMSD2.0.
+For MMSD2.0, our framework consistently outperforms or matches the performance of state-of-the-art methods, whether using InterCLIP or the original CLIP as the backbone, as shown in Table 2 (InterCLIP-MEP).
 This demonstrates the effectiveness of our training strategy and MEP.
-Methods InterCLIP-MEP w/ V2T and InterCLIP-MEP w/ T2V outperform InterCLIP-MEP w/o Inter, indicating that InterCLIP more effectively captures text-image interactions.
-Additionally, InterCLIP-MEP w/ T2V outperforms InterCLIP-MEP w/ V2T, likely due to the complexity of the visual space, which poses challenges for the projection layer in mapping visual representations into the text encoder space effectively.
-We observe that InterCLIP-MEP w/ TW performs worse than InterCLIP-MEP w/o Inter, possibly because embedding representations within both encoders increases learning difficulty compared to unidirectional methods.
-In summary, InterCLIP with T2V interaction, combined with our training strategy and MEP yields the most promising results.
-Ablation Study
-To further validate the effectiveness of InterCLIP-MEP, we remove the projection module $\mathcal{F}_{p}$ and train only the classification module $\mathcal{F}_{c}$ for prediction, denoted as w/o Proj.
-To test the necessity of using LoRA (Hu et al.
-2022) for fine-tuning, we keep the rest of InterCLIP-MEP unchanged and freeze all self-attention weight matrices of InterCLIP, denoted as w/o LoRA, always selecting the optimal memory size $L$ for the MEP during inference.
+Our results show that w/ V2T and w/ T2V outperform w/o Inter, demonstrating that InterCLIP captures text-image interactions more effectively.
+Furthermore, w/ T2V achieves superior performance compared to w/ V2T, likely due to the inherent complexity of the visual space, which presents challenges for the projection layer when mapping vision representations into the text encoder space.
+In contrast, w/ TW performs worse than other configurations, possibly because embedding representations within both encoders increases the learning difficulty.
+In summary, InterCLIP with T2V interaction, combined with our training strategy and MEP, delivers the most promising results.
+These findings underscore the robustness and adaptability of our framework, establishing it as a highly effective solution for capturing nuanced text-image interactions and addressing the complexities of multi-modal sarcasm detection.
+Performance on MMSD.
+For MMSD, the RoBERTa-based text modality baseline significantly outperforms other methods due to spurious cues in the text, enabling accurate predictions solely based on textual features Qin et al.
+(2023).
+Consequently, models like DynRT, G2SAM, and DIP, which utilize RoBERTa or BERT for text feature extraction, achieve high performance on MMSD but experience a significant drop on MMSD2.0.
+To further investigate, we conduct an additional experiment, InterCLIP-MEP w/ RoBERTa, replacing the original text encoder with RoBERTa.
+While this change led to state-of-the-art performance on MMSD, it disrupted InterCLIP’s modality alignment, causing a reasonable performance drop on MMSD2.0.
+This suggests that MMSD’s text data contains spurious cues that allow models to rely heavily on the text encoder, while MMSD2.0, having been cleaned, requires more robust multi-modal capabilities.
+We further find that the w/ V2T variant consistently outperforms others in both InterCLIP-MEP and InterCLIP-MEP w/ RoBERTa experiments, underscoring the model’s tendency to overly depend on text modality.
+Efficiency Comparison.
+Our training strategy demonstrates both remarkable effectiveness and outstanding efficiency.
+To validate this, we conducted a comprehensive comparative analysis against leading state-of-the-art methods, as detailed in Table 3.
+For instance, the Multi-view CLIP method Qin et al.
+(2023) employs a multi-layer Transformer encoder for feature fusion, which, while effective, introduces a significant number of trainable parameters.
+This results in slower training and inference speeds and greater memory consumption.
+Similarly, the DIP method Wen et al.
+(2023) caches historical samples during training, which hinders its ability to support large-batch training under limited resource conditions.
+In contrast, our method operates with a batch size of 128 while utilizing a significantly smaller number of trainable parameters, which translates to notably faster training and validation cycles.
+Furthermore, by incorporating minimal parameter modifications to adapt CLIP and utilizing simple yet effective linear layers for representation fusion, our approach achieves superior inference speeds and drastically reduced memory consumption.
+These results highlight the practicality of our framework, establishing it as a benchmark for both computational efficiency and performance in multi-modal sarcasm detection.
+4.3 Analysis of InterCLIP-MEP
+To robustly validate the effectiveness of InterCLIP-MEP, we conduct comprehensive ablation studies and case studies on the more reliable MMSD2.0 benchmark, offering deeper insights into its design and performance.
+In addition, we include visualization analyses to provide an intuitive understanding of how the framework processes multi-modal sarcasm cues.
+Ablation study.
+We remove the projection module $\mathcal{F}_{p}$ and train only the classification module $\mathcal{F}_{c}$ for prediction, denoted as w/o Proj.
+To test the necessity of using LoRA Hu et al.
+(2022) for fine-tuning, we keep the rest of InterCLIP-MEP unchanged and freeze all self-attention weights of InterCLIP, denoted as w/o LoRA, always selecting the optimal memory size $L$ for the MEP during inference.
 To evaluate the effectiveness of the MEP, we train both $\mathcal{F}_{p}$ and $\mathcal{F}_{c}$ but use only $\mathcal{F}_{c}$ during inference, denoted as w/o MEP.
 Table 4 reports all results.
-All variants show performance declines compared to the baseline, demonstrating the importance of each module in the InterCLIP-MEP framework.
-For methods InterCLIP-MEP w/ TW and InterCLIP-MEP w/o Inter, the w/o MEP variant performed worse than the w/o Proj variant.
-However, for methods InterCLIP-MEP w/ T2V and InterCLIP-MEP w/ V2T, the w/o MEP variant performs better than the w/o Proj variant.
+All variants show performance declines compared to the baseline, demonstrating the importance of each module in InterCLIP-MEP.
+For w/ TW and w/o Inter, the w/o MEP variant performed worse than the w/o Proj variant.
+However, for w/ T2V and w/ V2T, the w/o MEP variant performs better than the w/o Proj variant.
 This suggests that backbones with strong image-text interaction capabilities benefit from training the classification module $\mathcal{F}_{c}$ along with the projection module $\mathcal{F}_{p}$, even without using MEP during inference.
 We also find that not using LoRA to fine-tune the self-attention modules results in significant performance loss, indicating that the original CLIP’s vanilla space is not directly suitable for the sarcasm detection task.
-Hyperparameter Study
-In this section, we further investigate the method using Interactive-CLIP with T2V interaction as the backbone.
+Case study.
+As shown in Figure 4, we select three examples to further demonstrate the robustness of InterCLIP-MEP.
+We observe that direct predictions through the classification module $\mathcal{F}_{c}$ result in high prediction entropy and incorrect outcomes.
+However, MEP effectively mitigates the issue for cases where $\mathcal{F}_{c}$ fails to correctly identify the results by using the historical knowledge of the test samples.
+This integration ensures that even in complex situations, the model maintains a high level of accuracy.
+Figure 4: Case study of InterCLIP-MEP.
+In the figure, GT represents the labels annotated by human experts, MEP represents the labels predicted by the memory-enhanced predictor, and $\mathcal{F}_{c}$ represents the labels provided by the classification module.
+Figure 5: Visual comparison between InterCLIP and original CLIP.
+Both InterCLIP and the original CLIP were fine-tuned using the same training set and identical parameters.
+The key distinction is that the original CLIP does not incorporate interaction.
+Visulization.
+To further validate that InterCLIP is capable of more effectively capturing the interactive information between text and images compared to original CLIP, thereby aiding in the detection of sarcasm cues, we use GradCAM Selvaraju et al.
+(2017) to visualize the areas of focus during the inference process of the visual model in Figure 5.
+We observe that in the first example, which complains about the unpleasant odor caused by the sea lions, InterCLIP focuses more accurately on the location of the sea lions compared to the original CLIP.
+In the second example, which complains about traffic congestion, InterCLIP correctly focuses on the distribution of cars on the road, whereas the original CLIP’s focus is scattered.
+5 Conclusion
+In this paper, we propose InterCLIP-MEP, a novel framework for multi-modal sarcasm detection that directly addresses the challenges of modeling nuanced text-image interactions and managing prediction uncertainty.
+We design Interactive CLIP (InterCLIP) to embed cross-modal information within text and image encoders, enabling a deeper understanding of sarcasm cues.
+Additionally, we develop a Memory-Enhanced Predictor (MEP) to dynamically leverage historical sample knowledge, making our inference process more robust and adaptive.
+Through extensive experiments on MMSD and MMSD2.0 benchmarks, we demonstrate that InterCLIP-MEP achieves state-of-the-art performance while significantly reducing computational costs.
+By requiring fewer trainable parameters and less GPU memory, our method offers a lightweight, efficient, and scalable solution, setting a new benchmark for multi-modal sarcasm detection.
+Limitations
+While InterCLIP-MEP delivers strong performance, there remain areas for further refinement.
+For instance, the framework could benefit from additional techniques to better capture sarcasm that heavily relies on subtle cultural or highly specific contextual cues.
+Moreover, extending its application to more diverse and less structured datasets could be an interesting direction for future work, further broadening its practical applicability.
+Ethical Considerations
+This work focuses on advancing multi-modal sarcasm detection to improve understanding of complex communication in online content.
+While the proposed framework enhances detection accuracy and efficiency, potential misuse must be considered.
+Automated sarcasm detection could inadvertently amplify biases present in training data or be deployed for unethical purposes, such as targeted content moderation or surveillance.
+To mitigate these risks, we encourage the responsible use of this technology and emphasize the importance of using diverse and unbiased datasets during training to minimize unintended consequences.
+Furthermore, this research strictly adheres to ethical guidelines for data collection and usage.
+Appendix A Implementation Details
+The model training and testing were conducted using PyTorch Lightning [^3] https://lightning.ai/.
+InterCLIP was constructed by leveraging the Transformers library Wolf et al.
+(2020).
+For the MMSD2.0 experiments, the initial weights for InterCLIP are based on clip-vit-base-patch32444https://huggingface.co/openai/clip-vit-base-patch32.
+For the MMSD experiments, we utilized the roberta-ViT-B-32 model architecture provided by OpenCLIP [^5] https://github.com/mlfoundations/open_clip, with the pre-trained checkpoint laion2b_s12b_b32k [^6] https://huggingface.co/laion/CLIP-ViT-B-32-roberta-base-laion2B-s12B-b32k.
+Custom scripts were developed to adapt its format to the Transformers library, ensuring compatibility with our framework.
+The model parameters were optimized using AdamW Loshchilov and Hutter (2019), with a learning rate set to 1e-4 for the LoRA fine-tuning modules and 5e-4 for other trainable modules.
+A cosine annealing scheduler with warmup was employed to dynamically adjust the learning rate, where the warmup steps constituted the first 20% of the total optimization steps, and the minimum learning rate was set to 1% of the initial rate.
+For the modules $\mathcal{G}_{t/v}$, $\mathcal{F}_{t/v}$, $\mathcal{F}_{c}$, and $\mathcal{F}_{p}$, simple multi-layer perceptrons (MLPs) were utilized.
+The training processing was performed with a batch size of 64 for 3 epochs.
+All experiments were run on a machine equipped with an NVIDIA RTX 4090 GPU.
+Appendix B Experimental Details
+B.1 Hyperparameter details
+We summarize the hyperparameters involved in InterCLIP-MEP and their descriptions in Table 5.
+The hyperparameter settings for obtaining the main results in the paper are summarized in Table 6.
+For other baseline methods, we follow the optimal hyperparameter settings they reported.
+Parameter Description
+$r$ The rank of LoRA, determining the dimension of the low-rank update matrices.
+$\mathbf{W}$ The weight matrices in the self-attention module fine-tuned using LoRA, specifically targeting combinations of $W_{\{q,k,v,o\}}$.
+top-$n$ The number of top self-attention layers conditioned during fine-tuning.
+$d_{f}$ The dimensionality of the latent space for the projected features.
+$\mathbf{L}$ The configurable range of memory sizes maintained by the Memory-Enhanced Predictor (MEP).
+Table 5: Summary of hyperparameters.
+Param. Value
+For trainer
+epoch 3
+batch_size 64
+lr 5e-4
+lora_lr 1e-4
+warmup_ratio 0.2
+min_lr_rate 0.01
+For our model
+$r$ 8
+top-$n$ 4
+$d_{f}$ 1024
+$\mathbf{W}$ $W_{k},W_{v},W_{o}$
+$\mathbf{L}$128, 256, 384, 512, 640, 768, 896, 1024, 1152, 1280}
+Table 6: Hyperparameter settings.
+Figure 6: Hyperparameter study curves for w/ T2V. Panel (d) compares results with those from using only the classification module $\mathcal{F}_{c}$ for prediction.
+B.2 Hyperparameter study
+We further investigate the method using Interactive-CLIP with T2V interaction as the backbone.
 Keeping the other hyperparameters constant, we condition different top-$n$ layers of the self-attention modules.
-We also study the impact of different projection dimensions $d_{f}$, different LoRA ranks $r$, and different memory sizes $L$ on the InterCLIP-MEP w/ T2V method.
+We also study the impact of different projection dimensions $d_{f}$, different LoRA ranks $r$, and different memory sizes $L$ on the w/ T2V method.
 We present all results in Figure 6.
-Figure 4: Hyperparameter study curves for InterCLIP-MEP w/ T2V. Panel (d) compares results with those from using only the classification module $\mathcal{F}_{c}$ for prediction.
 Figure 6(a) shows that conditioning the top four self-attention layers yields the best results.
@@
 Figure 6(c) indicates the projection dimension is best at 256 or 1024.
-Figure 4(d) reveals that a memory size of 640 in MEP outperforms others, confirming the value of historical sample knowledge.
-Figure 5: Case study of InterCLIP-MEP.
-Figure 6: Visual comparison between InterCLIP and CLIP.
-Case Study and Visualization
-As shown in Figure 5, we select three examples to further demonstrate the robustness of InterCLIP-MEP.
-We observe that direct predictions through the classification module $\mathcal{F}_{c}$ result in high prediction entropy and incorrect outcomes.
-However, MEP effectively mitigates the issue for cases where $\mathcal{F}_{c}$ fails to correctly identify the results by using the historical knowledge of the test samples.
-To further validate that InterCLIP can more effectively capture the interactive information between text and images compared to original CLIP, thereby aiding in the detection of sarcasm cues, we visualize the areas of focus during the inference process of the visual model in Figure 6.
-We observe that in the first example, which complains about the unpleasant odor caused by the sea lions, InterCLIP focuses more accurately on the location of the sea lions compared to the original CLIP.
-In the second example, which complains about traffic congestion, InterCLIP correctly focuses on the distribution of cars on the road, whereas the original CLIP’s focus is scattered.
-Conclusion
-In this paper, we propose a novel framework for robust multi-modal sarcasm detection, termed InterCLIP-MEP.
-We introduce an Interactive CLIP (InterCLIP) as the backbone for extracting sample representations.
-Unlike the original CLIP, InterCLIP embeds representations from different modalities into the encoder, effectively capturing the intricate interactions between text and image.
-For inference, we introduce a Memory-Enhanced Predictor (MEP) that leverages historical knowledge of valuable test samples to enhance inference, providing a more robust and reliable multi-modal sarcasm detection.
-Our experimental results demonstrate that InterCLIP-MEP achieves state-of-the-art performance on the MMSD2.0 benchmark.
+Figure 6(d) reveals that a memory size of 640 in MEP outperforms the others, confirming the value of historical sample knowledge.
+B.3 Empirical study of self-attention fine-tuning and interaction modes
+Keeping the other hyperparameters constant as shown in Table 6, we fine-tune all possible weight matrices $\mathbf{W}$ and employ different interaction modes of InterCLIP as the backbone.
+We consistently select the optimal memory size from $\mathbf{L}$ for MEP.
+We calculate the average metrics for different methods and different weight matrices.
+The results are presented in Table 7 and Table 8.
+We observe that fine-tuning the weight matrices $W_{k},W_{v},W_{o}$ and using the T2V interaction mode of InterCLIP are the best choices for InterCLIP-MEP.
+$\mathbf{W}$ Mean Acc. (%) Mean F1 (%)
+$W_{q}$ 85.14 83.99
+$W_{k}$ 85.09 84.00
+$W_{v}$ 85.34 84.24
+$W_{o}$ 85.39 84.23
+$W_{q},W_{k}$ 85.36 84.16
+$W_{q},W_{v}$ 85.63 84.40
+$W_{q},W_{o}$ 85.73 84.49
+$W_{k},W_{o}$ 85.67 84.43
+$W_{v},W_{o}$ 85.73 84.54
+$W_{k},W_{v}$ 85.70 84.51
+$W_{q},W_{k},W_{o}$ 85.92 84.63
+$W_{q},W_{v},W_{o}$ 85.87 84.60
+$W_{q},W_{k},W_{v}$ 86.05 84.75
+$W_{k},W_{v},W_{o}$ 86.14 84.92
+$W_{q},W_{k},W_{v},W_{o}$ 85.82 84.55
+Table 7: Average results of fine-tuning different weight matrices $\mathbf{W}$ across four baseline methods.
+Method Mean Acc. (%) Mean F1 (%)
+w/o Inter 85.49 84.32
+w/ TW 85.66 84.44
+w/ V2T 85.62 84.41
+w/ T2V 85.78 84.55
+Table 8: Average results of four baseline methods for fine-tuning different weight matrices $\mathbf{W}$.
+Appendix C More Visual Examples
+Figure 7 presents additional examples illustrating the focus differences between InterCLIP and CLIP.
+These visualizations highlight InterCLIP’s improved ability to capture sarcasm-related cues by focusing on relevant areas in the images.
+Figure 7: Additional visual examples showcasing InterCLIP’s improved focus on sarcasm-related visual cues compared to CLIP.
+Appendix D Extended Experiments
+To further verify the performance of our framework, we conducted an extended set of experiments.
+Sarcastic Non-sarcastic All
+Train 4,014 46,265 50,279
+Validation 1,125 13,097 14,222
+Test 555 6,772 7,327
+Table 9: Statistics of DocMSU.
+Method Acc. F1 P R
+BERT-base* 87.12 86.51 77.61 70.37
+Swin-Transformer* 74.83 61.51 67.57 56.45
+CMGCN* 88.12 75.23 78.11 72.55
+CLIP* 96.19 77.62 78.99 76.30
+ViLT* 93.15 41.44 69.03 29.61
+Du et al. (2024)* 97.83 87.25 81.20 94.27
+InterCLIP-MEP (Ours)
+w/o Inter 97.84 87.48 78.08 99.45
+w/ TW 97.79 87.24 77.48 99.81
+w/ V2T 97.83 87.45 77.81 99.82
+w/ T2V 97.67 86.65 76.45 99.99
+Table 10: Results of the extended experiments.
+Underline results denote the compared SOTA baseline, boldface highlights results that surpass the baseline, and * indicates results sourced from Du et al.
+(2024).
+D.1 Benchmark
+DocMSU Du et al.
+(2024) is a recently introduced multi-modal sarcasm benchmark designed specifically for long-text analysis.
+It facilitates the evaluation of multi-modal sarcasm comprehension as well as detection tasks.
+In this study, we concentrate on the multi-modal sarcasm detection task.
+The benchmark statistics can be found in Table 9.
+D.2 Baselines
+We follow the evaluation protocol of Du et al.
+(2024).
+We compare against unimodal baselines: BERT-base (text-only) Devlin (2018) and Swin Transformer (image-only) Liu et al.
+(2021).
+For multi-modal approaches, we include CLIP Radford et al.
+(2021), Vision-and-Language Transformer (ViLT) Kim et al.
+(2021), and CMGCN Liang et al.
+(2022).
+The method proposed by Du et al.
+(2024) is taken as the state-of-the-art baseline.
+D.3 Results
+Our InterCLIP-MEP framework demonstrates strong performance across various configurations, as shown in Table 10.
+In particular, the w/o Inter and w/ V2T variants consistently achieve higher F1 scores compared to the baselines, thereby showcasing their robustness in handling multi-modal sarcasm detection tasks.
+Notably, all variants achieve nearly perfect recall, highlighting their outstanding capability in accurately identifying sarcasm across diverse datasets.
+The w/o Inter variant also achieves the highest accuracy, further demonstrating its effectiveness and precision.
+Overall, the comprehensive results in Table 10 affirm the unparalleled effectiveness of InterCLIP’s modality interaction mechanism and our proposed memory-enhanced predictor (MEP), particularly in surpassing existing state-of-the-art methods in both accuracy and F1 score, thus setting a new benchmark in the field.
+Appendix E Other Related Works
+E.1 CLIP adaptation
+The Contrastive Language-Image Pretraining (CLIP) model Radford et al.
+(2021) excels in vision-language tasks.
+Adapting CLIP for specific domains has shown substantial improvements, as demonstrated by Li et al.
+(2022) for phrase localization, Liang et al.
+(2023) for open-vocabulary semantic segmentation, and Wang et al.
+(2023) for action recognition.
+In this work, inspired by Ganz et al.
+(2024), we conditionally enhance both the text and vision encoders of CLIP, making it more effective in capturing the interplay between text and images to identify multi-modal sarcasm cues.
+Unlike Ganz et al.
+(2024), who focused solely on embedding text representations into the vision encoder, we also explore embedding image representations into the text encoder.
+Furthermore, their approach is limited to general classification tasks and does not address the complexities of multi-modal sarcasm detection.
+E.2 Memory-enhanced prediction
+Inspired by cognitive science Stokes (2015); Baddeley (2000), memory has been introduced to enhance neural networks Weston et al.
+(2014); Sukhbaatar et al.
+(2015).
+Several studies Wu et al.
+(2018); Wen et al.
+(2023) have used memory mechanisms to improve model training, and some Zhang et al.
+(2024); Wei et al.
+(2024) leverage memory to store historical knowledge, enhancing prediction accuracy.
+In this work, we introduce a memory-enhanced predictor for multi-modal sarcasm detection.
+In contrast to other methods, our memory dynamically updates during testing, utilizing relevant historical information for improved accuracy and robustness.
+Appendix F List of Symbols
+In Table 11, we have listed the main symbols used in the paper and their descriptions.
+Symbol Description
+$T$ $T$ denotes a short text.
+$I$ $I$ represents an image.
+$\mathcal{P}$ $\mathcal{P}$ denotes a text-image pair $(T,I)$.
+$\mathcal{T}$ $\mathcal{T}$ denotes CLIP’s text encoder.
+$\mathcal{V}$ $\mathcal{V}$ denotes CLIP’s vision encoder.
+$\mathbf{F}$ $\mathbf{F}$ represents the final layer representations encoded by either the text or vision encoder, with text representations as $\mathbf{F}_{t}$ and image representations as $\mathbf{F}_{v}$.
+$\tilde{\mathbf{F}}$ $\tilde{\mathbf{F}}$ represents the final layer representations encoded by either the text or vision encoder after embedding representations from another modality, with text representations as $\tilde{\mathbf{F}}_{t}$ and image representations as $\tilde{\mathbf{F}}_{v}$.
+$\mathbf{H}$ $\mathbf{H}$ represents the input representations for each sub-attention layer in the text or vision encoders.
+Each layer’s input comes from the output of the previous layer, denoted $\mathbf{H}_{t}$ for the text encoder and $\mathbf{H}_{v}$ for the vision encoder.
+$\mathcal{F}_{t/v}$ $\mathcal{F}_{t/v}$ denotes the adapting projection layer in the text or vision encoders used to project the embedded representations of the other modality into the current encoder space.
+It is denoted as $\mathcal{F}_{t}$ in the text encoder and $\mathcal{F}_{v}$ in the vision encoder.
+$\mathbf{F}^{{}^{\prime}}$ $\mathbf{F}^{\prime}$ represents the representations projected into the corresponding encoder space.
+For example, embedding visual representations $\mathbf{F}_{v}$ in the text encoder and projecting it through $\mathcal{F}_{t}$ results in $\mathbf{F}^{\prime}_{v}$.
+$\mathbf{H}^{{}^{\prime}}$ $\mathbf{H}^{\prime}$ represents the representations after embedding another modality’s representations and processing them through a self-attention layer.
+$\mathcal{H}_{t/v}$ $\mathcal{H}_{t/v}$ denotes the projection module in the self-attention layer used to transform the output of the self-attention module, denoted $\mathcal{H}_{t}$ for the text encoder and $\mathcal{H}_{v}$ for the vision encoder.
+$\mathcal{G}_{t/v}$ $\mathcal{G}_{t/v}$ denotes the projection module in the self-attention layer that has embedded representations from another modality, used to jointly transform the output representation in combination with $\mathcal{H}_{t/v}$.
+$\mathbf{H}^{{}^{\prime\prime}}$ $\mathbf{H}^{\prime\prime}$ represents the final representations in the self-attention layer.
+$\tilde{h}^{f}$ $\tilde{h}^{f}$ denotes the final fused feature obtained from a sample.
+$\mathcal{F}_{c}$ $\mathcal{F}_{c}$ denotes the classification module used to assign pseudo-labels to samples.
+$\mathcal{F}_{p}$ $\mathcal{F}_{p}$ denotes the projection module used to project samples into a latent space.
+$\hat{h}^{f}$ $\hat{h}^{f}$ represents the feature of a sample’s fused feature after transformation by $\mathcal{F}_{p}$ and L2 normalization.
+Table 11: List of symbols

### v7

@@
 InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection
-Abstract
-Sarcasm in social media, often expressed through text-image combinations, poses challenges for sentiment analysis and intention mining.
-Current multi-modal sarcasm detection methods have been demonstrated to overly rely on spurious cues within the textual modality, revealing a limited ability to genuinely identify sarcasm through nuanced text-image interactions.
-To solve this problem, we propose InterCLIP-MEP, which introduces Interactive CLIP (InterCLIP) with an efficient training strategy to extract enriched text-image representations by embedding cross-modal information directly into each encoder.
-Additionally, we design a Memory-Enhanced Predictor (MEP) with a dynamic dual-channel memory that stores valuable test sample knowledge during inference, acting as a non-parametric classifier for robust sarcasm recognition.
-Experiments on two benchmarks demonstrate that InterCLIP-MEP achieves state-of-the-art performance, with significant accuracy and F1 score improvements on MMSD and MMSD2.0.
-Our code is available at https://github.com/CoderChen01/InterCLIP-MEP.
-InterCLIP-MEP: Interactive CLIP and Memory-Enhanced Predictor for Multi-modal Sarcasm Detection
-Junjie Chen1,3thanks: Junjie is a research assistant at the EPIC Lab of Shanghai Jiao Tong University, working remotely with Linfeng Zhang.
-Hang Yu2thanks: Co-corresponding author.
-Subin Huang3thanks: Corresponding author.
-Sanmin Liu3    Linfeng Zhang1 1Shanghai Jiao Tong University, China    2Shanghai University, China 3Anhui Polytechnic University, China jorji.chen@gmail.com,  yuhang@shu.edu.cn,  subinhuang@ahpu.edu.cn, sanmin.liu@ahpu.edu.cn,  zhanglinfeng@sjtu.edu.cn
-1 Introduction
-Sarcasm, with its subtlety and complexity, plays a key role in communication by conveying irony, mockery, or hidden meanings Muecke (1982); Gibbs and O’Brien (1991); Gibbs and Colston (2007).
-Automatically detecting sarcasm in text has become an important research area, supporting tasks like sentiment analysis and intent mining Pang et al.
-(2008); Tsur et al.
-(2010); Bouazizi and Ohtsuki (2015).
-Figure 1: An overview of the shortcomings of existing multi-modal sarcasm detection pipelines.
-In panels (a) and (b), we present two main multi-modal sarcasm detection pipelines, with shortcomings indicated by a red question mark.
-In panel (c), we visually show an example of multi-modal sarcasm cues correctly or incorrectly recognized in a multi-modal sarcasm sample.
-With the rise of social media platforms like Twitter and Reddit, users often use text-image combinations to express their messages.
-As a result, multi-modal sarcasm detection is increasingly important, posing challenges in understanding the complex relationship between textual and visual cues to identify sarcasm.
-As shown in Figures 1(a) and 1(b), many methods rely on dual unimodal pre-trained encoders, such as ViTDosovitskiy et al.
-(2021) and BERT Devlin et al.
-(2019), as the backbone for encoding text-image pairs, followed by specific feature fusion Xu et al.
-(2020); Pan et al.
-(2020); Liang et al.
-(2021, 2022); Wen et al.
-(2023); Tian et al.
-(2023); Wei et al.
-(2024).
-However, this approach may not capture multi-modal sarcasm cues as effectively as multi-modal pre-trained models like CLIP Radford et al.
-(2021).
-In Figure 1(a), the use of a learnable classification head to predict labels from fused representations is common but often associated with high predictive entropy and significant uncertainty.
+Abstract.
+[…]
+Multi-modal Sarcasm Detection, Multi-modal Interaction, Cross-modal Representations, Memory-enhanced Predictor, Parameter-efficient Fine-tuning
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Multimedia information systemsccs: Computing methodologies Information extractionccs: Information systems Sentiment analysis
+1. Introduction
+Sarcasm, characterized by its nuanced and intricate nature, serves as a vital mechanism in communication, often employed to convey irony, mockery, or implicit meanings (Gibbs1991PsychologicalAO; tsur2010icwsm; JingSOJN23; LiuZS24).
+The automated identification of sarcasm in textual data has evolved into a significant research focus, offering substantial support for tasks such as sentiment analysis and intent mining (pang2008opinion; JoshiTPBC16; JoshiBC17).
+With the growing prevalence of social media platforms such as Twitter and Reddit, users increasingly rely on the fusion of textual and visual elements to articulate their messages.
+Consequently, multi-modal sarcasm detection has gained considerable prominence, presenting challenges in deciphering the intricate interplay between textual and visual cues to accurately identify sarcastic content (WangYJMXL24; ZhuangLZGZM25).
+Despite significant progress achieved by prior studies (xu2020reasoning; pan-etal-2020-modeling; liang2021multi; Liang2022MultiModalSD; 10.24963/ijcai.2024/887) in this domain, Qin et al.
+(qin2023mmsd2) have identified the presence of spurious cues within the MMSD benchmark (Cai2019MultiModalSD) utilized in these works.
+These cues contribute to the development of biased models, thereby overestimating their effectiveness in detecting sarcasm cues from multi-modal data.
+Through the introduction of MMSD2.0, a refined benchmark that eliminates spurious cues and rectifies mislabeled samples, Qin et al.
+(qin2023mmsd2) reveal a significant performance degradation of existing methods when evaluated on MMSD2.0.
+Although their proposed method, which leverages Multi-view CLIP for reliable multi-modal sarcasm detection, shows promise, it still exhibits limited capability in accurately discerning sarcasm through subtle text-image interactions.
+Figure 1.
+An overview of the limitations inherent in existing multi-modal sarcasm detection frameworks.
+Panels (a) and (b) illustrate two predominant multi-modal sarcasm detection pipelines, with their respective shortcomings highlighted by red question marks.
+Panel (c) provides a visual representation of multi-modal sarcasm cues, demonstrating instances where such cues are either accurately identified or misinterpreted in a multi-modal sarcasm sample.
+As depicted in Figures 1(a) and 1(b), numerous approaches utilize dual unimodal pre-trained encoders, such as ViT (dosovitskiy2020image) and BERT (devlin2019bert), as the backbone for encoding text-image pairs, followed by specialized feature fusion techniques (xu2020reasoning; pan-etal-2020-modeling; liang2021multi; Liang2022MultiModalSD; wen2023dip; Tian2023DynamicRT; wei2024g2sam; Wang2025AdaSFFuse).
+However, such methodologies may fail to effectively capture multi-modal sarcasm cues compared to multi-modal pre-trained models like CLIP (radford2021learning).
+In Figure 1(a), the adoption of a learnable classification head to predict labels from fused representations is prevalent but often accompanied by high predictive entropy and substantial uncertainty.
+The work by (wei2024g2sam) introduced the construction of a static semantic space utilizing historical training samples, where robust predictions during inference are achieved through KNN-based querying and voting, as illustrated in Figure 1(b).
+Nevertheless, despite CLIP’s demonstrated effectiveness as a text-image encoder for multi-modal sarcasm detection (qin2023mmsd2), it remains challenged in capturing multi-modal sarcasm cues due to the inconsistency of sarcasm, which contradicts CLIP’s alignment of text and image.
+Additionally, reliance on a static semantic space for inference is inadequate to address the dynamic nature of evolving sample distributions.
+Indeed, Qin et al.
+(qin2023mmsd2) have highlighted that many models depend on spurious cues within the MMSD benchmark (Cai2019MultiModalSD), resulting in biased outcomes.
+In this paper, building on these insights, we propose InterCLIP-MEP, a novel framework designed to enhance the reliability of multi-modal sarcasm detection.
+Our InterCLIP-MEP introduces a refined variant of CLIP, Interactive CLIP (InterCLIP), as the backbone to enhance sample encoding for detecting multi-modal sarcasm.
+InterCLIP directly embeds cross-modal representations into text and vision encoders, thereby enhancing the understanding of multi-modal sarcasm cues (as illustrated in Figure 2, on the left).
+To complement this, we further design a Memory-Enhanced Predictor (MEP), which dynamically leverages historical test sample features to construct a more adaptive and reliable non-parametric classifier (Figure 2, right).
+Importantly, the design of MEP is conceptually motivated by both cognitive science and machine learning theories.
+From the cognitive perspective, human working memory is known to have limited capacity and tends to preferentially retain information with higher certainty (Baddeley2000TheEB; anderson1991reflections).
+Inspired by this mechanism, MEP employs an entropy-based update strategy to selectively preserve low-uncertainty features, simulating human-like selective retention.
+From the machine learning perspective, entropy is a widely recognized measure of predictive uncertainty, and uncertainty-aware sample selection has been shown effective in active learning and memory-based inference (toneva2018empirical; settles2009active).
+Beyond these general motivations, the characteristics of sarcasm further justify our design: sarcastic expressions in social media typically fall into a limited set of semantic patterns (e.g., exaggeration, contrast, or ironic praise), and humans often rely on prior experiences or historical contexts to correctly interpret them (attardo2000irony; gibbs1986psycholinguistics; gibbs2007irony; bamman2015contextualized; wallace-etal-2014-humans).
+By maintaining low-entropy representations in memory, MEP effectively acts as a denoising mechanism, ensuring more robust predictions during inference.
+Collectively, these components constitute the proposed framework, InterCLIP-MEP.
+Additionally, InterCLIP-MEP incorporates an efficient training strategy that fine-tunes cross-modal interactions through a lightweight adaptation mechanism, ensuring computational efficiency while achieving state-of-the-art performance (Figure 2, left).
+In summary, the key contributions of this work are as follows:
+  • The proposed framework, InterCLIP-MEP, introduces a novel approach for multi-modal sarcasm detection.
+It integrates Interactive CLIP (InterCLIP) to enhance text-image interaction encoding and incorporates a Memory-Enhanced Predictor (MEP) to achieve more robust and reliable sarcasm predictions.
+[…]
+  • Extensive experiments conducted on the MMSD, MMSD2.0 and DocMSU benchmarks demonstrate that InterCLIP-MEP achieves significant improvements, outperforming state-of-the-art methods with an increase of 1.08% in accuracy and 1.51% in F1 score, particularly on the MMSD2.0 dataset.
+The remainder of this paper is structured as follows.
+Section 2 reviews related studies and discusses the limitations of existing methodologies.
+Section 3 introduces the proposed InterCLIP-MEP framework, including the Interactive CLIP backbone, efficient training strategy, and Memory-Enhanced Predictor.
+Section 4 presents the main experiments, which provide both quantitative and qualitative evaluations as well as an efficiency analysis to demonstrate the effectiveness of the proposed method.
+Section 5 reports the extended experiments, covering four aspects: performance on a document-level benchmark, evaluation under low-resource settings, error propagation robustness study, and robustness analysis under distributional shifts, to comprehensively verify the generalization and stability of InterCLIP-MEP.
+Section 6 discusses the limitations of the current study and outlines potential directions for future research.
+Finally, Section 7 concludes the paper.
+2. Related Work
+2.1. Multi-modal Sarcasm Detection
+Early research in sarcasm detection predominantly focused on textual data (bouazizi2015sarcasm; amir2016modelling; baziotis2018ntua).
+With the proliferation of social media platforms, identifying sarcasm in text-image pairs has emerged as a more complex challenge, prompting the advancement of multi-modal methodologies.
+Among the pioneering efforts, Schifanella et al.
+(schifanella2016detecting) explored multi-modal social media posts to detect sarcasm cues.
+Expanding on this work, Cai et al.
+(Cai2019MultiModalSD) proposed the MMSD benchmark, showcasing the efficacy of a hierarchical fusion model that incorporates image features.
+This benchmark has since served as a cornerstone for multi-modal sarcasm detection, motivating a series of subsequent studies (xu2020reasoning; pan-etal-2020-modeling; liang2021multi; Liang2022MultiModalSD; Liu2022TowardsMS; qin2023mmsd2; wen2023dip; Tian2023DynamicRT; wei2024g2sam; 10.24963/ijcai.2024/887; ding2022multi; 10477507; zhuang2025multi).
+However, subsequent studies revealed that the MMSD benchmark contained spurious cues, which could introduce bias into model predictions (qin2023mmsd2).
+To address this issue, Qin et al.
+(qin2023mmsd2) developed the MMSD2.0 benchmark, which eliminates these cues and rectifies mislabeled samples.
+Re-evaluations conducted on MMSD2.0 demonstrated significant performance declines in existing models, underscoring the necessity for more robust methodologies.
+[…]
+2.2. CLIP Adaptation
+The Contrastive Language-Image Pretraining (CLIP) model (radford2021learning) exhibits remarkable capabilities in cross-modal alignment for vision-language understanding tasks.
+Recent progress in domain-specific adaptations of CLIP has led to substantial performance enhancements across diverse applications, including phrase localization (li2022adapting), open-vocabulary semantic segmentation (Liang2023Open), and action recognition (wang2023seeing).
+These achievements highlight CLIP’s adaptability in addressing specialized visual-linguistic challenges.
+Closely related to our work, Ganz et al.
+(ganz2024question) proposed the Question-Aware Vision Transformer (QA-ViT), which augments frozen vision encoders with question embeddings to produce question-aware visual representations.
+However, QA-ViT was designed for general multi-modal reasoning tasks such as visual question answering rather than sarcasm detection.
+It only explores the text-to-vision direction, leaving the reverse integration of visual cues into textual encoding unexplored.
+More importantly, it does not address the pragmatic and incongruity-based nature of sarcasm, where subtle mismatches between modalities must be captured.
+In parallel, Qin et al.
+(qin2023mmsd2) introduced Multi-view CLIP as the first CLIP-based approach for multi-modal sarcasm detection, where frozen CLIP encoders are used to extract text-only, image-only, and joint representations, and the outputs are fused via a transformer.
+While pioneering, this design treats CLIP merely as a static feature extractor and lacks adaptive cross-modal interaction within the encoders.
+In contrast, our Interactive CLIP (InterCLIP) explicitly conditions both encoders on cross-modal signals through bidirectional interaction and lightweight LoRA tuning, enabling a richer alignment of subtle sarcastic cues.
+When combined with our Memory-Enhanced Predictor (MEP), which adaptively incorporates reliable exemplars at inference time, our framework directly targets the unique challenges of sarcasm detection and provides a principled, task-specific advancement over both QA-ViT and Multi-view CLIP.
+2.3. Memory-enhanced Prediction
+Inspired by the insights from cognitive science (stokes2015activity; Baddeley2000TheEB; anderson1991reflections; gibbs2007irony), the concept of memory has been ingeniously integrated into neural networks to enhance their capabilities (weston2014memory; sukhbaatar2015end).
+This integration has opened up new avenues for improving model training, as evidenced by several pioneering studies (wu2018unsupervised; wen2023dip; zhu2020label).
+For instance, Wen et al.
+(wen2023dip) proposed DIP, which maintains semantic distributions in memory banks to stabilize training and better model incongruity, while Zhu and Yang (zhu2020label) introduced a label-independent memory (LIM) to cache unlabeled video features for robust prototype construction in semi-supervised few-shot learning.
+Despite differences in domains, both DIP and LIM adopt training-time memory to act as stabilizers or prototype aggregators, and their memories vanish once the model enters the inference stage.
+In contrast, our Memory-Enhanced Predictor (MEP) introduces a fundamentally different usage of memory: it is a test-time, entropy-aware predictor rather than a training aid.
+Instead of serving as a regularizer during optimization, MEP dynamically maintains a compact buffer of reliable test exemplars and integrates them into the prediction process through cosine matching.
+This makes memory a core inference component rather than an auxiliary training tool.
+Furthermore, some recent works (zhang2024dual; wei2024g2sam) have also explored inference-time memories.
+Zhang et al.
+(zhang2024dual) proposed Dual Memory Networks for general vision-language adaptation, where a static memory caches training features and a dynamic memory accumulates test representations, with cross-attention used to derive adaptive classifiers.
 Wei et al.
-(2024) pioneered the construction of a static semantic space using historical training samples, where more robust predictions are obtained during inference through KNN-based querying and voting, as illustrated in Figure 1(b).
-However, while CLIP has been proven to be effective in serving as a text-image encoder for multi-modal sarcasm detection Qin et al.
-(2023), it still struggles to capture multi-modal sarcasm cues due to the inherent inconsistency of sarcasm, which conflicts with CLIP’s direct alignment of text and image.
-Furthermore, relying on a static semantic space for inference is ill-suited to handle the dynamic nature of evolving sample distributions.
-In fact, Qin et al.
-(2023) have shown that many models rely on spurious cues in the MMSD benchmark Cai et al.
-(2019), leading to biased results.
-Building on the limitations of prior multi-modal sarcasm detection approaches, we propose Interactive CLIP (InterCLIP) as the backbone, embedding cross-modal representations directly into text and vision encoders to enhance the understanding of multi-modal sarcasm cues (Figure 2, left).
-To complement this, we design a Memory-Enhanced Predictor (MEP) that dynamically utilizes historical test sample features to create a more adaptive and reliable non-parametric classifier for final predictions (Figure 2, right).
-Together, these components form the proposed framework, InterCLIP-MEP.
-Furthermore, InterCLIP-MEP employs an efficient training strategy that fine-tunes cross-modal interactions through a lightweight adaptation mechanism, ensuring computational efficiency while delivering state-of-the-art performance (Figure 2, left).
-Overall, our contributions are as follows:
-  • We introduce InterCLIP-MEP [^1] Our code is available in the supplementary material., a novel framework for multi-modal sarcasm detection, which combines Interactive CLIP (InterCLIP) for enhanced text-image interaction encoding and Memory-Enhanced Predictor (MEP) for more robust and reliable sarcasm predictions.
-  • We propose an efficient training strategy that significantly reduces computational overhead compared to state-of-the-art methods.
-By introducing approximately 20.6x fewer trainable parameters, our approach reduces GPU memory usage by about 2.5x and accelerates computation by roughly 8.7x with a batch size of 128, all while maintaining superior performance on a single NVIDIA RTX 4090 GPU.
-  • Through extensive experiments on the MMSD and MMSD2.0 benchmarks, we show that InterCLIP-MEP improves accuracy by 1.08% and F1 score by 1.51% over state-of-the-art methods, especially on MMSD2.0.
-2 Related Work
-Early research in sarcasm detection focused primarily on text data Bouazizi and Ohtsuki (2015); Amir et al.
-(2016); Baziotis et al.
-(2018).
-With the rise of social media, detecting sarcasm in text-image pairs has become more challenging, driving the development of multi-modal approaches.
-Schifanella et al.
-(2016) were among the first to explore multi-modal social media posts for identifying sarcasm cues.
-Building on this, Cai et al.
-(2019) introduced the MMSD benchmark, demonstrating the effectiveness of a hierarchical fusion model that integrates image features.
-This benchmark has since become a foundation for multi-modal sarcasm detection, inspiring subsequent studies Xu et al.
-(2020); Pan et al.
-(2020); Liang et al.
-(2021, 2022); Liu et al.
-(2022); Qin et al.
-(2023); Wen et al.
-(2023); Tian et al.
-(2023); Wei et al.
-(2024).
-However, the MMSD benchmark was later found to contain spurious cues that could lead to model bias Qin et al.
-(2023).
-To mitigate this, Qin et al.
-(2023) introduced the MMSD2.0 benchmark, which removes these cues and corrects mislabeled samples.
-Re-evaluations on MMSD2.0 revealed significant performance drops in existing models, emphasizing the need for more robust approaches.
-In parallel, Tang et al.
-(2024) explored the use of large language models (LLMs) in multi-modal sarcasm detection, incorporating instruction templates and retrieval modules.
-While promising, the performance improvements were modest compared to the substantial computational cost.
-In this work, we present InterCLIP-MEP, a lightweight and efficient framework that achieves competitive performance without the high resource demands of LLM-based approaches.
-By overcoming the limitations of current methods, our approach offers a practical and scalable solution for multi-modal sarcasm detection.
-Figure 2: Overview of our framework.
-(I) Training Interactive CLIP (InterCLIP): Vision and text representations are extracted using separate encoders and embedded into the top-$n$ layers of the opposite modality’s encoder for interaction.
-The top-$n$ layers are fine-tuned with LoRA, while the rest of the encoder remains frozen.
-Final vision and text representations are concatenated and used to train a classification module for identifying multi-modal sarcasm.
-A projection module is also trained to project representations into a latent space.
-(II) Memory-Enhanced Predictor (MEP): During inference, InterCLIP generates interactive representations.
-The classification module assigns pseudo-labels, and the projection module provides projection features.
-MEP updates dynamic memory with these features and pseudo-labels.
-The final prediction of the current sample is made by comparing its projected feature with those in memory.
-3 Methodology
-An overview of InterCLIP-MEP is illustrated in Figure 2.
-Initially, we elaborate on the Interactive CLIP (InterCLIP) and its training strategy, followed by an in-depth explanation of the Memory-Enhanced Predictor (MEP).
-3.1 Interactive CLIP
-The input to Interactive CLIP (InterCLIP) is a text-image pair $\mathcal{P}=(T,I)$, where $T$ represents a piece of text and $I$ represents an image.
-Here, for simplicity, we do not consider the case of batch inputs.
-The text encoder $\mathcal{T}$ extracts the vanilla text representations $\mathbf{F}_{t}$:
-$\mathbf{F}_{t}$ $=\mathcal{T}(T)$
-=$=\{h^{t}_{\textrm{bos}}(t_{\textrm{bos}}),h^{t}_{1}(t_{1}),\dots,h^{t}_{n}(t_{n}),h^{t}_{\textrm{eos}}(t_{\textrm{eos}})\},$ (1)
-where $t_{i}$ denotes a text token, $n$ is the length of $T$ after tokenization, $t_{\textrm{bos}}$ and $t_{\textrm{eos}}$ are special tokens required by the text encoder.
-Here, $h^{t}_{i}(\cdot)\in\mathbb{R}^{d_{t}}$ represents the $d_{t}$-dimensional encoded representation of the corresponding token $t_{i}$, with $i$ ranging from 1 to $n$, including the beginning-of-sequence (bos) and end-of-sequence (eos) tokens.
-The vision encoder $\mathcal{V}$ extracts the vanilla image representations $\mathbf{F}_{v}$:
-$\mathbf{F}_{v}=\mathcal{V}(I)=\{h^{v}_{\textrm{cls}}(p_{\textrm{cls}}),h^{v}_{1}(p_{1}),\dots,h^{v}_{m}(p_{m})\},$ (2)
-where $I$ is processed into multiple patches $p_{i}$, $m$ is the number of patches, and $p_{\textrm{cls}}$ is a special token required by the visual encoder.
-Here, $h^{v}_{i}(\cdot)\in\mathbb{R}^{d_{v}}$ represents the $d_{v}$-dimensional encoded representation of the corresponding $p_{i}$, with $i$ ranging from 1 to $m$, including the classification (cls) token.
-Specifically, both $\mathbf{F}_{t}$ and $\mathbf{F}_{v}$ are representations from the final layer outputs of their respective encoders.
-Conditioning on $\mathbf{F}_{t}$ or $\mathbf{F}_{v}$, we can obtain the interactive text representations $\tilde{\mathbf{F}}_{t}$ or the interactive image representations $\tilde{\mathbf{F}}_{v}$:
-$\tilde{\mathbf{F}}_{t}=\mathcal{T}(T|\mathbf{F}_{v}),\tilde{\mathbf{F}}_{v}=\mathcal{V}(V|\mathbf{F}_{t}).$ (3)
-We use $\tilde{h}^{t}_{i}(\cdot)\in\mathbb{R}^{d_{t}}$ and $\tilde{h}^{v}_{i}(\cdot)\in\mathbb{R}^{d_{v}}$ to denote the re-encoded interactive representations of each text token and image patch, respectively.
-To be specific, we condition only the top-$n$ self-attention layers of the text or vision encoder, where $n$ is a hyperparameter that will be analyzed in the experiment section.
-Figure 3 illustrates the structure of the conditioned self-attention layers.
-Given that the text and vision encoder in CLIP share a similar architecture, for brevity, we denote the input representations to the self-attention layers of the text or vision encoder as $\mathbf{H}_{t/v}$, which are derived from the outputs of the previous layer.
-The previous layer can either be conditioned or non-conditioned.
-Due to the dimensional mismatch between the embedded representations $\mathbf{F}_{v/t}$ and the corresponding encoder representation space, we introduce an adapting projection layer $\mathcal{F}_{t/v}$ to project $\mathbf{F}_{v/t}$ into the appropriate representation space.
-To fuse the input representations $\mathbf{H}_{t/v}$ with the projected embedded representations $\mathbf{F}_{v/t}^{{}^{\prime}}=\mathcal{F}_{t/v}(\mathbf{F}_{v/t})$, we concatenate them and feed them into the attention layer to obtain the transformed representations.
-We then extract the transformed input representations $\mathbf{H}^{{}^{\prime}}_{t/v}$ from the output.
-Following Ganz et al.
-(2024), we apply a gated projection layer $\mathcal{G}_{t/v}$ along with the self-attention’s projection head $\mathcal{H}_{t/v}$ using a learnable gating mechanism to compute the self-attention output representations $\mathbf{H}^{{}^{\prime\prime}}_{t/v}$.
-Given the similarity between the self-attention layers of the vision encoder and the text encoder, we use the text encoder $\mathcal{T}$ to illustrate the process as follows:
-$\begin{split}\mathbf{F}_{v}^{{}^{\prime}}&=\mathcal{F}_{t}(\mathbf{F}_{v}),\quad\mathbf{F}_{v}\in\mathbb{R}^{m\times d_{v}},\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{m\times d_{t}},\\ \mathbf{H}^{{}^{\prime}}_{t}&=\textrm{Attn}_{t}(\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}})_{[:n]},\\ &\phantom{=}\mathbf{H}_{t},\mathbf{H}^{{}^{\prime}}_{t}\in\mathbb{R}^{n\times d_{t}},\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{(n+m)\times d_{t}},\\ \mathbf{H}^{{}^{\prime\prime}}_{t}&=\mathcal{H}_{t}(\mathbf{H}^{{}^{\prime}}_{t})+\mathcal{G}_{t}(\mathbf{H}^{{}^{\prime}}_{t})\cdot\textrm{tanh}(\beta_{t}),\mathbf{H}^{{}^{\prime\prime}}_{t}\in\mathbb{R}^{n\times d_{t}}.\end{split}$ (4)
-Here, $\oplus$ denotes the concatenation operation, and $\beta_{t}$ is a learnable gating parameter initialized to 0 to ensure training stability.
-The subsequent computation follows the original CLIP Radford et al.
-(2021), ultimately yielding the interactive representations $\tilde{\mathbf{F}}_{t}$.
-Figure 3: Structure of the conditional self-attention.
-InterCLIP supports three interaction modes for fusing text and image features into the final representation $\tilde{h}^{f}\in\mathbb{R}^{d_{t}+d_{v}}$:
-  • T2V: Text representations $\mathbf{F}_{t}$ are embedded into the vision encoder to produce interactive image representations $\tilde{\mathbf{F}}_{v}$.
-$\tilde{h}^{f}$ is formed by concatenating $h^{t}_{\textrm{eos}}$ and $\tilde{h}^{v}_{\textrm{cls}}$.
-  • V2T: Image representations $\mathbf{F}_{v}$ are embedded into the text encoder to produce interactive text representations $\tilde{\mathbf{F}}_{t}$.
-$\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}$ and $h^{v}_{\textrm{cls}}$.
-  • Two-way (TW): Both text and image representations $\mathbf{F}_{t}$ and $\mathbf{F}_{v}$ are embedded into each other’s encoders, resulting in $\tilde{\mathbf{F}}_{t}$ and $\tilde{\mathbf{F}}_{v}$.
-$\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}$ and $\tilde{h}^{v}_{\textrm{cls}}$.
-We will analyze the effectiveness of these three interaction modes in the experimental analysis.
-Training Strategy.
-As shown in Figure 2 (left), to adapt InterCLIP for MEP, we introduce an efficient training strategy.
-Using InterCLIP as the backbone to obtain fused features of the samples, we introduce a classification module and a projection module.
-Given the fused features of a batch of samples $\tilde{H}^{f}\in\mathbb{R}^{{N}\times(d_{t}+d_{v})}$, the classification module $\mathcal{F}_{c}$ calculates the probabilities $\hat{y}$ of these samples being sarcastic or non-sarcastic:
-$\hat{y}=\textrm{softmax}(\mathcal{F}_{c}(\tilde{H}^{f})),\quad\hat{y}\in\mathbb{R}^{N\times 2},$ (5)
-where $N$ denotes the batch size. We optimize $\mathcal{F}_{c}$ using binary cross-entropy loss:
-$\mathcal{L}^{c}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i,1})+(1-y_{i})\log(1-\hat{y}_{i,0})\right],$
-where $y_{i}$ denotes the label of the $i$-th sample, with sarcastic labeled as 1 and non-sarcastic as 0, and $\hat{y}_{i}$ denotes the prediction for the $i$-th sample.
-The projection module $\mathcal{F}_{p}$ maps $\tilde{H}^{f}$ into a latent feature space:
-$\hat{H}^{f}=\textrm{norm}(\mathcal{F}_{p}(\tilde{H}^{f})),\quad\hat{H}^{f}\in\mathbb{R}^{{N}\times d_{f}},$ (7)
-where $\textrm{norm}(\cdot)$ denotes L2 normalization, and $d_{f}$ represents the dimension of the projected features.
-In this space, the cosine distance between features of the same class is minimized, while the distance between features of different classes is maximized.
-We use a label-aware cosine similarity loss to optimize $\mathcal{F}_{p}$:
-$\mathcal{L}^{p}=\textrm{mean}(\hat{H}^{f}_{P}\cdot\hat{H}^{f^{T}}_{N})+\textrm{mean}(1-\hat{H}^{f}_{P}\cdot\hat{H}^{f^{T}}_{P})\\ +\textrm{mean}(1-\hat{H}^{f}_{N}\cdot\hat{H}^{f^{T}}_{N}),$ (8)
-where $\hat{H}^{f}_{P}$ and $\hat{H}^{f}_{N}$ represent the projected features of positive and negative samples, respectively.
-We fully train the modules $\mathcal{F}_{c}$, $\mathcal{F}_{p}$, the adapting projection layers ($\mathcal{F}_{t}$ and $\mathcal{F}_{v}$), the gated projection layers ($\mathcal{G}_{t}$ and $\mathcal{G}_{v}$), and the learnable gating parameters ($\beta_{t}$ and $\beta_{v}$).
-We use LoRA Hu et al.
-(2022) to fine-tune parts of the weight matrices $\mathbf{W}$ in the self-attention modules of all encoders, specifically various combinations of $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$.
-We consider $\mathbf{W}$ and the rank $r$ of LoRA as hyperparameters for our study.
-All learnable parts are optimized by minimizing the joint loss:
-$\mathcal{L}=\mathcal{L}^{c}+\mathcal{L}^{p}.$ (9)
-Algorithm 1 Memory-Enhanced Predictor
-Input: Memory size $L$, Learned InterCLIP model, classification module $\mathcal{F}_{c}$ and projection module $\mathcal{F}_{p}$
-Output: Final prediction $\hat{y}^{p}$
-1:  Initialize memory $\mathcal{M}\in\mathbf{0}^{2\times L\times d_{f}}$
-2:  Initialize index $\mathcal{I}\in\mathbf{0}^{2}$
-3:  Initialize entropy records $\mathcal{C}\in\mathbf{0}^{2\times L}$
-4:  for $i\leftarrow 1$ to $N_{\textrm{test}}$ do
-5:     $\tilde{h}^{f}_{i}\leftarrow\textrm{InterCLIP}(\mathcal{P}_{i})$
-6:     $\hat{y}_{i}\leftarrow\textrm{softmax}(\mathcal{F}_{c}(\tilde{h}^{f}_{i}))$
-7:     $\ell_{\textrm{pse}_{i}}\leftarrow\arg\max_{j}(\hat{y}_{i,j}),\;j\in\{0,1\}$
-8:     $c_{i}\leftarrow-\hat{y}_{i,0}\log{\hat{y}_{i,0}}-\hat{y}_{i,1}\log{\hat{y}_{i,1}}$
-9:     $\hat{h}^{f}_{i}\leftarrow\textrm{norm}(\mathcal{F}_{p}(\tilde{h}^{f}_{i}))$
-10:     if $\mathcal{I}[\ell_{\textrm{pse}_{i}}]<L$ then
-11:        $\mathcal{M}[\ell_{\textrm{pse}_{i}}][\mathcal{I}[\ell_{\textrm{pse}_{i}}]]\leftarrow\hat{h}^{f}_{i}$
-12:        $\mathcal{C}[\ell_{\textrm{pse}_{i}}][\mathcal{I}[\ell_{\textrm{pse}_{i}}]]\leftarrow c_{i}$
-13:        $\mathcal{I}[\ell_{\textrm{pse}_{i}}]\leftarrow\mathcal{I}[\ell_{\textrm{pse}_{i}}]+1$
-14:     else
-15:        $j\leftarrow\textrm{GetMaxIdx}(\mathcal{C}[\ell_{\textrm{pse}_{i}}])$
-16:        if $c_{i}<\mathcal{C}[\ell_{\textrm{pse}_{i}}][j]$ then
-17:           $\mathcal{M}[\ell_{\textrm{pse}_{i}}][j]\leftarrow\hat{h}^{f}_{i}$
-18:           $\mathcal{C}[\ell_{\textrm{pse}_{i}}][j]\leftarrow c_{i}$
-19:        end if
-20:     end if
-21:     $\textrm{logits}\leftarrow\left[\sum_{k=0}^{\mathcal{I}[0]}(\hat{h}^{f}_{i}\mathcal{M}[0]^{T})_{k},\sum_{k=0}^{\mathcal{I}[1]}(\hat{h}^{f}_{i}\mathcal{M}[1]^{T})_{k}\right]$
-22:     $\hat{y}^{p}_{i}\leftarrow\textrm{softmax}(\textrm{logits})$
-23:     yield $\hat{y}^{p}_{i}$
-24:  end for
-3.2 Memory-Enhanced Predictor
-As depicted in Figure 2 (right), we present the Memory-Enhanced Predictor (MEP) that builds upon the learned InterCLIP, along with the classification module and the projection module, leveraging the valuable historical knowledge of test samples to enhance multi-modal sarcasm detection.
-The detailed computational process of MEP is provided in Algorithm 1, where $N_{\textrm{test}}$ denotes the number of test samples.
-MEP uses the trained InterCLIP to extract fused features of the samples.
-It utilizes the classification module $\mathcal{F}_{c}$ to assign a pseudo-label $\ell_{\textrm{pse}_{i}}$ to each sample $\mathcal{P}_{i}$ and the projection module $\mathcal{F}_{p}$ to obtain the sample’s projected feature $\hat{h}^{f}_{i}$.
-To store valuable projected features of test samples as historical knowledge, MEP maintains a dynamic fixed-length dual-channel memory $\mathcal{M}\in\mathcal{R}^{2\times L\times d_{f}}$, where $L$ is the memory length per channel.
-The first channel stores projected features of non-sarcastic samples, while the second channel stores those of sarcastic samples.
-Based on the pseudo-label $\ell_{\textrm{pse}_{i}}$, the appropriate memory channel $\mathcal{M}[\ell_{\textrm{pse}_{i}}]$ is selected for updating.
-If the selected channel has available space, the sample’s projected features are added directly, and the prediction entropy is recorded.
-If the memory is full, the prediction entropy of all samples in the memory is compared with that of the current sample.
-Samples with the highest entropy are replaced, ensuring the retained samples have lower entropy.
-Finally, the current sample’s projected feature is combined with the historical features stored in both memory channels $\mathcal{M}$ using cosine similarity to yield the final prediction.
-4 Experiment
-MMSD/MMSD2.0 Sarcastic Non-sarcastic All
-Train 8,642/9,576 11,174/10,240 19,816/19,816
-Validation 959/1,042 1,451/1,368 2,410/2,410
-Test 959/1,037 1,450/1,372 2,409/2,409
-Table 1: Statistics of MMSD and MMSD2.0.
-Method MMSD2.0 MMSD
-Acc. (%) F1 (%) P (%) R (%) Acc. (%) F1 (%) P (%) R (%)
-Text
-TextCNN Kim (2014) 71.61* 69.52* 64.62* 75.22* 80.03* 75.32* 74.29* 76.39*
-Bi-LSTM Zhou et al. (2016) 72.48* 68.05* 68.02* 68.08* 81.90* 77.53* 76.66* 78.42*
-SMSD Xiong et al. (2019) 73.56* 69.97* 68.45* 71.55* 80.90* 75.82* 76.46* 75.18*
-RoBERTa Liu et al. (2019) 79.66* 76.21* 76.74* 75.70* 93.97* 92.45* 90.39* 94.59*
-Image
-ResNet He et al. (2016) 65.50* 57.58* 61.17* 54.39* 64.76* 61.53* 54.41* 70.80*
-ViT Dosovitskiy et al. (2021) 72.02* 69.72* 65.26* 74.83* 67.83* 63.40* 57.93* 70.07*
-Text-Image
-HFM Cai et al. (2019) 70.57* 66.88* 64.84* 69.05* 83.44* 80.18* 76.57* 84.15*
-Att-BERT Pan et al. (2020) 80.03* 77.04* 76.28* 77.82* 86.05* 82.92* 80.87* 85.08*
-CMGCN Liang et al. (2022) 79.83* 76.90* 75.82* 78.01* 86.54* 84.09* -* -*
-HKE Liu et al. (2022) 76.50* 72.25* 73.48* 71.07* 87.36* 72.25* 81.84* 86.48*
-DIP Wen et al. (2023) 80.59 78.23 75.52 81.14 89.59 87.17 87.76 86.58
-DynRT Tian et al. (2023) 70.37 68.55 63.02 75.15 93.59 91.93 90.30 93.62
-Multi-view CLIP Qin et al. (2023) 85.64* 84.10* 80.33* 88.24* 88.33* 85.55* 82.66* 88.65*
-G2SAM Wei et al. (2024) 79.43 78.07 72.04 85.20 90.48 88.48 87.95 89.02
-InterCLIP-MEP (Ours)
-w/o Inter ($L_{2}=1024,L_{1}=1280$) 86.05 84.81 79.83 90.45 88.75 86.31 83.73 89.05
-w/ TW ($L_{2}=128,L_{1}=1152$) 85.51 84.26 79.15 90.07 88.54 86.32 82.25 90.82
-w/ V2T ($L_{2}=640,L_{1}=1024$) 86.26 85.00 80.17 90.45 88.92 86.66 83.21 90.41
-w/ T2V ($L_{2}=1024,L_{1}=1152$) 86.72 85.61 80.20 91.80 88.83 86.37 84.02 88.84
-InterCLIP-MEP w/ RoBERTa (Ours)
-w/o Inter ($L_{2}=640,L_{1}=128$) 77.21 75.55 70.20 81.77 93.94 92.54 90.77 94.37
-w/ TW ($L_{2}=896,L_{1}=256$) 81.98 80.78 74.69 87.95 93.73 92.28 90.48 94.16
-w/ V2T ($L_{2}=640,L_{1}=512$) 76.96 75.26 69.98 81.39 93.94 92.54 90.69 94.47
-w/ T2V ($L_{2}=1024,L_{1}=384$) 82.81 81.55 75.81 88.24 93.73 92.28 90.56 94.06
-Table 2: Main results.
-We use * to indicate that the results are taken from Qin et al.
-(2023).
-- indicates that results are not reported.
-$\dagger$ indicates our reproduced results.
-Underlined values represent the best multi-modal baseline for comparison.
-Bold values indicate those that surpass the underlined baseline.
-$L_{2}$ for MMSD2.0 and $L_{1}$ for MMSD denote the optimal MEP memory sizes.
-Method Accuracy (%) Trainable Parameters (M) Fitting Time / Epoch (s) Inference Time (s) GPU Memory Peak (GB)
-Multi-view CLIP  85.64  165  488  51  15.59
-DIP  80.59  196  OOM  OOM  OOM
-G2SAM  79.43  116  90  13  18.32
-DynRT  70.37  25  370  26  8.03
-InterCLIP-MEP  86.72  8  55  6  6.14
-Table 3: Efficiency comparison of different methods.
-To demonstrate the efficiency of InterCLIP-MEP, we selected several recent baselines for comparison.
-The analysis was conducted using the MMSD2.0 dataset on a single NVIDIA RTX 4090 GPU with a batch size of 128.
-In the table, Fitting Time / Epoch indicates the time required for each epoch during training and validation and OOM indicates Out of Memory, referring to GPU memory overflow.
-w/o Inter w/ TW w/ V2T w/ T2V
-Variant Acc. (%) F1 (%) Acc. (%) F1 (%) Acc. (%) F1 (%) Acc. (%) F1 (%)
-Baseline 86.05 84.81 85.51 84.26 86.26 85.00 86.72 85.61
-w/o Proj 85.76 84.43 85.43 84.05 85.68 84.22 86.22 84.51
-w/o MEP 85.39 83.99 85.22 83.79 86.26 84.78 86.26 84.82
-w/o LoRA 82.44 77.73 76.42 74.37 73.31 72.22 75.13 71.79
-Table 4: Ablation study of InterCLIP-MEP, with Baseline denoting results without ablation.
-4.1 Experimental Settings
-Datasets and metrics.
-Following Qin et al.
-(2023), we evaluate performance on MMSD Cai et al.
-(2019) and MMSD2.0 Qin et al.
-(2023) using accuracy (Acc.), precision (P), recall (R), and F1-score (F1) as metrics.
-We present the statistics of the two datasets in Table 1.
-Baselines.
-We compare the effectiveness of the InterCLIP-MEP framework against several unimodal and multi-modal methods.
-For text modality methods, we compare with TextCNN Kim (2014), Bi-LSTM Zhou et al.
-(2016), SMSD Xiong et al.
-(2019), and RoBERTa Liu et al.
-(2019).
-For image modality methods, we compare with ResNet He et al.
-(2016) and ViT Dosovitskiy et al.
-(2021).
-We compare with state-of-the-art multi-modal methods, including HFM Cai et al.
-(2019), Att-BERT Pan et al.
-(2020), CMGCN Liang et al.
-(2022), HKE Liu et al.
-(2022), DIP Wen et al.
-(2023), DynRT Tian et al.
-(2023), Multi-view CLIP Qin et al.
-(2023), and G2SAM Wei et al.
-(2024), which employ various techniques such as hierarchical fusion, graph neural networks, and dynamic routing for multi-modal sarcasm detection.
-4.2 Main Results
-To validate the effectiveness of our InterCLIP-MEP framework, we conduct experiments using the original CLIP as the backbone instead of InterCLIP, referred to as w/o Inter.
-We compare this configuration with three interaction modes of InterCLIP: w/ V2T, w/ T2V, and w/ TW.
-For each experiment, we condition only the top four layers of the self-attention modules, with the projection dimension $d_{f}$ set to 1024.
-We set the LoRA rank $r$ to 8, fine-tuning the self-attention weight matrices $\mathbf{W}$, specifically $W_{k}$, $W_{v}$, and $W_{o}$.
-For the memory size $L$, we select the optimal size from a fixed set of candidate values $\mathbf{L}$  [^2] The fixed candidate values are {128, 256, 384, 512, 640, 768, 896, 1024, 1152, 1280}..
-The main results are shown in Table 2.
-Performance on MMSD2.0.
-For MMSD2.0, our framework consistently outperforms or matches the performance of state-of-the-art methods, whether using InterCLIP or the original CLIP as the backbone, as shown in Table 2 (InterCLIP-MEP).
-This demonstrates the effectiveness of our training strategy and MEP.
-Our results show that w/ V2T and w/ T2V outperform w/o Inter, demonstrating that InterCLIP captures text-image interactions more effectively.
-Furthermore, w/ T2V achieves superior performance compared to w/ V2T, likely due to the inherent complexity of the visual space, which presents challenges for the projection layer when mapping vision representations into the text encoder space.
-In contrast, w/ TW performs worse than other configurations, possibly because embedding representations within both encoders increases the learning difficulty.
-In summary, InterCLIP with T2V interaction, combined with our training strategy and MEP, delivers the most promising results.
-These findings underscore the robustness and adaptability of our framework, establishing it as a highly effective solution for capturing nuanced text-image interactions and addressing the complexities of multi-modal sarcasm detection.
-Performance on MMSD.
-For MMSD, the RoBERTa-based text modality baseline significantly outperforms other methods due to spurious cues in the text, enabling accurate predictions solely based on textual features Qin et al.
-(2023).
-Consequently, models like DynRT, G2SAM, and DIP, which utilize RoBERTa or BERT for text feature extraction, achieve high performance on MMSD but experience a significant drop on MMSD2.0.
-To further investigate, we conduct an additional experiment, InterCLIP-MEP w/ RoBERTa, replacing the original text encoder with RoBERTa.
-While this change led to state-of-the-art performance on MMSD, it disrupted InterCLIP’s modality alignment, causing a reasonable performance drop on MMSD2.0.
-This suggests that MMSD’s text data contains spurious cues that allow models to rely heavily on the text encoder, while MMSD2.0, having been cleaned, requires more robust multi-modal capabilities.
-We further find that the w/ V2T variant consistently outperforms others in both InterCLIP-MEP and InterCLIP-MEP w/ RoBERTa experiments, underscoring the model’s tendency to overly depend on text modality.
-Efficiency Comparison.
-Our training strategy demonstrates both remarkable effectiveness and outstanding efficiency.
-To validate this, we conducted a comprehensive comparative analysis against leading state-of-the-art methods, as detailed in Table 3.
-For instance, the Multi-view CLIP method Qin et al.
-(2023) employs a multi-layer Transformer encoder for feature fusion, which, while effective, introduces a significant number of trainable parameters.
-This results in slower training and inference speeds and greater memory consumption.
-Similarly, the DIP method Wen et al.
-(2023) caches historical samples during training, which hinders its ability to support large-batch training under limited resource conditions.
-In contrast, our method operates with a batch size of 128 while utilizing a significantly smaller number of trainable parameters, which translates to notably faster training and validation cycles.
-Furthermore, by incorporating minimal parameter modifications to adapt CLIP and utilizing simple yet effective linear layers for representation fusion, our approach achieves superior inference speeds and drastically reduced memory consumption.
-These results highlight the practicality of our framework, establishing it as a benchmark for both computational efficiency and performance in multi-modal sarcasm detection.
-4.3 Analysis of InterCLIP-MEP
-To robustly validate the effectiveness of InterCLIP-MEP, we conduct comprehensive ablation studies and case studies on the more reliable MMSD2.0 benchmark, offering deeper insights into its design and performance.
-In addition, we include visualization analyses to provide an intuitive understanding of how the framework processes multi-modal sarcasm cues.
-Ablation study.
-We remove the projection module $\mathcal{F}_{p}$ and train only the classification module $\mathcal{F}_{c}$ for prediction, denoted as w/o Proj.
-To test the necessity of using LoRA Hu et al.
-(2022) for fine-tuning, we keep the rest of InterCLIP-MEP unchanged and freeze all self-attention weights of InterCLIP, denoted as w/o LoRA, always selecting the optimal memory size $L$ for the MEP during inference.
-To evaluate the effectiveness of the MEP, we train both $\mathcal{F}_{p}$ and $\mathcal{F}_{c}$ but use only $\mathcal{F}_{c}$ during inference, denoted as w/o MEP.
-Table 4 reports all results.
-All variants show performance declines compared to the baseline, demonstrating the importance of each module in InterCLIP-MEP.
-For w/ TW and w/o Inter, the w/o MEP variant performed worse than the w/o Proj variant.
-However, for w/ T2V and w/ V2T, the w/o MEP variant performs better than the w/o Proj variant.
-This suggests that backbones with strong image-text interaction capabilities benefit from training the classification module $\mathcal{F}_{c}$ along with the projection module $\mathcal{F}_{p}$, even without using MEP during inference.
-We also find that not using LoRA to fine-tune the self-attention modules results in significant performance loss, indicating that the original CLIP’s vanilla space is not directly suitable for the sarcasm detection task.
-Case study.
-As shown in Figure 4, we select three examples to further demonstrate the robustness of InterCLIP-MEP.
-We observe that direct predictions through the classification module $\mathcal{F}_{c}$ result in high prediction entropy and incorrect outcomes.
-However, MEP effectively mitigates the issue for cases where $\mathcal{F}_{c}$ fails to correctly identify the results by using the historical knowledge of the test samples.
-This integration ensures that even in complex situations, the model maintains a high level of accuracy.
-Figure 4: Case study of InterCLIP-MEP.
-In the figure, GT represents the labels annotated by human experts, MEP represents the labels predicted by the memory-enhanced predictor, and $\mathcal{F}_{c}$ represents the labels provided by the classification module.
-Figure 5: Visual comparison between InterCLIP and original CLIP.
-Both InterCLIP and the original CLIP were fine-tuned using the same training set and identical parameters.
-The key distinction is that the original CLIP does not incorporate interaction.
-Visulization.
-To further validate that InterCLIP is capable of more effectively capturing the interactive information between text and images compared to original CLIP, thereby aiding in the detection of sarcasm cues, we use GradCAM Selvaraju et al.
-(2017) to visualize the areas of focus during the inference process of the visual model in Figure 5.
-We observe that in the first example, which complains about the unpleasant odor caused by the sea lions, InterCLIP focuses more accurately on the location of the sea lions compared to the original CLIP.
-In the second example, which complains about traffic congestion, InterCLIP correctly focuses on the distribution of cars on the road, whereas the original CLIP’s focus is scattered.
-5 Conclusion
-In this paper, we propose InterCLIP-MEP, a novel framework for multi-modal sarcasm detection that directly addresses the challenges of modeling nuanced text-image interactions and managing prediction uncertainty.
-We design Interactive CLIP (InterCLIP) to embed cross-modal information within text and image encoders, enabling a deeper understanding of sarcasm cues.
-Additionally, we develop a Memory-Enhanced Predictor (MEP) to dynamically leverage historical sample knowledge, making our inference process more robust and adaptive.
-Through extensive experiments on MMSD and MMSD2.0 benchmarks, we demonstrate that InterCLIP-MEP achieves state-of-the-art performance while significantly reducing computational costs.
-By requiring fewer trainable parameters and less GPU memory, our method offers a lightweight, efficient, and scalable solution, setting a new benchmark for multi-modal sarcasm detection.
-Limitations
-While InterCLIP-MEP delivers strong performance, there remain areas for further refinement.
-For instance, the framework could benefit from additional techniques to better capture sarcasm that heavily relies on subtle cultural or highly specific contextual cues.
-Moreover, extending its application to more diverse and less structured datasets could be an interesting direction for future work, further broadening its practical applicability.
-Ethical Considerations
-This work focuses on advancing multi-modal sarcasm detection to improve understanding of complex communication in online content.
-While the proposed framework enhances detection accuracy and efficiency, potential misuse must be considered.
-Automated sarcasm detection could inadvertently amplify biases present in training data or be deployed for unethical purposes, such as targeted content moderation or surveillance.
-To mitigate these risks, we encourage the responsible use of this technology and emphasize the importance of using diverse and unbiased datasets during training to minimize unintended consequences.
-Furthermore, this research strictly adheres to ethical guidelines for data collection and usage.
-Appendix A Implementation Details
-The model training and testing were conducted using PyTorch Lightning [^3] https://lightning.ai/.
-InterCLIP was constructed by leveraging the Transformers library Wolf et al.
-(2020).
-For the MMSD2.0 experiments, the initial weights for InterCLIP are based on clip-vit-base-patch32444https://huggingface.co/openai/clip-vit-base-patch32.
-For the MMSD experiments, we utilized the roberta-ViT-B-32 model architecture provided by OpenCLIP [^5] https://github.com/mlfoundations/open_clip, with the pre-trained checkpoint laion2b_s12b_b32k [^6] https://huggingface.co/laion/CLIP-ViT-B-32-roberta-base-laion2B-s12B-b32k.
-Custom scripts were developed to adapt its format to the Transformers library, ensuring compatibility with our framework.
-The model parameters were optimized using AdamW Loshchilov and Hutter (2019), with a learning rate set to 1e-4 for the LoRA fine-tuning modules and 5e-4 for other trainable modules.
-A cosine annealing scheduler with warmup was employed to dynamically adjust the learning rate, where the warmup steps constituted the first 20% of the total optimization steps, and the minimum learning rate was set to 1% of the initial rate.
-For the modules $\mathcal{G}_{t/v}$, $\mathcal{F}_{t/v}$, $\mathcal{F}_{c}$, and $\mathcal{F}_{p}$, simple multi-layer perceptrons (MLPs) were utilized.
-The training processing was performed with a batch size of 64 for 3 epochs.
-All experiments were run on a machine equipped with an NVIDIA RTX 4090 GPU.
-Appendix B Experimental Details
-B.1 Hyperparameter details
-We summarize the hyperparameters involved in InterCLIP-MEP and their descriptions in Table 5.
-The hyperparameter settings for obtaining the main results in the paper are summarized in Table 6.
-For other baseline methods, we follow the optimal hyperparameter settings they reported.
-Parameter Description
-$r$ The rank of LoRA, determining the dimension of the low-rank update matrices.
-$\mathbf{W}$ The weight matrices in the self-attention module fine-tuned using LoRA, specifically targeting combinations of $W_{\{q,k,v,o\}}$.
-top-$n$ The number of top self-attention layers conditioned during fine-tuning.
-$d_{f}$ The dimensionality of the latent space for the projected features.
-$\mathbf{L}$ The configurable range of memory sizes maintained by the Memory-Enhanced Predictor (MEP).
-Table 5: Summary of hyperparameters.
-Param. Value
-For trainer
-epoch 3
-batch_size 64
-lr 5e-4
-lora_lr 1e-4
-warmup_ratio 0.2
-min_lr_rate 0.01
-For our model
-$r$ 8
-top-$n$ 4
-$d_{f}$ 1024
-$\mathbf{W}$ $W_{k},W_{v},W_{o}$
-$\mathbf{L}$128, 256, 384, 512, 640, 768, 896, 1024, 1152, 1280}
-Table 6: Hyperparameter settings.
-Figure 6: Hyperparameter study curves for w/ T2V. Panel (d) compares results with those from using only the classification module $\mathcal{F}_{c}$ for prediction.
-B.2 Hyperparameter study
-We further investigate the method using Interactive-CLIP with T2V interaction as the backbone.
-Keeping the other hyperparameters constant, we condition different top-$n$ layers of the self-attention modules.
-We also study the impact of different projection dimensions $d_{f}$, different LoRA ranks $r$, and different memory sizes $L$ on the w/ T2V method.
-We present all results in Figure 6.
-Figure 6(a) shows that conditioning the top four self-attention layers yields the best results.
-From Figure 6(b), a rank of 8 is optimal.
-Figure 6(c) indicates the projection dimension is best at 256 or 1024.
-Figure 6(d) reveals that a memory size of 640 in MEP outperforms the others, confirming the value of historical sample knowledge.
-B.3 Empirical study of self-attention fine-tuning and interaction modes
-Keeping the other hyperparameters constant as shown in Table 6, we fine-tune all possible weight matrices $\mathbf{W}$ and employ different interaction modes of InterCLIP as the backbone.
-We consistently select the optimal memory size from $\mathbf{L}$ for MEP.
-We calculate the average metrics for different methods and different weight matrices.
-The results are presented in Table 7 and Table 8.
-We observe that fine-tuning the weight matrices $W_{k},W_{v},W_{o}$ and using the T2V interaction mode of InterCLIP are the best choices for InterCLIP-MEP.
-$\mathbf{W}$ Mean Acc. (%) Mean F1 (%)
-$W_{q}$ 85.14 83.99
-$W_{k}$ 85.09 84.00
-$W_{v}$ 85.34 84.24
-$W_{o}$ 85.39 84.23
-$W_{q},W_{k}$ 85.36 84.16
-$W_{q},W_{v}$ 85.63 84.40
-$W_{q},W_{o}$ 85.73 84.49
-$W_{k},W_{o}$ 85.67 84.43
-$W_{v},W_{o}$ 85.73 84.54
-$W_{k},W_{v}$ 85.70 84.51
-$W_{q},W_{k},W_{o}$ 85.92 84.63
-$W_{q},W_{v},W_{o}$ 85.87 84.60
-$W_{q},W_{k},W_{v}$ 86.05 84.75
-$W_{k},W_{v},W_{o}$ 86.14 84.92
-$W_{q},W_{k},W_{v},W_{o}$ 85.82 84.55
-Table 7: Average results of fine-tuning different weight matrices $\mathbf{W}$ across four baseline methods.
-Method Mean Acc. (%) Mean F1 (%)
-w/o Inter 85.49 84.32
-w/ TW 85.66 84.44
-w/ V2T 85.62 84.41
-w/ T2V 85.78 84.55
-Table 8: Average results of four baseline methods for fine-tuning different weight matrices $\mathbf{W}$.
-Appendix C More Visual Examples
-Figure 7 presents additional examples illustrating the focus differences between InterCLIP and CLIP.
-These visualizations highlight InterCLIP’s improved ability to capture sarcasm-related cues by focusing on relevant areas in the images.
-Figure 7: Additional visual examples showcasing InterCLIP’s improved focus on sarcasm-related visual cues compared to CLIP.
-Appendix D Extended Experiments
-To further verify the performance of our framework, we conducted an extended set of experiments.
-Sarcastic Non-sarcastic All
-Train 4,014 46,265 50,279
-Validation 1,125 13,097 14,222
-Test 555 6,772 7,327
-Table 9: Statistics of DocMSU.
-Method Acc. F1 P R
-BERT-base* 87.12 86.51 77.61 70.37
-Swin-Transformer* 74.83 61.51 67.57 56.45
-CMGCN* 88.12 75.23 78.11 72.55
-CLIP* 96.19 77.62 78.99 76.30
-ViLT* 93.15 41.44 69.03 29.61
-Du et al. (2024)* 97.83 87.25 81.20 94.27
-InterCLIP-MEP (Ours)
-w/o Inter 97.84 87.48 78.08 99.45
-w/ TW 97.79 87.24 77.48 99.81
-w/ V2T 97.83 87.45 77.81 99.82
-w/ T2V 97.67 86.65 76.45 99.99
-Table 10: Results of the extended experiments.
-Underline results denote the compared SOTA baseline, boldface highlights results that surpass the baseline, and * indicates results sourced from Du et al.
-(2024).
-D.1 Benchmark
-DocMSU Du et al.
-(2024) is a recently introduced multi-modal sarcasm benchmark designed specifically for long-text analysis.
-It facilitates the evaluation of multi-modal sarcasm comprehension as well as detection tasks.
-In this study, we concentrate on the multi-modal sarcasm detection task.
-The benchmark statistics can be found in Table 9.
-D.2 Baselines
-We follow the evaluation protocol of Du et al.
-(2024).
-We compare against unimodal baselines: BERT-base (text-only) Devlin (2018) and Swin Transformer (image-only) Liu et al.
-(2021).
-For multi-modal approaches, we include CLIP Radford et al.
-(2021), Vision-and-Language Transformer (ViLT) Kim et al.
-(2021), and CMGCN Liang et al.
-(2022).
-The method proposed by Du et al.
-(2024) is taken as the state-of-the-art baseline.
-D.3 Results
-Our InterCLIP-MEP framework demonstrates strong performance across various configurations, as shown in Table 10.
-In particular, the w/o Inter and w/ V2T variants consistently achieve higher F1 scores compared to the baselines, thereby showcasing their robustness in handling multi-modal sarcasm detection tasks.
-Notably, all variants achieve nearly perfect recall, highlighting their outstanding capability in accurately identifying sarcasm across diverse datasets.
-The w/o Inter variant also achieves the highest accuracy, further demonstrating its effectiveness and precision.
-Overall, the comprehensive results in Table 10 affirm the unparalleled effectiveness of InterCLIP’s modality interaction mechanism and our proposed memory-enhanced predictor (MEP), particularly in surpassing existing state-of-the-art methods in both accuracy and F1 score, thus setting a new benchmark in the field.
-Appendix E Other Related Works
-E.1 CLIP adaptation
-The Contrastive Language-Image Pretraining (CLIP) model Radford et al.
-(2021) excels in vision-language tasks.
-Adapting CLIP for specific domains has shown substantial improvements, as demonstrated by Li et al.
-(2022) for phrase localization, Liang et al.
-(2023) for open-vocabulary semantic segmentation, and Wang et al.
-(2023) for action recognition.
-In this work, inspired by Ganz et al.
-(2024), we conditionally enhance both the text and vision encoders of CLIP, making it more effective in capturing the interplay between text and images to identify multi-modal sarcasm cues.
-Unlike Ganz et al.
-(2024), who focused solely on embedding text representations into the vision encoder, we also explore embedding image representations into the text encoder.
-Furthermore, their approach is limited to general classification tasks and does not address the complexities of multi-modal sarcasm detection.
-E.2 Memory-enhanced prediction
-Inspired by cognitive science Stokes (2015); Baddeley (2000), memory has been introduced to enhance neural networks Weston et al.
-(2014); Sukhbaatar et al.
-(2015).
-Several studies Wu et al.
-(2018); Wen et al.
-(2023) have used memory mechanisms to improve model training, and some Zhang et al.
-(2024); Wei et al.
-(2024) leverage memory to store historical knowledge, enhancing prediction accuracy.
-In this work, we introduce a memory-enhanced predictor for multi-modal sarcasm detection.
-In contrast to other methods, our memory dynamically updates during testing, utilizing relevant historical information for improved accuracy and robustness.
-Appendix F List of Symbols
-In Table 11, we have listed the main symbols used in the paper and their descriptions.
+(wei2024g2sam) introduced G2SAM for multi-modal sarcasm detection, which leverages a graph-based memory to capture global semantic relations.
+However, both approaches rely on static or pre-constructed memory structures that do not evolve with incoming test samples.
+By contrast, our MEP employs a dynamic, entropy-aware buffer that continuously updates with reliable test exemplars, making memory an adaptive and lightweight inference component tailored specifically for the challenges of sarcasm detection (gibbs2007irony; bamman2015contextualized).
+Table 1. List of symbols.
 Symbol Description
 $T$ $T$ denotes a short text.
-$I$ $I$ represents an image.
+II II represents an image.
 $\mathcal{P}$ $\mathcal{P}$ denotes a text-image pair $(T,I)$.
@@
 $\hat{h}^{f}$ $\hat{h}^{f}$ represents the feature of a sample’s fused feature after transformation by $\mathcal{F}_{p}$ and L2 normalization.
-Table 11: List of symbols
+3. Methodology
+An overview of InterCLIP-MEP is illustrated in Figure 2.
+The framework is first introduced through a detailed exposition of Interactive CLIP (InterCLIP) and its associated efficient training strategy, followed by a comprehensive analysis of the Memory-Enhanced Predictor (MEP).
+For clarity, Table 1 provides a summary of the primary mathematical symbols employed throughout this work.
+Figure 2.
+Overview of our framework.
+(I) Training Interactive CLIP (InterCLIP): Vision and text representations are extracted using separate encoders and embedded into the top-$n$ layers of the opposite modality’s encoder for interaction.
+The top-$n$ layers are fine-tuned with LoRA, while the rest of the encoder remains frozen.
+Final vision and text representations are concatenated and used to train a classification module for identifying multi-modal sarcasm.
+A projection module is also trained to project representations into a latent space.
+(II) Memory-Enhanced Predictor (MEP): During inference, InterCLIP generates interactive representations.
+The classification module assigns pseudo-labels, and the projection module provides projection features.
+MEP updates dynamic dual-channel memory with these features and pseudo-labels.
+The final prediction of the current sample is made by comparing its projected feature with those in memory.
+3.1. Interactive CLIP
+The input to Interactive CLIP (InterCLIP) is a text-image pair $\mathcal{P}=(T,I)$, where $T$ represents a piece of text and $I$ represents an image.
+Here, for simplicity, we do not consider the case of batch inputs.
+The text encoder $\mathcal{T}$ extracts the vanilla text representations $\mathbf{F}_{t}$:
+(1) $\mathbf{F}_{t}=\mathcal{T}(T)=\{h^{t}_{\textrm{bos}}(t_{\textrm{bos}}),h^{t}_{1}(t_{1}),\dots,h^{t}_{n}(t_{n}),h^{t}_{\textrm{eos}}(t_{\textrm{eos}})\},$
+where $t_{i}$ denotes a text token, $n$ is the length of $T$ after tokenization, $t_{\textrm{bos}}$ and $t_{\textrm{eos}}$ are special tokens required by the text encoder.
+Here, $h^{t}_{i}(\cdot)\in\mathbb{R}^{d_{t}}$ represents the $d_{t}$-dimensional encoded representation of the corresponding token $t_{i}$, with $i$ ranging from 1 to $n$, including the beginning-of-sequence (bos) and end-of-sequence (eos) tokens.
+The vision encoder $\mathcal{V}$ extracts the vanilla image representations $\mathbf{F}_{v}$:
+(2) $\mathbf{F}_{v}=\mathcal{V}(I)=\{h^{v}_{\textrm{cls}}(p_{\textrm{cls}}),h^{v}_{1}(p_{1}),\dots,h^{v}_{m}(p_{m})\},$
+where $I$ is processed into multiple patches $p_{i}$, $m$ is the number of patches, and $p_{\textrm{cls}}$ is a special token required by the visual encoder.
+Here, $h^{v}_{i}(\cdot)\in\mathbb{R}^{d_{v}}$ represents the $d_{v}$-dimensional encoded representation of the corresponding $p_{i}$, with $i$ ranging from 1 to $m$, including the classification (cls) token.
+Specifically, both $\mathbf{F}_{t}$ and $\mathbf{F}_{v}$ are representations from the final layer outputs of their respective encoders.
+Conditioning on $\mathbf{F}_{t}$ or $\mathbf{F}_{v}$, we can obtain the interactive text representations $\tilde{\mathbf{F}}_{t}$ or the interactive image representations $\tilde{\mathbf{F}}_{v}$:
+(3) $\tilde{\mathbf{F}}_{t}=\mathcal{T}(T|\mathbf{F}_{v}),\tilde{\mathbf{F}}_{v}=\mathcal{V}(V|\mathbf{F}_{t}).$
+Figure 3. Structure of the conditional self-attention.
+We use $\tilde{h}^{t}_{i}(\cdot)\in\mathbb{R}^{d_{t}}$ and $\tilde{h}^{v}_{i}(\cdot)\in\mathbb{R}^{d_{v}}$ to denote the re-encoded interactive representations of each text token and image patch, respectively.
+To be specific, we condition only the top-$n$ self-attention layers of the text or vision encoder, where $n$ is a hyperparameter that will be analyzed in the experiment section.
+Figure 3 illustrates the structure of the conditioned self-attention layers.
+Given that the text and vision encoder in CLIP share a similar architecture, for brevity, we denote the input representations to the self-attention layers of the text or vision encoder as $\mathbf{H}_{t/v}$, which are derived from the outputs of the previous layer.
+The previous layer can either be conditioned or non-conditioned.
+Due to the dimensional mismatch between the embedded representations $\mathbf{F}_{v/t}$ and the corresponding encoder representation space, we introduce an adapting projection layer $\mathcal{F}_{t/v}$ to project $\mathbf{F}_{v/t}$ into the appropriate representation space.
+To fuse the input representations $\mathbf{H}_{t/v}$ with the projected embedded representations $\mathbf{F}_{v/t}^{{}^{\prime}}=\mathcal{F}_{t/v}(\mathbf{F}_{v/t})$, we concatenate them and feed them into the attention layer to obtain the transformed representations.
+We then extract the transformed input representations $\mathbf{H}^{{}^{\prime}}_{t/v}$ from the output.
+Following (ganz2024question), we apply a gated projection layer $\mathcal{G}_{t/v}$ along with the self-attention’s projection head $\mathcal{H}_{t/v}$ using a learnable gating mechanism to compute the self-attention output representations $\mathbf{H}^{{}^{\prime\prime}}_{t/v}$.
+Given the similarity between the self-attention layers of the vision encoder and the text encoder, we use the text encoder $\mathcal{T}$ to illustrate the process as follows:
+(4) $\mathbf{F}_{v}^{{}^{\prime}}=\mathcal{F}_{t}(\mathbf{F}_{v}),\quad\mathbf{F}_{v}\in\mathbb{R}^{m\times d_{v}},\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{m\times d_{t}},$
+(5) $\mathbf{H}^{{}^{\prime}}_{t}=\textrm{Attn}_{t}(\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}})_{[:n]},\quad\mathbf{H}_{t},\mathbf{H}^{{}^{\prime}}_{t}\in\mathbb{R}^{n\times d_{t}},\mathbf{H}_{t}\oplus\mathbf{F}_{v}^{{}^{\prime}}\in\mathbb{R}^{(n+m)\times d_{t}},$
+(6) $\mathbf{H}^{{}^{\prime\prime}}_{t}=\mathcal{H}_{t}(\mathbf{H}^{{}^{\prime}}_{t})+\mathcal{G}_{t}(\mathbf{H}^{{}^{\prime}}_{t})\cdot\textrm{tanh}(\beta_{t}),\mathbf{H}^{{}^{\prime\prime}}_{t}\in\mathbb{R}^{n\times d_{t}}.$
+Here, $\oplus$ denotes the concatenation operation, and $\beta_{t}$ is a learnable gating parameter initialized to 0 to ensure training stability.
+The subsequent computation follows the original CLIP, ultimately yielding the interactive representations $\tilde{\mathbf{F}}_{t}$.
+InterCLIP supports three interaction modes for fusing text and image features into the final representation $\tilde{h}^{f}\in\mathbb{R}^{d_{t}+d_{v}}$:
+  • T2V: Text representations $\mathbf{F}_{t}$ are embedded into the vision encoder to produce interactive image representations $\tilde{\mathbf{F}}_{v}$.
+$\tilde{h}^{f}$ is formed by concatenating $h^{t}_{\textrm{eos}}$ and $\tilde{h}^{v}_{\textrm{cls}}$.
+  • V2T: Image representations $\mathbf{F}_{v}$ are embedded into the text encoder to produce interactive text representations $\tilde{\mathbf{F}}_{t}$.
+$\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}$ and $h^{v}_{\textrm{cls}}$.
+  • Two-way (TW): Both text and image representations $\mathbf{F}_{t}$ and $\mathbf{F}_{v}$ are embedded into each other’s encoders, resulting in $\tilde{\mathbf{F}}_{t}$ and $\tilde{\mathbf{F}}_{v}$.
+$\tilde{h}^{f}$ is formed by concatenating $\tilde{h}^{t}_{\textrm{eos}}$ and $\tilde{h}^{v}_{\textrm{cls}}$.
+We will analyze the effectiveness of these three interaction modes in the experimental analysis.
+3.2. Training Strategy
+As shown in Figure 2 (left), to adapt InterCLIP for MEP, we introduce an efficient training strategy.
+Using InterCLIP as the backbone to obtain fused features of the samples, we introduce a classification module and a projection module.
+Given the fused features of a batch of samples $\tilde{H}^{f}\in\mathbb{R}^{{N}\times(d_{t}+d_{v})}$, the classification module $\mathcal{F}_{c}$ calculates the probabilities $\hat{y}$ of these samples being sarcastic or non-sarcastic:
+(7) $\hat{y}=\textrm{softmax}(\mathcal{F}_{c}(\tilde{H}^{f})),\quad\hat{y}\in\mathbb{R}^{N\times 2},$
+where $N$ denotes the batch size. We optimize $\mathcal{F}_{c}$ using binary cross-entropy loss:
+(8) $\mathcal{L}^{c}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i,1})+(1-y_{i})\log(1-\hat{y}_{i,0})\right],$
+where $y_{i}$ denotes the label of the $i$-th sample, with sarcastic labeled as 1 and non-sarcastic as 0, and $\hat{y}_{i}$ denotes the prediction for the $i$-th sample.
+The projection module $\mathcal{F}_{p}$ maps $\tilde{H}^{f}$ into a latent feature space:
+(9) $\hat{H}^{f}=\textrm{norm}(\mathcal{F}_{p}(\tilde{H}^{f})),\quad\hat{H}^{f}\in\mathbb{R}^{{N}\times d_{f}},$
+where $\textrm{norm}(\cdot)$ denotes L2 normalization, and $d_{f}$ represents the dimension of the projected features.
+In this space, the cosine distance between features of the same class is minimized, while the distance between features of different classes is maximized.
+We use a label-aware cosine similarity loss to optimize $\mathcal{F}_{p}$:
+(10) $\mathcal{L}^{p}=\textrm{mean}(\hat{H}^{f}_{P}\cdot\hat{H}^{f^{T}}_{N})+\textrm{mean}(1-\hat{H}^{f}_{P}\cdot\hat{H}^{f^{T}}_{P})+\textrm{mean}(1-\hat{H}^{f}_{N}\cdot\hat{H}^{f^{T}}_{N}),$
+where $\hat{H}^{f}_{P}$ and $\hat{H}^{f}_{N}$ represent the projected features of positive and negative samples, respectively.
+We fully train the modules $\mathcal{F}_{c}$, $\mathcal{F}_{p}$, the adapting projection layers ($\mathcal{F}_{t}$ and $\mathcal{F}_{v}$), the gated projection layers ($\mathcal{G}_{t}$ and $\mathcal{G}_{v}$), and the learnable gating parameters ($\beta_{t}$ and $\beta_{v}$).
+We use LoRA (edward2022lora) to fine-tune parts of the weight matrices $\mathbf{W}$ in the self-attention modules of all encoders, specifically various combinations of $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$.
+We consider $\mathbf{W}$ and the rank $r$ of LoRA as hyperparameters for our study.
+All learnable parts are optimized by minimizing the joint loss:
+(11) $\mathcal{L}=\mathcal{L}^{c}+\mathcal{L}^{p}.$
+3.3. Memory-Enhanced Predictor
+As depicted in Figure 2 (right), we present the Memory-Enhanced Predictor (MEP) that builds upon the learned InterCLIP, along with the classification module and the projection module, leveraging the valuable historical knowledge of test samples to enhance multi-modal sarcasm detection.
+Algorithm 1 Memory-Enhanced Predictor
+1:Input: Memory size $L$, Learned InterCLIP model, classification module $\mathcal{F}_{c}$ and projection module $\mathcal{F}_{p}$
+2:Output: Final prediction $\hat{y}^{p}$
+3: Initialize memory $\mathcal{M}\in\mathbf{0}^{2\times L\times d_{f}}$
+4: Initialize index $\mathcal{I}\in\mathbf{0}^{2}$
+5: Initialize entropy records $\mathcal{C}\in\mathbf{0}^{2\times L}$
+6:for $i\leftarrow 1$ to $N_{\textrm{test}}$ do
+7:  $\tilde{h}^{f}_{i}\leftarrow\textrm{InterCLIP}(\mathcal{P}_{i})$
+8:  $\hat{y}_{i}\leftarrow\textrm{softmax}(\mathcal{F}_{c}(\tilde{h}^{f}_{i}))$
+9:  $\ell_{\textrm{pse}_{i}}\leftarrow\arg\max_{j}(\hat{y}_{i,j}),\;j\in\{0,1\}$
+10:  $c_{i}\leftarrow-\hat{y}_{i,0}\log{\hat{y}_{i,0}}-\hat{y}_{i,1}\log{\hat{y}_{i,1}}$
+11:  $\hat{h}^{f}_{i}\leftarrow\textrm{norm}(\mathcal{F}_{p}(\tilde{h}^{f}_{i}))$
+12:  if $\mathcal{I}[\ell_{\textrm{pse}_{i}}]<L$ then
+13:   $\mathcal{M}[\ell_{\textrm{pse}_{i}}][\mathcal{I}[\ell_{\textrm{pse}_{i}}]]\leftarrow\hat{h}^{f}_{i}$
+14:   $\mathcal{C}[\ell_{\textrm{pse}_{i}}][\mathcal{I}[\ell_{\textrm{pse}_{i}}]]\leftarrow c_{i}$
+15:   $\mathcal{I}[\ell_{\textrm{pse}_{i}}]\leftarrow\mathcal{I}[\ell_{\textrm{pse}_{i}}]+1$
+16:  else
+17:   $j\leftarrow\textrm{GetMaxIdx}(\mathcal{C}[\ell_{\textrm{pse}_{i}}])$
+18:   if $c_{i}<\mathcal{C}[\ell_{\textrm{pse}_{i}}][j]$ then
+19:    $\mathcal{M}[\ell_{\textrm{pse}_{i}}][j]\leftarrow\hat{h}^{f}_{i}$
+20:    $\mathcal{C}[\ell_{\textrm{pse}_{i}}][j]\leftarrow c_{i}$
+21:   end if
+22:  end if
+23:  $\textrm{logits}\leftarrow\left[\sum_{k=0}^{\mathcal{I}[0]}(\hat{h}^{f}_{i}\mathcal{M}[0]^{T})_{k},\sum_{k=0}^{\mathcal{I}[1]}(\hat{h}^{f}_{i}\mathcal{M}[1]^{T})_{k}\right]$
+24:  $\hat{y}^{p}_{i}\leftarrow\textrm{softmax}(\textrm{logits})$
+25:  yield $\hat{y}^{p}_{i}$
+26:end for
+The detailed computational process of MEP is provided in Algorithm 1, where $N_{\textrm{test}}$ denotes the number of test samples.
+MEP uses the trained InterCLIP to extract fused features of the samples.
+It utilizes the classification module $\mathcal{F}_{c}$ to assign a pseudo-label $\ell_{\textrm{pse}_{i}}$ to each sample $\mathcal{P}_{i}$ and the projection module $\mathcal{F}_{p}$ to obtain the sample’s projected feature $\hat{h}^{f}_{i}$.
+To store valuable projected features of test samples as historical knowledge, MEP maintains a dynamic fixed-length dual-channel memory $\mathcal{M}\in\mathcal{R}^{2\times L\times d_{f}}$, where $L$ is the memory length per channel.
+The first channel stores projected features of non-sarcastic samples, while the second channel stores those of sarcastic samples.
+Based on the pseudo-label $\ell_{\textrm{pse}_{i}}$, the appropriate memory channel $\mathcal{M}[\ell_{\textrm{pse}_{i}}]$ is selected for updating.
+If the selected channel has available space, the sample’s projected features are added directly, and the prediction entropy is recorded.
+If the memory is full, the prediction entropy of all samples in the memory is compared with that of the current sample.
+Samples with the highest entropy are replaced, ensuring the retained samples have lower entropy.
+Finally, the current sample’s projected feature is combined with the historical features stored in both memory channels $\mathcal{M}$ using cosine similarity to yield the final prediction.
+Table 2. Statistics of MMSD and MMSD2.0.
+MMSD/MMSD2.0 Sarcastic Non-sarcastic All
+Train 8,642/9,576 11,174/10,240 19,816/19,816
+Validation 959/1,042 1,451/1,368 2,410/2,410
+Test 959/1,037 1,450/1,372 2,409/2,409
+4. Experiments
+4.1. Experimental Settings
+Datasets and metrics.
+Following (qin2023mmsd2), we evaluate performance on MMSD (Cai2019MultiModalSD) and MMSD2.0 (qin2023mmsd2) using accuracy (Acc.), precision (P), recall (R), and F1-score (F1) as metrics.
+We present the statistics of the two datasets in Table 2.
+Baselines.
+We compare the effectiveness of the InterCLIP-MEP framework against several unimodal and multi-modal methods.
+For text modality methods, we compare with TextCNN (kim-2014-convolutional), Bi-LSTM (zhou-etal-2016-attention), SMSD (xiong2019sarcasm), and RoBERTa (liu2019roberta).
+For image modality methods, we compare with ResNet (he2015deep) and ViT (dosovitskiy2020image).
+We compare with state-of-the-art multi-modal methods, including HFM (Cai2019MultiModalSD), Att-BERT (pan-etal-2020-modeling), CMGCN (Liang2022MultiModalSD), HKE (Liu2022TowardsMS), DIP (wen2023dip), DynRT (Tian2023DynamicRT), Multi-view CLIP (qin2023mmsd2), and G2SAM (wei2024g2sam), which employ various techniques such as hierarchical fusion, graph neural networks, and dynamic routing for multi-modal sarcasm detection.
+4.2. Implementation Details
+The model training and testing were conducted using PyTorch Lightning [^1] https://lightning.ai/.
+InterCLIP was constructed by leveraging the Transformers library (wolf-etal-2020-transformers).
+For the MMSD2.0 experiments, the initial weights for InterCLIP are based on clip-vit-base-patch32222https://hf.co/openai/clip-vit-base-patch32.
+For the MMSD experiments, we utilized the roberta-ViT-B-32 model architecture provided by OpenCLIP [^3] https://github.com/mlfoundations/open_clip, with the pre-trained checkpoint laion2b_s12b_b32k [^4] https://hf.co/laion/CLIP-ViT-B-32-roberta-base-laion2B-s12B-b32k.
+Custom scripts were developed to adapt its format to the Transformers library, ensuring compatibility with our framework.
+The model parameters were optimized using AdamW (loshchilov2017decoupled), with a learning rate set to 1e-4 for the LoRA fine-tuning modules and 5e-4 for other trainable modules.
+A cosine annealing scheduler with warmup was employed to dynamically adjust the learning rate, where the warmup steps constituted the first 20% of the total optimization steps, and the minimum learning rate was set to 1% of the initial rate.
+For the modules $\mathcal{G}_{t/v}$, $\mathcal{F}_{t/v}$, $\mathcal{F}_{c}$, and $\mathcal{F}_{p}$, simple multi-layer perceptrons (MLPs) were utilized.
+The training processing was performed with a batch size of 64 for 3 epochs.
+All experiments were run on a machine equipped with an NVIDIA RTX 4090 GPU.
+Table 3. Summary of hyperparameters.
+Parameter Description
+$r$ The rank of LoRA, determining the dimension of the low-rank update matrices.
+$\mathbf{W}$ The weight matrices in the self-attention module fine-tuned using LoRA, specifically targeting combinations of $W_{\{q,k,v,o\}}$.
+top-$n$ The number of top self-attention layers conditioned during fine-tuning.
+$d_{f}$ The dimensionality of the latent space for the projected features.
+$\mathbf{L}$ The configurable range of memory sizes maintained by the Memory-Enhanced Predictor (MEP).
+Table 4. Hyperparameter settings.
+Parameter Value
+For trainer
+epoch 3
+batch_size 64
+lr 5e-4
+lora_lr 1e-4
+warmup_ratio 0.2
+min_lr_rate 0.01
+For our model
+$r$ 8
+top-$n$ 4
+$d_{f}$ 1024
+$\mathbf{W}$ $W_{k},W_{v},W_{o}$
+$\mathbf{L}$ {128, 256, 384, 512, 640, 768, 896, 1024, 1152, 1280}
+4.3. Hyperparameter Details
+We summarize the hyperparameters involved in InterCLIP-MEP and their descriptions in Table 4.
+The hyperparameter settings for obtaining the main results in the paper are summarized in Table 4.
+For each experiment, we condition only the top four layers of the self-attention modules, with the projection dimension $d_{f}$ set to 1024.
+We set the LoRA rank $r$ to 8, fine-tuning the self-attention weight matrices $\mathbf{W}$, specifically $W_{k}$, $W_{v}$, and $W_{o}$.
+For the memory size $L$, we select the optimal size from a fixed set of candidate values $\mathbf{L}$.
+For other baseline methods, we follow the optimal hyperparameter settings they reported.
+4.4. Main Results
+To begin our experimental analysis, we first evaluate the overall effectiveness of the proposed InterCLIP-MEP framework.
+To further assess its robustness, we design a control variant, denoted as w/o Inter, where the original CLIP is used as the backbone in place of InterCLIP.
+This variant is compared against three interaction modes of InterCLIP: w/ V2T, w/ T2V, and w/ TW.
+The overall results and baseline comparisons are summarized in Table 5, while the effectiveness of different interaction modes is further analyzed in Table 6.
+Performance on MMSD2.0
+As shown in Table 5, our framework consistently surpasses or matches state-of-the-art baselines on MMSD2.0, confirming the effectiveness of the proposed training strategy and the Memory-Enhanced Predictor (MEP).
+Table 6 further compares different interaction settings.
+Both w/ V2T and w/ T2V outperform the w/o Inter variant, indicating that InterCLIP more effectively captures cross-modal sarcasm cues than the original CLIP backbone.
+Among them, w/ T2V yields the best results.
+This is because projecting compact, low-dimensional text embeddings into the higher-dimensional vision space allows redundancy in visual features to accommodate textual semantics, preserving information more effectively.
+In contrast, projecting rich, high-dimensional vision embeddings into the text space requires compression into a compact semantic space, which inevitably loses details and hampers alignment.
+This asymmetry explains why T2V consistently outperforms V2T.
+Meanwhile, the w/ TW configuration performs worse than either single-direction variant, as simultaneously embedding both modalities exacerbates the alignment asymmetry and further increases the learning burden.
+In summary, InterCLIP with T2V interaction, together with our lightweight training strategy and MEP, achieves the most promising results on MMSD2.0, underscoring the robustness and adaptability of our framework in modeling subtle text–image interactions.
+Table 5. Baseline comparison on MMSD2.0 and MMSD, where InterCLIP-MEP uses the best interaction mode for each dataset.
+Method MMSD2.0 MMSD
+Acc. (%) F1 (%) P (%) R (%) Acc. (%) F1 (%) P (%) R (%)
+Text
+TextCNN (kim-2014-convolutional) 71.61* 69.52* 64.62* 75.22* 80.03* 75.32* 74.29* 76.39*
+Bi-LSTM (zhou-etal-2016-attention) 72.48* 68.05* 68.02* 68.08* 81.90* 77.53* 76.66* 78.42*
+SMSD (xiong2019sarcasm) 73.56* 69.97* 68.45* 71.55* 80.90* 75.82* 76.46* 75.18*
+RoBERTa (liu2019roberta) 79.66* 76.21* 76.74* 75.70* 93.97* 92.45* 90.39* 94.59*
+Image
+ResNet (he2015deep) 65.50* 57.58* 61.17* 54.39* 64.76* 61.53* 54.41* 70.80*
+ViT (dosovitskiy2020image) 72.02* 69.72* 65.26* 74.83* 67.83* 63.40* 57.93* 70.07*
+Text-Image
+HFM (Cai2019MultiModalSD) 70.57* 66.88* 64.84* 69.05* 83.44* 80.18* 76.57* 84.15*
+Att-BERT (pan-etal-2020-modeling) 80.03* 77.04* 76.28* 77.82* 86.05* 82.92* 80.87* 85.08*
+CMGCN (Liang2022MultiModalSD) 79.83* 76.90* 75.82* 78.01* 86.54* 84.09* -* -*
+HKE (Liu2022TowardsMS) 76.50* 72.25* 73.48* 71.07* 87.36* 72.25* 81.84* 86.48*
+DIP (wen2023dip) 80.59 78.23 75.52 81.14 89.59 87.17 87.76 86.58
+DynRT (Tian2023DynamicRT) 70.37 68.55 63.02 75.15 93.59 91.93 90.30 93.62
+Multi-view CLIP (qin2023mmsd2) 85.64* 84.10* 80.33* 88.24* 88.33* 85.55* 82.66* 88.65*
+G2SAM (wei2024g2sam) 79.43 78.07 72.04 85.20 90.48 88.48 87.95 89.02
+InterCLIP-MEP (Ours) 86.72 85.61 80.20 91.80 93.94 92.54 90.77 94.37
+    We use * to indicate that the results are taken from (qin2023mmsd2).
+indicates that results are not reported.
+$\dagger$ indicates our reproduced results.
+Underlined values represent the best multi-modal baseline for comparison.
+Bold values indicate those that surpass the underlined baseline.
+The InterCLIP-MEP results use the T2V interaction on MMSD2.0 and the w/o Inter setting on MMSD.
+These choices follow our analysis (Section 4.4, Table 6), which shows T2V is most effective on MMSD2.0, while MMSD’s spurious textual cues make w/o Inter more suitable.
+Table 6. Performance comparison of InterCLIP-MEP under different interaction settings.
+Method MMSD2.0 MMSD
+Acc. (%) F1 (%) P (%) R (%) Acc. (%) F1 (%) P (%) R (%)
+InterCLIP-MEP
+w/o Inter ($L_{2}=1024,L_{1}=1280$) 86.05 84.81 79.83 90.45 88.75 86.31 83.73 89.05
+w/ TW ($L_{2}=128,L_{1}=1152$) 85.51 84.26 79.15 90.07 88.54 86.32 82.25 90.82
+w/ V2T ($L_{2}=640,L_{1}=1024$) 86.26 85.00 80.17 90.45 88.92 86.66 83.21 90.41
+w/ T2V ($L_{2}=1024,L_{1}=1152$) 86.72 85.61 80.20 91.80 88.83 86.37 84.02 88.84
+InterCLIP-MEP w/ RoBERTa
+w/o Inter ($L_{2}=640,L_{1}=128$) 77.21 75.55 70.20 81.77 93.94 92.54 90.77 94.37
+w/ TW ($L_{2}=896,L_{1}=256$) 81.98 80.78 74.69 87.95 93.73 92.28 90.48 94.16
+w/ V2T ($L_{2}=640,L_{1}=512$) 76.96 75.26 69.98 81.39 93.94 92.54 90.69 94.47
+w/ T2V ($L_{2}=1024,L_{1}=384$) 82.81 81.55 75.81 88.24 93.73 92.28 90.56 94.06
+    Bold values indicate those that surpass the underlined baseline. $L_{2}$ for MMSD2.0 and $L_{1}$ for MMSD denote the optimal MEP memory sizes.
+Performance on MMSD
+Table 7.
+Performance comparison on the MMSD dataset in terms of Macro-average F1 score, following the standard evaluation protocol used in recent literature (zhuang2025multi).
+Results marked with * are reported from (zhuang2025multi).
+Method Macro-F1 (%)
+DIP (wen2023dip)* 89.01
+G2SAM (wei2024g2sam)* 89.65
+KFGC-Net (zhuang2025multi)* 91.02
+InterCLIP-MEP 93.72
+As shown in Table 5, the RoBERTa-based text baseline significantly outperforms other methods on MMSD due to spurious textual cues, which allow models to achieve high accuracy using text features alone (qin2023mmsd2).
+Consequently, multi-modal models such as DynRT (Tian2023DynamicRT), G2SAM (wei2024g2sam), and DIP (wen2023dip), which rely on RoBERTa or BERT for text encoding, also obtain strong results on MMSD but exhibit a substantial performance drop on MMSD2.0.
+To further investigate this effect, we introduce an additional variant, InterCLIP-MEP w/ RoBERTa, by replacing the original text encoder with RoBERTa (Table 6).
+Although this variant achieves state-of-the-art performance on MMSD, it disrupts InterCLIP’s modality alignment and thus leads to a noticeable decline on MMSD2.0.
+These findings suggest that MMSD’s text data contains spurious cues that encourage over-reliance on textual signals, whereas MMSD2.0, being cleaned, requires stronger multi-modal reasoning.
+We also observe that the w/ V2T variant performs best in both InterCLIP-MEP and InterCLIP-MEP w/ RoBERTa, highlighting the model’s tendency to over-depend on textual features.
+Following (zhuang2025multi), we also report the Macro-F1 score on the MMSD dataset to ensure consistency with community evaluation standards.
+As shown in Table 7, InterCLIP-MEP achieves the highest Macro-F1 of 93.72%, outperforming DIP (wen2023dip), G2SAM (wei2024g2sam), and KFGC-Net (zhuang2025multi).
+Table 8. Efficiency comparison of different methods.
+Method Accuracy (%) Trainable Parameters (M) Fitting Time / Epoch (s) Inference Time (s) GPU Memory Peak (GB)
+Multi-view CLIP  85.64  165  488  51  15.59
+DIP  80.59  196  OOM  OOM  OOM
+G2SAM  79.43  116  90  13  18.32
+DynRT  70.37  25  370  26  8.03
+InterCLIP-MEP  86.72  8  55  6  6.14
+    To demonstrate the efficiency of InterCLIP-MEP, we selected several recent baselines for comparison.
+The analysis was conducted using the MMSD2.0 dataset on a single NVIDIA RTX 4090 GPU with a batch size of 128.
+In the table, Fitting Time / Epoch indicates the time required for each epoch during training and validation and OOM indicates Out of Memory, referring to GPU memory overflow.
+Efficiency Comparison.
+Our training strategy demonstrates both remarkable effectiveness and outstanding efficiency.
+To validate this, we conducted a comprehensive comparative analysis against leading state-of-the-art methods, as detailed in Table 8.
+For instance, the Multi-view CLIP method (qin2023mmsd2) employs a multi-layer Transformer encoder for feature fusion, which, while effective, introduces a significant number of trainable parameters.
+This results in slower training and inference speeds and greater memory consumption.
+Similarly, the DIP method (wen2023dip) caches historical samples during training, which hinders its ability to support large-batch training under limited resource conditions.
+In contrast, our method operates with a batch size of 128 while utilizing a significantly smaller number of trainable parameters, which translates to notably faster training and validation cycles.
+Furthermore, by incorporating minimal parameter modifications to adapt CLIP and utilizing simple yet effective linear layers for representation fusion, our approach achieves superior inference speeds and drastically reduced memory consumption.
+These results highlight the practicality of our framework, establishing it as a benchmark for both computational efficiency and performance in multi-modal sarcasm detection.
+Table 9. Ablation study of InterCLIP-MEP.
+w/o Inter w/ TW w/ V2T w/ T2V
+Variant Acc. (%) F1 (%) Acc. (%) F1 (%) Acc. (%) F1 (%) Acc. (%) F1 (%)
+Baseline 86.05 84.81 85.51 84.26 86.26 85.00 86.72 85.61
+w/o Proj 85.76 84.43 85.43 84.05 85.68 84.22 86.22 84.51
+w/o MEP 85.39 83.99 85.22 83.79 86.26 84.78 86.26 84.82
+w/o LoRA 82.44 77.73 76.42 74.37 73.31 72.22 75.13 71.79
+    Baseline denotes the results without ablation.
+4.5. Analysis of InterCLIP-MEP
+To robustly validate the effectiveness of InterCLIP-MEP, we conduct comprehensive ablation studies and case studies on the more reliable MMSD2.0 benchmark, offering deeper insights into its design and performance.
+In addition, we include visualization analyses to provide an intuitive understanding of how the framework processes multi-modal sarcasm cues.
+Ablation study.
+We remove the projection module $\mathcal{F}_{p}$ and train only the classification module $\mathcal{F}_{c}$ for prediction, denoted as w/o Proj.
+To test the necessity of using LoRA (edward2022lora) for fine-tuning, we keep the rest of InterCLIP-MEP unchanged and freeze all self-attention weights of InterCLIP, denoted as w/o LoRA, always selecting the optimal memory size $L$ for the MEP during inference.
+To evaluate the effectiveness of the MEP, we train both $\mathcal{F}_{p}$ and $\mathcal{F}_{c}$ but use only $\mathcal{F}_{c}$ during inference, denoted as w/o MEP.
+Table 9 reports all results.
+All variants show performance declines compared to the baseline, demonstrating the importance of each module in InterCLIP-MEP.
+For w/ TW and w/o Inter, the w/o MEP variant performed worse than the w/o Proj variant.
+However, for w/ T2V and w/ V2T, the w/o MEP variant performs better than the w/o Proj variant.
+This suggests that backbones with strong image-text interaction capabilities benefit from training the classification module $\mathcal{F}_{c}$ along with the projection module $\mathcal{F}_{p}$, even without using MEP during inference.
+We also find that not using LoRA to fine-tune the self-attention modules results in significant performance loss, indicating that the original CLIP’s vanilla space is not suitable for the sarcasm detection task.
+Figure 4. Hyperparameter study curves for w/ T2V. Panel (d) compares results with those from using only the classification module $\mathcal{F}_{c}$ for prediction.
+Figure 5.
+Case study of InterCLIP-MEP.
+In the figure, the emojis in the Sample column denote the ground-truth labels from the dataset.
+MEP represents the labels predicted by the memory-enhanced predictor, and $\mathcal{F}_{c}$ represents the labels predicted by the classification module.
+Hyperparameter study.
+We further investigate the method using Interactive-CLIP with T2V interaction as the backbone.
+Keeping the other hyperparameters constant, we condition different top-$n$ layers of the self-attention modules.
+We also study the impact of different projection dimensions $d_{f}$, different LoRA ranks $r$, and different memory sizes $L$ on the w/ T2V method.
+We present all results in Figure 5.
+Figure 5(a) shows that conditioning the top four self-attention layers yields the best results.
+From Figure 5(b), a rank of 8 is optimal.
+Figure 5(c) indicates the projection dimension is best at 256 or 1024.
+Figure 5(d) reveals that a memory size of 640 in MEP outperforms the others, confirming the value of historical sample knowledge.
+Figure 6.
+Visual examples showcasing InterCLIP’s improved focus on sarcasm-related visual cues compared to CLIP.
+In the figure, the first column shows the original input images, the second and third columns present the GradCAM visualizations from InterCLIP and the original CLIP respectively, and the last column provides human-readable explanations comparing their behaviors.
+Case study.
+As shown in Figure 5, we provide illustrative examples to examine both the strengths and the limitations of InterCLIP-MEP.
+In the first two cases, direct predictions from the classification module $\mathcal{F}_{c}$ yield uncertain and incorrect outcomes with relatively high entropy, whereas the Memory-Enhanced Predictor (MEP) leverages historical sample knowledge to successfully correct the results.
+These examples highlight the robustness and effectiveness of our design.
+At the same time, we also include two challenging failure cases.
+In the third example, $\mathcal{F}_{c}$ makes a correct prediction, while MEP mistakenly overrides it.
+This situation arises when the memory stores similar but misleading representations, which can bias the final decision.
+In the fourth example, both $\mathcal{F}_{c}$ and MEP fail, reflecting the intrinsic difficulty of multi-modal sarcasm detection when subtle ironic cues are either extremely implicit or context-dependent beyond the available modalities.
+These observations naturally raise two questions: how does the memory mechanism handle potential error propagation, and how robust is it under distribution shifts?
+To answer these, we provide dedicated analyses in Section 5.3 and Section 5.4, where we examine these challenges in a systematic manner.
+With these results in mind, it is important to note that the identified failures do not undermine the overall effectiveness of InterCLIP-MEP, but rather point to promising directions for refinement.
+For instance, incorporating context-aware memory update strategies or external commonsense knowledge may further reduce such errors.
+Overall, the inclusion of both successful and unsuccessful cases provides a more balanced perspective, demonstrating that while our method substantially improves robustness, multi-modal sarcasm detection remains a highly challenging task.
+Visualization.
+To further validate that InterCLIP is capable of more effectively capturing the interactive information between text and images compared to original CLIP, thereby aiding in the detection of sarcasm cues, we use GradCAM (Selvaraju2017GradCAM) to visualize the areas of focus during the inference process of the visual model in Figure 6.
+These visualizations highlight InterCLIP’s improved ability to capture sarcasm-related cues by focusing on relevant areas in the images.
+4.6. Empirical Study of Self-attention Fine-tuning and Interaction Modes
+Keeping the other hyperparameters constant as shown in Table 4, we fine-tune all possible weight matrices $\mathbf{W}$ and employ different interaction modes of InterCLIP as the backbone.
+We consistently select the optimal memory size from $\mathbf{L}$ for MEP.
+We calculate the average metrics for different methods and different weight matrices.
+The results are presented in Table 12 and Table 12.
+We observe that fine-tuning the weight matrices $W_{k},W_{v},W_{o}$ and using the T2V interaction mode of InterCLIP are the best choices for InterCLIP-MEP.
+Table 10. Average results of fine-tuning different weight matrices $\mathbf{W}$ across four baseline methods.
+$\mathbf{W}$ Mean Acc. (%) Mean F1 (%)
+$W_{q}$ 85.14 83.99
+$W_{k}$ 85.09 84.00
+$W_{v}$ 85.34 84.24
+$W_{o}$ 85.39 84.23
+$W_{q},W_{k}$ 85.36 84.16
+$W_{q},W_{v}$ 85.63 84.40
+$W_{q},W_{o}$ 85.73 84.49
+$W_{k},W_{o}$ 85.67 84.43
+$W_{v},W_{o}$ 85.73 84.54
+$W_{k},W_{v}$ 85.70 84.51
+$W_{q},W_{k},W_{o}$ 85.92 84.63
+$W_{q},W_{v},W_{o}$ 85.87 84.60
+$W_{q},W_{k},W_{v}$ 86.05 84.75
+$W_{k},W_{v},W_{o}$ 86.14 84.92
+$W_{q},W_{k},W_{v},W_{o}$ 85.82 84.55
+Table 11. Average results of four baseline methods for fine-tuning different weight matrices $\mathbf{W}$.
+Method Mean Acc. (%) Mean F1 (%)
+w/o Inter 85.49 84.32
+w/ TW 85.66 84.44
+w/ V2T 85.62 84.41
+w/ T2V 85.78 84.55
+Table 12. Statistics of DocMSU.
+Sarcastic Non-sarcastic All
+Train 4,014 46,265 50,279
+Validation 1,125 13,097 14,222
+Test 555 6,772 7,327
+5. Extended Experiments
+To further verify the performance and robustness of our framework, we extended our evaluation in three complementary directions.
+First, we conducted experiments on a document-level benchmark to examine its capability in handling long-text sarcasm detection.
+Second, we performed low-resource experiments to assess generalization and adaptability under data scarcity.
+Third, we systematically investigated the impact of our memory mechanism under error propagation and distribution shift conditions, thereby providing deeper insights into its stability and reliability.
+5.1. Document-level Benchmark
+DocMSU Benchmark
+DocMSU (du2024docmsu) is a recently introduced multi-modal sarcasm benchmark designed specifically for long-text analysis.
+It facilitates the evaluation of multi-modal sarcasm comprehension as well as detection tasks.
+In this study, we concentrate on the multi-modal sarcasm detection task.
+The benchmark statistics can be found in Table 12.
+Baselines
+We follow the evaluation protocol of (du2024docmsu).
+We compare against unimodal baselines: BERT-base (text-only) (devlin2019bert) and Swin Transformer (image-only) (liu2021swin).
+For multi-modal approaches, we include CLIP (radford2021learning), Vision-and-Language Transformer (ViLT) (kim2021vilt), and CMGCN (Liang2022MultiModalSD).
+The method proposed by (du2024docmsu) is taken as the state-of-the-art baseline.
+Results
+Table 13. Results on DocMSU.
+Method Acc. F1 P R
+BERT-base* 87.12 86.51 77.61 70.37
+Swin-Transformer* 74.83 61.51 67.57 56.45
+CMGCN* 88.12 75.23 78.11 72.55
+CLIP* 96.19 77.62 78.99 76.30
+ViLT* 93.15 41.44 69.03 29.61
+DocMSU Baseline* 97.83 87.25 81.20 94.27
+InterCLIP-MEP (Ours)
+w/o Inter 97.84 87.48 78.08 99.45
+w/ TW 97.79 87.24 77.48 99.81
+w/ V2T 97.83 87.45 77.81 99.82
+w/ T2V 97.67 86.65 76.45 99.99
+    Underline results denote the compared SOTA baseline, boldface highlights results that surpass the baseline, and * indicates results sourced from (du2024docmsu).
+Our InterCLIP-MEP framework demonstrates strong performance across various configurations, as shown in Table 13.
+In particular, the w/o Inter and w/ V2T variants consistently achieve higher F1 scores compared to the baselines, thereby showcasing their robustness in handling multi-modal sarcasm detection tasks.
+Notably, all variants achieve nearly perfect recall, highlighting their outstanding capability in accurately identifying sarcasm across diverse datasets.
+The w/o Inter variant also achieves the highest accuracy, further demonstrating its effectiveness and precision.
+Overall, the comprehensive results in Table 13 affirm the unparalleled effectiveness of InterCLIP’s modality interaction mechanism and our proposed memory-enhanced predictor (MEP), particularly in surpassing existing state-of-the-art methods in both accuracy and F1 score, thus setting a new benchmark in the field.
+5.2. Low-resource Setting
+To assess the robustness of InterCLIP-MEP under low-resource conditions, we followed the experimental protocol of Multi-view CLIP (qin2023mmsd2) and conducted experiments on the MMSD2.0 dataset.
+Specifically, we subsampled the training set by selecting both positive and negative samples at equal intervals of 10%, and employed the hyperparameter configuration that yielded the best performance in Table 5.
+The models were then evaluated on the test set, and the results are presented in Figure 9.
+Figure 7. Accuracy on MMSD2.0 with different training ratios.
+Figure 8. Accuracy under different levels of pseudo-label noise. The dashed line indicates the result without MEP.
+Figure 9. Evaluation on a mixed MMSD2.0–DocMSU test set illustrates that interactive representations and the memory mechanism jointly enhance robustness under distributional shifts.
+As shown in the figure, when 50% or more of the training data is available, InterCLIP-MEP achieves a clear advantage over Multi-view CLIP.
+This indicates that our framework is capable of leveraging moderately limited supervision to reconstruct a more effective multi-modal representation space.
+Moreover, while our method performs slightly worse than Multi-view CLIP when training with less than 50% of the data, it exhibits an emergent performance gain once this threshold is reached.
+Specifically, the accuracy of InterCLIP-MEP increases by 14.24% from its lowest to highest point, whereas Multi-view CLIP only improves by 2.91%.
+We attribute this phenomenon to the distinct training strategies.
+Multi-view CLIP relies on original CLIP representations, which provide relatively strong performance even with limited supervision but yield diminishing returns as more data becomes available.
+In contrast, InterCLIP-MEP employs lightweight fine-tuning to restructure the representation space of CLIP, making it more aligned with the specific requirements of multi-modal sarcasm detection.
+This process requires a sufficient amount of data to stabilize the new cross-modal alignment, but once the data threshold is met, the framework benefits substantially more from additional training samples.
+5.3. Error Propagation Robustness Study
+To further assess the robustness of the proposed memory mechanism against error propagation, we designed a stress test by randomly flipping a certain proportion of pseudo-labels generated by the classification module.
+This operation introduces deliberately corrupted low-entropy predictions into the memory, thereby simulating potential error accumulation during inference.
+For consistency, all hyperparameters in this experiment follow the optimal settings reported in Table 5.
+The results are shown in Figure 9.
+Even when up to 40% of pseudo-labels are randomly flipped, the accuracy of InterCLIP-MEP remains consistently above the baseline system without memory (86.26%).
+The observed variations are minimal, indicating that the memory-enhanced predictor maintains stable performance under noisy pseudo-labels.
+These findings demonstrate that the memory-enhanced predictor does not suffer from cascading error amplification.
+Instead, it maintains stable performance under noisy pseudo-labels, confirming that the dynamic update and entropy-based filtering strategy effectively suppress the influence of mislabeled entries.
+5.4. Robustness under Distributional Shift
+To rigorously examine robustness under distributional shifts, we compared InterCLIP-MEP, its variant without the memory mechanism (w/o MEP), and the state-of-the-art Multi-view CLIP baseline.
+All models were trained on the MMSD2.0 (qin2023mmsd2) training set and evaluated on a mixed testing set constructed by shuffling the MMSD2.0 and DocMSU (du2024docmsu) test sets.
+This heterogeneous evaluation setting introduces significant domain variation beyond the training distribution and provides a stringent test of model stability.
+As shown in Figure 9, InterCLIP-MEP attains the highest accuracy of 73.96%, substantially outperforming its variant without the memory mechanism (63.97%) and Multi-view CLIP (58.62%).
+The nearly 10% improvement over the w/o MEP variant demonstrates that the dynamic memory mechanism significantly mitigates performance degradation by suppressing error accumulation and leveraging historical knowledge during inference.
+Moreover, even without the memory component, InterCLIP itself still surpasses Multi-view CLIP by more than 5%, highlighting the effectiveness of the interactive representation design.
+These results confirm that InterCLIP-MEP achieves superior stability and adaptability when faced with distributional shifts.
+6. Limitations and Future Work
+While InterCLIP-MEP demonstrates strong robustness and consistent improvements across multiple evaluations, several limitations remain that highlight promising directions for future research.
+Limitations.
+Despite its effectiveness, InterCLIP-MEP can be sensitive to memory noise when misleading or weakly relevant representations are retrieved, which may bias the inference process and occasionally override correct predictions.
+Moreover, the framework still struggles with highly implicit or context-dependent sarcasm, especially when such cues rely on external commonsense or situational background that is not explicitly encoded in the available modalities.
+In addition, the current design assumes a relatively static memory and homogeneous data distribution, which may limit adaptability when facing unseen domains or evolving data patterns.
+Future Work.
+Future studies may explore context-aware memory updating strategies that dynamically adjust to semantic drift and data uncertainty.
+Integrating external commonsense or world knowledge could further enhance the model’s ability to capture subtle pragmatic cues and hidden irony.
+Another potential direction is to introduce memory regularization or selective forgetting mechanisms to mitigate the effect of noisy retrievals and reduce error propagation under domain shifts.
+Finally, extending the proposed memory-augmented paradigm to other multimodal reasoning tasks, such as emotion understanding or visual entailment, could provide a broader understanding of how long-term memory interacts with perception and language in open-world scenarios.
+7. Conclusion
+This paper introduces InterCLIP-MEP, an innovative framework designed to address the critical challenges of modeling subtle text-image interactions and mitigating prediction uncertainty in multi-modal sarcasm detection.
+The proposed framework incorporates Interactive CLIP (InterCLIP), which integrates cross-modal information directly into text and image encoders.
+This enhances the model’s ability to interpret multi-modal sarcasm cues.
+Furthermore, a Memory-Enhanced Predictor (MEP) is developed to dynamically utilize knowledge from historical samples, resulting in a more robust and adaptive inference process.
+Extensive evaluations conducted on the MMSD, MMSD2.0, and DocMSU benchmarks demonstrate that InterCLIP-MEP achieves state-of-the-art performance while substantially reducing computational overhead.
+By requiring fewer trainable parameters and less GPU memory, InterCLIP-MEP offers a lightweight, efficient, and scalable solution, setting a new benchmark for multi-modal sarcasm detection.
+Acknowledgements
+The research presented in this paper was funded by the Anhui Provincial Natural Science Foundation under Grant No.
+2308085MF220.
+It also received additional support from the Anhui University Natural Science Foundation under Grant No.
+2023AH050914.